### v1

# Interacting phantom dark energy: new accelerating scaling attractors

Abstract
We perform a detailed investigation of interacting phantom cosmology, by applying the powerful method of dynamical system analysis. We consider two well-studied interaction forms, namely one global and one local one, while the novel ingredient of our work is the examination of new potentials for the phantom field. Our analysis shows the existence of saddle matter-dominated points, stable dark-energy dominated points, and scaling accelerating solutions, that can attract the Universe at late times. As we show, some of the stable accelerating scaling attractors, in which dark matter and dark energy can co-exist, alleviating the cosmic coincidence problem, are totally new, even for the previously studied interaction rates, and arise purely from the novel potential forms.
pacs:
98.80.-k, 95.36.+x, 95.35.+d, 98.80.Es
I Introduction
Dark matter (DM) and dark energy (DE) are two main ingredients of the universe, constituting of 95% of its total energy budget. The DM sector is responsible for the observed structure formation, while the DE sector is in charge of the late time accelerated expansion Copeland et al. (2006); Bamba et al. (2012). The nature, origin and dynamics of both DE and DM are not yet clearly understood, nevertheless the $\Lambda$-Cold Dark Matter ($\Lambda$CDM) cosmological model, based on General Relativity, matches quite well with a large span of the astronomical datasets in which DE is a simple cosmological constant and DM is a cold pressureless sector. However, $\Lambda$CDM faces many challenges at both theoretical and observational levels Di Valentino et al. (2021a); Perivolaropoulos and Skara (2022); Abdalla et al. (2022), and among possible improvements is the consideration of a time varying nature in the DE sector.
In general, DE is assumed to evolve independently and interact only with gravity. However, there is no fundamental reason to exclude possible interactions between DE and other fields, especially with the DM sector whose microscopic nature is also unknown. The interaction between DE and DM can alleviate the cosmic coincidence problem Amendola (2000); Chimento et al. (2003); Cai and Wang (2005); Pavon and Zimdahl (2005); Huey and Wandelt (2006); Hu and Ling (2006); del Campo et al. (2008, 2009), while it can potentially alleviate the current cosmological tensions too Di Valentino et al. (2021a); Perivolaropoulos and Skara (2022); Abdalla et al. (2022); Kumar and Nunes (2017); Di Valentino et al. (2017); Yang et al. (2018); Kumar et al. (2019); Pan et al. (2019a, b); Di Valentino et al. (2020a); Pan et al. (2020a); Pan and Yang (2023). Hence, over the last years interacting DM-DE scenarios have gained significant attention in the literature for their attractive features Guo et al. (2005); Sussman et al. (2005); Zimdahl (2005); Wang et al. (2005, 2006); Barrow and Clifton (2006); Setare (2006); Poplawski (2006); Sadjadi and Honardoost (2007); Chang et al. (2007); Kim et al. (2007); Rosenfeld (2007); Zimdahl and Pavon (2007); Feng et al. (2007); Cruz et al. (2008); Chen et al. (2008); Jamil et al. (2010); Jackson et al. (2009); Valiviita et al. (2010); Suwa and Nihei (2010); Karami and Ghaffari (2010); Wei (2010); Martinelli et al. (2010); Gavela et al. (2010); Lopez Honorez et al. (2010); Baldi (2011); Chen et al. (2014); Baldi and Salucci (2012); Chimento et al. (2013); Harko and Lobo (2013); Sun and Yue (2013); Li and Zhang (2014); Tamanini et al. (2014); Yang and Xu (2014); Li et al. (2014); Pan et al. (2015); Duniya et al. (2015); Li et al. (2016); Odderskov et al. (2016); Nunes et al. (2016); Ferreira et al. (2017); Biswas et al. (2017); Costa et al. (2017); Linton et al. (2018); Yang et al. (2019); Di Valentino et al. (2020b); Li et al. (2020); Yang et al. (2020); Pan et al. (2020b, c); Di Valentino et al. (2021b); Yang et al. (2021); Gao et al. (2021); Chatzidakis et al. (2022); Hou et al. (2023); Pan et al. (2023); Yang et al. (2023); Zhai et al. (2023); Bernui et al. (2023); Escamilla et al. (2023); Benisty et al. (2024); Halder et al. (2024a); Giarè et al. (2024a, b); Sabogal et al. (2024); Ghedini et al. (2024); Paliathanasis et al. (2024) (for reviews see Bolotin et al. (2014); Wang et al. (2016)).
The interacting DM-DE models, widely known as Interacting DE (IDE) or Coupled DE, induce a richer dynamics in the universe evolution, recovering the non-interacting scenarios as special cases. The heart of the IDE models is the interaction function which characterizes the rate of energy and/or momentum transfer between the dark components, and consequently it affects the evolution of the universe. In order to investigate the dynamics of interacting cosmologies, one can confront them with observational datasets from various astronomical surveys and extract constraints on the involved functions and parameters. However, one can also apply the powerful approach of dynamical systems, which allows to extract analytical information independently of the initial conditions. The techniques of dynamical system analysis have been widely applied in the context of non-interacting and interacting dynamics Copeland et al. (1998); Billyard and Coley (2000); Gong et al. (2006); Chen and Gong (2009); Chen et al. (2009); Boehmer et al. (2008); Karwan (2008); Caldera-Cabral et al. (2009); Leon et al. (2009); Leon and Saridakis (2010); Xu et al. (2012); Leon and Saridakis (2013); Avelino et al. (2013); Boehmer et al. (2015a, b); Biswas and Chakraborty (2015a); Fadragas et al. (2014); Biswas and Chakraborty (2015b); Mahata and Chakraborty (2015); Banerjee and Roy (2015); Paliathanasis et al. (2015, 2016); Singh and Singh (2016); Dutta et al. (2017); Carneiro and Borges (2018); Odintsov et al. (2017); Zonunmawia et al. (2017); Bahamonde et al. (2018); Odintsov and Oikonomou (2018); Paliathanasis et al. (2019); Hernández-Almada et al. (2020); Sá (2020); Chakraborty et al. (2022); Paliathanasis et al. (2021); Sá (2021); Samart et al. (2021); Arevalo and Cid (2022); Sá (2024); Halder et al. (2024b).
In the present work we are interested in performing a dynamical-system analysis of interacting cosmology considering a phantom dark energy Caldwell (2002); Dabrowski et al. (2003); Nojiri et al. (2005); Singh et al. (2003); Capozziello et al. (2006); Cai et al. (2010); Saridakis (2009). In particular, we consider various interacting phantom DE scenarios, characterized by four different potentials and two specific interacting functions, one corresponding to the global interaction rate and the other corresponding to the local interaction rate. The examination of new potentials even for the known and studied cases of interaction form reveals new accelerating scaling attractors that were never studied in the literature. This is an important result, since it was believed that in general, due to the phantom nature of the acceleration, accelerating scaling solutions in phantom cosmology were very rare.
The manuscript is structured as follows. In section II we present the basic equations of general IDE scenarios and we construct the various specific models. In section III we perform the detailed phase space analysis extracting all possible solutions. Finally, section IV is devoted to the conclusions.
II Phantom Interacting Dark Energy
In this section we briefly present the scenario of phantom interacting dark energy, providing the main equations. We consider a flat Friedmann-Lemaître-Robertson-Walker (FLRW) line element given by
$ds^{2}=-dt^{2}+a^{2}(t)\left[dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})\right],$ (1)
where $a(t)$ is the expansion scale factor of the universe. Concerning the gravitational theory we consider standard general relativity, and we additionally consider the standard (dark plus baryonic) matter sector of the universe corresponding to a perfect fluid. Concerning the dark energy sector, we attribute it to a phantom scalar field $\phi$ with action
$S_{\phi}=\int d^{4}x\sqrt{-g}\left[\frac{1}{2}(\nabla\phi)^{2}-V(\phi)\right],$ (2)
where $V(\phi)$ is the potential. In the case of FLRW geometry, the energy density $\rho_{\phi}$ and pressure $p_{\phi}$ of the phantom fluid are given by
$\rho_{\phi}=-\frac{\dot{\phi}^{2}}{2}+V(\phi),$ (3)
$p_{\phi}=-\frac{\dot{\phi}^{2}}{2}-V(\phi),$ (4)
where a dot denotes derivative with respect to time. Note that the energy density of the phantom scalar field is unbounded from below and thus it may take negative values Bahamonde et al. (2018). Thus, the Friedmann equations are written as
$3H^{2}=\kappa^{2}(\rho_{m}+\rho_{\phi})$ (5)
$2\dot{H}+3H^{2}=-\kappa^{2}(p_{m}+p_{\phi}),$ (6)
where $H=\dot{a}/a$ is the Hubble function (dots denote time-derivatives), $\kappa^{2}=8\pi G$ is the gravitational constant, and $\rho_{m}$, $p_{m}$ are respectively the energy density and pressure of the matter fluid.
The key idea of interacting cosmology is that although the total energy content of the universe is conserved (as a requirement of the diffeomorphism invariance) the separate DE and DM sectors are not conserved, but mutually interact. In particular, their conservation equations take the general form
$\dot{\rho}_{m}+3H(p_{m}+\rho_{m})=-Q,$ (7)
$\dot{\rho}_{\phi}+3H(\rho_{\phi}+p_{\phi})=Q,$ (8)
where $Q$ is the interaction function which characterizes the transfer of energy between the dark sectors. For $Q>0$, the energy flow occurs from DM to DE, while for $Q<0$ energy flows in the reverse direction, namely from DE to DM. Notice that eqn. (8) can alternatively be expressed in a Klein-Gordon-like form, namely
$\ddot{\phi}+3H\dot{\phi}-\frac{\partial V(\phi)}{\partial\phi}=-\frac{Q}{\dot{\phi}}.$ (9)
In this way, one can see that if the interaction function $Q(t)$ and the potential $V(\phi)$ of the scalar field are prescribed, then the evolution of the interacting fluids can be in principle determined. Hence, the choice of both of these functions is the starting point for any analysis.
In the following we focus on two interacting functions. The first interaction model is based on the most well-studied interaction form, namely Bolotin et al. (2014); Wang et al. (2016)
$\text{Model I}:\ \ \ Q=\alpha H\rho_{m},$ (10)
where $\alpha$ is a dimensionless parameter. In this scenario the interaction rate is proportional to the energy density, which is expected to be the case since interactions in physics and chemistry are typically stronger if the number density of the involved particles increases. Nevertheless, the interaction rate is proportional to the Hubble function too, which although very convenient at the calculation stage raises the criticism that the DM-DE interaction rate, which is expected to depend on local physics, depends on a global feature of the whole Universe, namely the Hubble function Shafieloo et al. (2018). Hence, in this work we will additionally consider a second interaction function of the local form Shafieloo et al. (2018); Li et al. (2019)
$\text{Model II}:\ \ \ Q=\Gamma\rho_{m},$ (11)
where $\Gamma$ is the dimensionfull coupling parameter. Finally, concerning the potential choices, we consider two forms that have been widely used in the literature, namely the well known exponential potential Copeland et al. (1998); Amendola (2000); Elizalde et al. (2004); Chen et al. (2009); Basilakos et al. (2011); Halder et al. (2024b) and the hyperbolic potential Paliathanasis et al. (2015), given by
$V(\phi)=V_{0}e^{-\kappa\lambda\phi},$ (12)
$V(\phi)=V_{0}\cosh(\kappa\eta\phi),$ (13)
where $V_{0}>0$, and $\lambda$, $\eta$ are dimensionless free parameters. Another interesting potential form arises from the conformal transformation of Starobinsky $R^{2}$ gravity, and takes the form Ivanov et al. (2022)
$V(\phi)=\frac{1}{8\beta}\left(1-e^{-\kappa\mu\phi}\right)^{2},$ (14)
with $\mu$, $\beta$ dimensionless parameters. Finally, in order to cover the more general case $R+R^{2}+\Lambda$, we will also consider the potential Sebastiani et al. (2014)
$V(\phi)=c_{0}+c_{1}e^{\sqrt{\frac{2\kappa^{2}}{3}}\phi}+c_{2}e^{2\sqrt{\frac{2\kappa^{2}}{3}}\phi},$ (15)
with $c_{0}$,$c_{1}$,$c_{2}$, the model parameters. Note that the last two potentials have not been studied in detail in the framework of phantom cosmology.
In summary, the background evolution of the scenario at hand is determined by the two Friedmann equations (5),(6), together with the conservation equations (7),(8) and the above choices for $Q$ and $V$. As usual, one can introduce the dark matter and dark energy density parameters as
$\Omega_{\phi}$$\equiv$ $\frac{\kappa^{2}\rho_{\phi}}{3H^{2}}$ (16)
$\Omega_{m}$$\equiv$ $\frac{\kappa^{2}\rho_{m}}{3H^{2}}\,,$ (17)
as well as the total, effective, equation-of-state parameter as
$w_{tot}$$\equiv$ $\frac{p_{\phi}+p_{m}}{\rho_{\phi}+\rho_{m}}=\frac{-\frac{1}{2}\dot{\phi}^{2}-V+w_{m}\rho_{m}}{-\frac{1}{2}\dot{\phi}^{2}+V+\rho_{m}}\,,$ (18)
with $w_{m}\equiv p_{m}/\rho_{m}$ the matter equation-of-state parameter.
III Dynamical system analysis
In this section we will perform a detailed phase-space and stability analysis of interacting phantom cosmology. This approach allows one to obtain information on the global dynamics and the behavior of a cosmological scenario, independently of the initial conditions and the specific universe evolution Copeland et al. (1998). As usual, one first transforms the equations into an autonomous dynamical system, and extracts its critical points. Then, for each critical point one determines its stability by examining the sign of the eigenvalues of the perturbation matrix. The dynamical system analysis of interacting phantom cosmology has been performed in various articles Guo et al. (2005); Chen et al. (2009); Leon and Saridakis (2010); Shahalam et al. (2017); Chakraborty et al. (2022); Halder et al. (2024b), however it was restricted in specific potential choices. In this work we will perform it considering new potential forms and interaction models. In the following we focus on the case $w_{m}=0$, namely DM is pressureless and cold, and we examine the various potentials and interaction forms separately.
Point $x$ $y$ Existence $E_{1}$ $E_{2}$ Stability $\Omega_{\phi}$ $w_{tot}$
$A_{1}$ $-\frac{\lambda}{\sqrt{6}}$ $\sqrt{1+\frac{\lambda^{2}}{6}}$ $\forall$ $\alpha$, $\lambda$ $-(3+\frac{\lambda^{2}}{2})$ $-(\lambda^{2}+\alpha+3)$ stable node for $-\alpha<3+\lambda^{2}$ $1$ $-1-\frac{\lambda^{2}}{3}$
for $-3<\alpha<0$
$A_{2}$ $\frac{3+\alpha}{\sqrt{6}\lambda}$ $\sqrt{-\frac{\alpha}{3}-\frac{(3+\alpha)^{2}}{6\lambda^{2}}}$ $\lambda^{2}\geq\frac{(3+\alpha)^{2}}{-\alpha}$ $\frac{R+\sqrt{S}}{4(3+\alpha)\lambda^{2}}$ $\frac{R-\sqrt{S}}{4(3+\alpha)\lambda^{2}}$ stable node for $\alpha<-3$ $-\frac{\alpha\lambda^{2}+(\alpha+3)^{2}}{3\lambda^{2}}$ $\frac{\alpha}{3}$
and for $\alpha<-3$ with $-(\alpha+3)>\lambda^{2}\geq\frac{(\alpha+3)^{2}}{-\alpha}$
$-(\alpha+3)\geq\lambda^{2}\geq\frac{(\alpha+3)^{2}}{-\alpha}$ (see Fig. 1)
Table 1: The critical points for the case of scalar-field potential $V(\phi)=V_{0}e^{-\kappa\lambda\phi}$ and interaction rate $Q=\alpha H\rho_{m}$, alongside their existence, eigenvalues at these points and stability conditions. Additionally, we present the corresponding values of the dark-energy density parameter $\Omega_{\phi}$ and the total equation-of-state parameter $w_{tot}$. We have defined $S=-1944\lambda^{2}-3240\alpha\lambda^{2}-2160\alpha^{2}\lambda^{2}-720\alpha^{3}\lambda^{2}-120\alpha^{4}\lambda^{2}-8\alpha^{5}\lambda^{2}-567\lambda^{4}-1404\alpha\lambda^{4}-882\alpha^{2}\lambda^{4}-204\alpha^{3}\lambda^{4}-15\alpha^{4}\lambda^{4}-180\alpha\lambda^{6}-72\alpha^{2}\lambda^{6}-4\alpha^{3}\lambda^{6}+4\alpha^{2}\lambda^{8}$ and $R=-9\lambda^{2}+6\alpha\lambda^{2}+3\alpha^{2}\lambda^{2}+2\alpha\lambda^{4}$.
III.1 Potential $V(\phi)=V_{0}e^{-\kappa\lambda\phi}$
In this subsection we focus on the analysis of the case of the exponential potential (12) and the two interaction models, one corresponding to the global interaction (10), and one corresponding to the local interaction (11).
III.1.1 Model I: $Q=\alpha H\rho_{m}$
In order to transform the cosmological equations into an autonomous form we introduce the dimensionless variables
$x=\frac{\kappa\dot{\phi}}{\sqrt{6}H},~{}~{}y=\frac{\kappa\sqrt{V}}{\sqrt{3}H}.$ (19)
Using these variables we can express the density parameters as
$\Omega_{\phi}=-x^{2}+y^{2},~{}~{}~{}~{}~{}\Omega_{m}=1+x^{2}-y^{2},$ (20)
while for the total equation-of-state parameter we have
$w_{tot}=-x^{2}-y^{2}.$ (21)
Finally, it proves convenient to introduce the deceleration parameter $q\equiv-1-\frac{\dot{H}}{H^{2}}$, which in the case of dust matter becomes simply $q=\frac{3w_{tot}+1}{2}$, and thus in terms of the dimensionless variables:
$q=\frac{1}{2}\left[1-3(x^{2}+y^{2})\right].$ (22)
In terms of the dimensionless variables (19), the autonomous system takes the form:
$x^{\prime}$ $=$ $-3x-\frac{\sqrt{6}}{2}\lambda y^{2}+\frac{3}{2}x\left(1-x^{2}-y^{2}\right)$ (23)
$-\frac{\alpha}{2x}\left(1+x^{2}-y^{2}\right),$
$y^{\prime}$ $=$ $-\frac{\sqrt{6}}{2}\lambda xy+\frac{3}{2}y\left(1-x^{2}-y^{2}\right),$ (24)
where primes denote derivatives with respect to $N\equiv\ln a$. The above system is invariant under the transformation $y\longrightarrow-y$, therefore the phase space is defined as
$\mathbb{D}=\left\{(x,y)\in\mathbb{R}^{2}:0\leq-x^{2}+y^{2}\leq 1,y\geq 0\right\}.$ (25)
In order to extract the critical points, we impose $x^{\prime}=0$ and $y^{\prime}=0$. The critical points corresponding to an expanding Universe (namely having $y>0$), their existence conditions, their eigenvalues and their stability conditions are given in Table 1, alongside the corresponding values of the physically interesting quantities $\Omega_{\phi}$ and $w_{\rm tot}$. As we observe, for this scenario we have the following critical points.
Figure 1: The stability region of critical point $A_{2}$, for the case of scalar-field potential $V(\phi)=V_{0}e^{-\kappa\lambda\phi}$ and interaction rate $Q=\alpha H\rho_{m}$.
  • The critical point $A_{1}$ exists for all values of the parameters $\alpha$ and $\lambda$. It represents the dark-energy dominated ($\Omega_{\phi}=1$) solution with total equation of state $-1-\frac{\lambda^{2}}{3}$, and thus it corresponds to a phantom accelerating universe. It is a stable node when the condition $-\alpha<3+\lambda^{2}$ is satisfied, and as a late-time attractor this point can describe the late-time universe.
  • The critical point $A_{2}$ exists in the parameter region shown in Table 1. Since the eigenvalues of the corresponding linearized matrix are complicated, the stability conditions have to be determined numerically, and are presented in Fig. 1. When $\alpha<-1$, it corresponds to an accelerating universe. Moreover, at $A_{2}$ we have $\Omega_{m}/\Omega_{\phi}\neq 0$, and thus this point represents late-time accelerating scaling solution. Hence, in this case, the coincidence problem can be alleviated, as reported earlier in Chen et al. (2009). Additionally, we mention that the critical points $A_{2}$ can represent purely matter dominated solutions (namely having $\Omega_{m}=1$) for the parameters $\lambda$ and $\alpha$ satisfying $\lambda^{2}=-(\alpha+3)^{2}/\alpha$. Thus, depending on the parameter values, point $A_{2}$ can correspond to the above possibilities. As we shall see below, for different potential choices such late-time stable accelerating scaling solutions can be obtained in a larger parameter region compared to this case.
Point $x$ $y$ $z$ Existence $E_{1}$ $E_{2}$ $E_{3}$ Stability $\Omega_{\phi}$ $w_{tot}$
$B_{1}$ $-\frac{\lambda}{\sqrt{6}}$ $\sqrt{1+\frac{\lambda^{2}}{6}}$ 0 $\forall$ $\gamma$, $\lambda$ $-(3+\lambda^{2})$ $-(\frac{\lambda^{2}}{2}+3)$ $-\frac{\lambda^{2}}{2}$ always stable node 1 $-1-\frac{\lambda^{2}}{3}$
$B_{2}$ $-\frac{\lambda}{\sqrt{6}}$ $\sqrt{1+\frac{\lambda^{2}}{6}}$ 1 $\forall$ $\gamma$, $\lambda$ $-sgn(\gamma)\infty$ $-(\frac{\lambda^{2}}{2}+3)$ $\frac{\lambda^{2}}{2}$ always saddle 1 $-1-\frac{\lambda^{2}}{3}$
Table 2: The critical points for the case of scalar-field potential $V(\phi)=V_{0}e^{-\kappa\lambda\phi}$ and interaction rate $Q=\Gamma\rho_{m}$, alongside their eigenvalues and their existence and stability conditions. Additionally, we present the corresponding values of the dark-energy density parameter $\Omega_{\phi}$ and the total equation-of-state parameter $w_{tot}$.
III.1.2 Model II: $Q=\Gamma\rho_{m}$
In this case, apart from the dimensionless variables (19) we need one more, namely
$z=\frac{H_{0}}{H+H_{0}},$ (26)
with $H_{0}$ the value of the Hubble function at present. Hence, the autonomous system becomes:
$x^{\prime}$ $=$ $-3x-\frac{\sqrt{6}}{2}\lambda y^{2}+\frac{3}{2}x\left(1-x^{2}-y^{2}\right)$ (27)
$-\frac{\gamma z\left(1+x^{2}-y^{2}\right)}{2x(1-z)},$
$y^{\prime}$ $=$ $-\frac{\sqrt{6}}{2}\lambda xy+\frac{3}{2}y\left(1-x^{2}-y^{2}\right),$ (28)
$z^{\prime}$ $=$ $\frac{3}{2}z(1-z)\left(1-x^{2}-y^{2}\right),$ (29)
where $\gamma=\Gamma/H_{0}$ is the dimensionless coupling parameter. It is clear that the autonomous system (27)-(29) is invariant under the transformation $y\longrightarrow-y$, and therefore the physical region is defined as
$\mathbb{D}=\bigg{\{}(x,y,z)\in\mathbb{R}^{3}:0\leq-x^{2}+y^{2}\leq 1,y\geq 0,0\leq z\leq 1\bigg{\}}.$ (30)
The critical points, their existence conditions, their eigenvalues and their stability conditions are given in Table 2, alongside the values of $\Omega_{\phi}$ and $w_{tot}$. As we observe, for this scenario we have the following critical points.
  • The critical point $B_{1}$ exists always and corresponds to dark-energy dominated solutions. Since the total equation-of-state parameter is always less than $-1$, it corresponds to a phantom universe. Since, all the eigenvalues evaluated at $B_{1}$ are negative, it is always stable node. Thus, this point depicts a dark-energy dominated accelerated solution. However, note that it corresponds to $H\longrightarrow\infty$, as expected for a phantom, super-accelerating (i.e. with $\dot{H}>0$) universe.
  • As before, point $B_{2}$ corresponds to a dark-energy dominated solution, and exists always. It is always saddle and it has $H\longrightarrow 0$. Note that since it differs from point $B_{1}$ only in the $z$ value, both $B_{1}$ and $B_{2}$ have the same $\Omega_{\phi}$ and $w_{tot}$.
Finally, we mention that this local interaction model was investigated in Chen et al. (2009), considering the exponential potential, and the absence of accelerating scaling attractors was also reported.
Point $x$ $y$ $\lambda$ Existence $E_{1}$ $E_{2}$ $E_{3}$ Stability $\Omega_{\phi}$ $w_{tot}$
$C_{1}$ $-\frac{\eta}{\sqrt{6}}$ $\sqrt{1+\frac{\eta^{2}}{6}}$ $\eta$ $\forall$ $\alpha$, $\eta$ $-(3+\alpha+\eta^{2})$ $-(3+\frac{\eta^{2}}{2})$ $-2\eta^{2}$ stable node for $-\alpha<3+\eta^{2}$ 1 $-1-\frac{\eta^{2}}{3}$
for $-3<\alpha<0$
$C_{2}$ $\frac{3+\alpha}{\sqrt{6}\eta}$   $\sqrt{-\frac{\alpha}{3}-\frac{(3+\alpha)^{2}}{6\eta^{2}}}$   $\eta$   $\eta^{2}\geq\frac{(3+\alpha)^{2}}{-\alpha}$ $\frac{P+\sqrt{S}}{4(3+\alpha)\eta^{2}}$ $\frac{P-\sqrt{S}}{4(3+\alpha)\eta^{2}}$ $2(3+\alpha)$ stable node for $\alpha<-3$ $-\frac{\alpha\eta^{2}+(\alpha+3)^{2}}{3\eta^{2}}$ $\frac{\alpha}{3}$
and for $\alpha<-3$ $-(\alpha+3)>\eta^{2}\geq\frac{(\alpha+3)^{2}}{-\alpha}$
$-(\alpha+3)\geq\eta^{2}\geq\frac{(\alpha+3)^{2}}{-\alpha}$ (see Fig. 2)
$C_{3}$ $\frac{\eta}{\sqrt{6}}$ $\sqrt{1+\frac{\eta^{2}}{6}}$ $-\eta$ $\forall$ $\alpha$, $\eta$ $-(3+\alpha+\eta^{2})$ $-(3+\frac{\eta^{2}}{2})$ $-2\eta^{2}$ stable node for $-\alpha<3+\eta^{2}$ 1 $-1-\frac{\eta^{2}}{3}$
for $-3<\alpha<0$
$C_{4}$ $-\frac{3+\alpha}{\sqrt{6}\eta}$ $\sqrt{-\frac{\alpha}{3}-\frac{(3+\alpha)^{2}}{6\eta^{2}}}$ $-\eta$ $\eta^{2}\geq\frac{(3+\alpha)^{2}}{-\alpha}$ $\frac{P+\sqrt{S}}{4(3+\alpha)\eta^{2}}$ $\frac{P-\sqrt{S}}{4(3+\alpha)\eta^{2}}$ $2(3+\alpha)$ stable node for $\alpha<-3$ $-\frac{\alpha\eta^{2}+(\alpha+3)^{2}}{3\eta^{2}}$ $\frac{\alpha}{3}$
and for $\alpha<-3$ $-(\alpha+3)>\eta^{2}\geq\frac{(\alpha+3)^{2}}{-\alpha}$
$-(\alpha+3)\geq\eta^{2}\geq\frac{(\alpha+3)^{2}}{-\alpha}$ (see Fig. 2)
Table 3: The critical points for the case of $V(\phi)=V_{0}\cosh{(\kappa\eta\phi)}$ and interaction rate $Q=\alpha H\rho_{m}$, alongside their existence, their eigenvalues and stability conditions. Additionally, we present the corresponding values of the dark-energy density parameter $\Omega_{\phi}$ and the total equation-of-state parameter $w_{tot}$. We have defined $S=-1944\eta^{2}-3240\alpha\eta^{2}-2160\alpha^{2}\eta^{2}-720\alpha^{3}\eta^{2}-120\alpha^{4}\eta^{2}-8\alpha^{5}\eta^{2}-567\eta^{4}-1404\alpha\eta^{4}-882\alpha^{2}\eta^{4}-204\alpha^{3}\eta^{4}-15\alpha^{4}\eta^{4}-180\alpha\eta^{6}-72\alpha^{2}\eta^{6}-4\alpha^{3}\eta^{6}+4\alpha^{2}\eta^{8}$ and $P=-9\eta^{2}+6\alpha\eta^{2}+3\alpha^{2}\eta^{2}+2\alpha\eta^{4}$.
III.2 Potential: $V(\phi)=V_{0}\cosh{(\kappa\eta\phi)}$
In this subsection we focus on the phase space analysis of the autonomous system for the hyperbolic potential (13) considering the global interaction rate (10) and the local interaction rate (11).
III.2.1 Model I: $Q=\alpha H\rho_{m}$
In order to transform the cosmological equations into an autonomous form, apart from the dimensionless variables (19) we need one more, namely
$\lambda=-\frac{V_{\phi}}{\kappa V},$ (31)
where a $\phi$-subscript denotes derivative with respect to $\phi$. If $\lambda(\phi)$ is invertible, and one can extract the form $\phi=\phi(\lambda)$, then the extraction of an autonomous system can be obtained by simply calculating the parameter $\Xi=\frac{VV_{\phi\phi}}{{V_{\phi}}^{2}}$. Using the potential $V(\phi)=V_{0}\cosh{(\kappa\eta\phi)}$, we find $\lambda=-\eta\tanh{(\kappa\eta\phi)}$ and $\Xi=\frac{\eta^{2}}{\lambda^{2}}$ Therefore, the range of $\lambda$ is $[-|\eta|,|\eta|]$. Finally, the autonomous system takes the form:
$x^{\prime}$ $=$ $-3x-\frac{\sqrt{6}}{2}\lambda y^{2}+\frac{3}{2}x\left(1-x^{2}-y^{2}\right)$ (32)
$-\frac{\alpha}{2x}\left(1+x^{2}-y^{2}\right),$
$y^{\prime}$ $=$ $-\frac{\sqrt{6}}{2}\lambda xy+\frac{3}{2}y\left(1-x^{2}-y^{2}\right),$ (33)
$\lambda^{\prime}$ $=$ $\sqrt{6}\left(\lambda^{2}-\eta^{2}\right)x.$ (34)
Since the above autonomous system is invariant under the transformation $y\longrightarrow-y$, the phase space domain $\mathbb{D}$ is defined as $\mathbb{D}=\left\{(x,y,\lambda)\in\mathbb{R}^{3}:0\leq-x^{2}+y^{2}\leq 1,y\geq 0,-|\eta|\leq\lambda\leq|\eta|\right\}$.
The critical points, their existence conditions, their eigenvalues and their stability conditions are given in Table 3, alongside the corresponding values of $\Omega_{\phi}$ and $w_{tot}$. For this scenario we have the following critical points.
  • The critical point $C_{1}$ exists for all $\alpha$ and $\eta$ and it corresponds to an accelerating universe, and it is stable in the region $-\alpha<3+\eta^{2}$. Moreover, at this point we have $\Omega_{\phi}=1$, which implies complete dark energy domination. Therefore, it describes the late-time dark-energy dominated accelerated universe.
  • The critical point $C_{2}$ describes an accelerating universe, since in their stability region, i.e. for $\alpha<-1$, we have $w_{tot}<-1/3$. The stability region is graphically shown in Fig. 2. As we observe, in this case we find almost similar stability region as in Fig. 1 for the exponential potential, which was expected since the present hyperbolic cosinus potential is related to the exponential potential.
    Again, we notice that at $C_{2}$ the ratio of $\Omega_{m}$ and $\Omega_{\phi}$ is non zero (namely $\Omega_{m}/\Omega_{\phi}\neq 0$), and thus this point represents a late-time accelerating scaling solution, which is stable in the aforementioned specific parameter region. Thus, for this case, the coincidence problem can be alleviated. Finally, one can further notice that $C_{2}$ can represent completely matter dominated solutions for a specific parameter space satisfying $\eta^{2}=-(\alpha+3)^{2}/\alpha$.
  • The critical point $C_{3}$ is stable for $-\alpha<3+\eta^{2}$, and the values of $\Omega_{\phi}$ and $w_{\rm tot}$ are $1$ and $-1-\frac{\eta^{2}}{3}<-1$ respectively, therefore this point represents dark-energy dominated accelerated solutions. Hence, point $C_{3}$, similar to $C_{1}$, is a candidate for the late-time universe.
  • The critical point $C_{4}$ has $w_{\rm tot}=\frac{\alpha}{3}<-\frac{1}{3}$ for $\alpha<-1$, and thus it corresponds to accelerated solutions. The stability region is graphically shown in Fig. 2, which is the same with point $C_{2}$ since they have the same eigenvalues. At this critical point $\Omega_{m}/\Omega_{\phi}\neq 0$ and therefore it corresponds to late-time accelerating stable solutions that can alleviate the coincidence problem. Finally, completely matter-dominated solutions are obtained for $\eta^{2}=-(\alpha+3)^{2}/\alpha$.
Figure 2: The stability region of the critical points $C_{2}$ and $C_{4}$, for the case of scalar-field potential $V(\phi)=V_{0}\cosh{(\kappa\eta\phi)}$ and interaction rate $Q=\alpha H\rho_{m}$.
III.2.2 Model II: $Q=\Gamma\rho_{m}$
Point $x$ $y$ $\lambda$ $z$ Existence $E_{1}$ $E_{2}$ $E_{3}$ $E_{4}$ Stability $\Omega_{\phi}$ $w_{tot}$
$D_{1}$ $-\frac{\eta}{\sqrt{6}}$   $\sqrt{1+\frac{\eta^{2}}{6}}$   $\eta$   0   $\forall$ $\gamma$, $\eta$ $-(3+\eta^{2})$ $-(\frac{\eta^{2}}{2}+3)$ $-2\eta^{2}$ $-\frac{\eta^{2}}{2}$ stable node 1 $-1-\frac{\eta^{2}}{3}$
$D_{2}$ $-\frac{\eta}{\sqrt{6}}$   $\sqrt{1+\frac{\eta^{2}}{6}}$   $\eta$   1   $\forall$ $\gamma$, $\eta$ $-sgn(\gamma)\infty$ $-(\frac{\eta^{2}}{2}+3)$ $-2\eta^{2}$ $\frac{\eta^{2}}{2}$ always saddle 1 $-1-\frac{\eta^{2}}{3}$
$D_{3}$ $\frac{\eta}{\sqrt{6}}$   $\sqrt{1+\frac{\eta^{2}}{6}}$   $-\eta$   0   $\forall$ $\gamma$, $\eta$ $-(3+\eta^{2})$ $-(\frac{\eta^{2}}{2}+3)$ $-2\eta^{2}$ $-\frac{\eta^{2}}{2}$ stable node 1 $-1-\frac{\eta^{2}}{3}$
$D_{4}$ $\frac{\eta}{\sqrt{6}}$   $\sqrt{1+\frac{\eta^{2}}{6}}$   $-\eta$   1   $\forall$ $\gamma$, $\eta$ $-sgn(\gamma)\infty$ $-(\frac{\eta^{2}}{2}+3)$ $-2\eta^{2}$ $\frac{\eta^{2}}{2}$ always saddle 1 $-1-\frac{\eta^{2}}{3}$
Table 4: The critical points for the case of $V(\phi)=V_{0}\cosh{(\kappa\eta\phi)}$ and interaction rate $Q=\Gamma\rho_{m}$, alongside their eigenvalues and their existence and stability conditions. Additionally, we present the corresponding values of the dark-energy density parameter $\Omega_{\phi}$ and the total equation-of-state parameter $w_{tot}$.
Using the dimensionless variables $x$, $y$ defined in (19) alongside the dimensionless variables $z$ and $\lambda$ of (26) and (31), we obtain the autonomous system as
$x^{\prime}$ $=$ $-3x-\frac{\sqrt{6}}{2}\lambda y^{2}+\frac{3}{2}x\left(1-x^{2}-y^{2}\right)$ (35)
$-\frac{\gamma z\left(1+x^{2}-y^{2}\right)}{2x(1-z)},$
$y^{\prime}$ $=$ $-\frac{\sqrt{6}}{2}\lambda xy+\frac{3}{2}y\left(1-x^{2}-y^{2}\right),$ (36)
$\lambda^{\prime}$ $=$ $\sqrt{6}\left(\lambda^{2}-\eta^{2}\right)x,$ (37)
$z^{\prime}$ $=$ $\frac{3}{2}z(1-z)\left(1-x^{2}-y^{2}\right),$ (38)
where $\gamma=\Gamma/H_{0}$ is the dimensionless coupling parameter. The above autonomous system is invariant under the transformation $y\longrightarrow-y$, and the phase space $\mathbb{D}$ is given by
$\mathbb{D}=\bigg{\{}(x,y,\lambda,z)\in\mathbb{R}^{4}:0\leq-x^{2}+y^{2}\leq 1,y\geq 0,$
$-|\eta|\leq\lambda\leq|\eta|,0\leq z\leq 1\bigg{\}}.$ (39)
The critical points, their existence conditions, their stability conditions, and the values of $\Omega_{\phi}$ and $w_{tot}$, are given in Table 4. We have the following critical points.
  • The critical point $D_{1}$ exists always and it corresponds to super-accelerated dark-energy domination, exhibiting $H\longrightarrow\infty$. It is always stable and it can describe the Universe at late times.
  • The critical point $D_{2}$ exists always, corresponding to dark energy dominated accelerated solutions, with $H\longrightarrow 0$, and it is always saddle.
  • Point $D_{3}$ exists always, and it corresponds to dark-energy dominated accelerated solutions, with $H\longrightarrow\infty$. Since it is stable, it is also a suitable candidate for late-time evolution of the universe.
  • Point $D_{4}$ exists always, it is saddle, and it corresponds to dark-energy dominated accelerated solutions with $H\longrightarrow 0$.
In summary, the hyperbolic potential with the local interaction rate leads to results similar to the exponential potential, since the two potentials are related. Thus, in this scenario, late-time scaling attractors are not possible.
III.3 Potential: $V(\phi)=\frac{1}{8\beta}\left(1-e^{-\kappa\mu\;\phi}\right)^{2}$
In this subsection we focus on the phase space analysis of the autonomous system for the potential (14), considering as usual both the global interaction rate (10) and the local interaction rate (11).
III.3.1 Model I: $Q=\alpha H\rho_{m}$
Point $x$ $y$ $\lambda$ Existence $E_{1}$ $E_{2}$ $E_{3}$ Stability $\Omega_{\phi}$ $w_{tot}$
$F_{1}$ $-\frac{2\mu}{\sqrt{6}}$ $\sqrt{1+\frac{2\mu^{2}}{3}}$ $2\mu$ $\forall$ $\alpha$, $\mu$ $-(3+\alpha+4\mu^{2})$ $-3-2\mu^{2}$ $-2\mu^{2}$ stable node for $-\alpha<3+4\mu^{2}$ 1 $-1-\frac{4\mu^{2}}{3}$
for $-3<\alpha<0$
$F_{2}$ $\frac{3+\alpha}{2\sqrt{6}\mu}$ $\sqrt{-\frac{\alpha}{3}-\frac{(3+\alpha)^{2}}{24\mu^{2}}}$ $2\mu$ $\mu^{2}\geq\frac{(3+\alpha)^{2}}{-4\alpha}$ $\frac{P+\sqrt{S}}{4(3+\alpha)\mu^{2}}$ $\frac{P-\sqrt{S}}{4(3+\alpha)\mu^{2}}$ $\frac{1}{2}(3+\alpha)$ stable node for $\alpha<-3$ $-\frac{4\alpha\mu^{2}+(\alpha+3)^{2}}{12\mu^{2}}$ $\frac{\alpha}{3}$
and for $\alpha<-3$ $-\frac{(\alpha+3)}{4}>\mu^{2}\geq\frac{(\alpha+3)^{2}}{-4\alpha}$
$-\frac{(\alpha+3)}{4}\geq\mu^{2}\geq\frac{(\alpha+3)^{2}}{-4\alpha}$ (see Fig. 3)
$F_{3}$ $x_{c}$ $\sqrt{1-x_{c}^{2}}$ 00 $\forall\mu$, $\alpha=-3$, 00 $-6$ $-\sqrt{6}\mu x_{c}$ stable for $\mu>0$ and $x_{c}>0$ $1-2x_{c}^{2}$ $-1$
$|x_{c}|\leq\frac{1}{\sqrt{2}}$ with $x_{c}\neq 0$ or for $\mu<0$ and $x_{c}<0$
Table 5: The critical points for the case of $V(\phi)=\frac{1}{8\beta}\left(1-e^{-\kappa\mu\;\phi}\right)^{2}$ and interaction rate $Q=\alpha H\rho_{m}$, alongside their existence, eigenvalues and stability conditions. Furthermore, we present the corresponding values of the dark-energy density parameter $\Omega_{\phi}$ and the total equation-of-state parameter $w_{tot}$. We have defined $S=-486\mu^{2}-810\alpha\mu^{2}-540\alpha^{2}\mu^{2}-180\alpha^{3}\mu^{2}-30\alpha^{4}\mu^{2}-2\alpha^{5}\mu^{2}-567\mu^{4}-1404\alpha\mu^{4}-882\alpha^{2}\mu^{4}-204\alpha^{3}\mu^{4}-15\alpha^{4}\mu^{4}-720\alpha\mu^{6}-288\alpha^{2}\mu^{6}-16\alpha^{3}\mu^{6}+24\alpha^{2}\mu^{8}$ and $P=-9\mu^{2}+6\alpha\mu^{2}+3\alpha^{2}\mu^{2}+8\alpha\mu^{4}$.
In this case we derive the autonomous system in terms of the dimensionless variables $x$, $y$, $\lambda$, defined in (19) and (31). As we can see, the expressions of the parameters $\lambda=-\frac{V_{\phi}}{\kappa V}$ and $\Xi=\frac{VV_{\phi\phi}}{V_{\phi}^{2}}$ take the form $\lambda=-2\mu\frac{e^{-\kappa\mu\phi}}{1-e^{-\kappa\mu\phi}}$ ($\mu\neq 0$) and $\Xi=\frac{\mu}{\lambda}+\frac{1}{2}$, respectively. Therefore, the autonomous system becomes:
$x^{\prime}$ $=$ $-3x-\frac{\sqrt{6}}{2}\lambda y^{2}+\frac{3}{2}x\left(1-x^{2}-y^{2}\right)$ (40)
$-\frac{\alpha}{2x}\left(1+x^{2}-y^{2}\right),$
$y^{\prime}$ $=$ $-\frac{\sqrt{6}}{2}\lambda xy+\frac{3}{2}y\left(1-x^{2}-y^{2}\right),$ (41)
$\lambda^{\prime}$ $=$ $\frac{\sqrt{6}}{2}\left(\lambda^{2}-2\mu\lambda\right)x.$ (42)
The above system is invariant under $y\longrightarrow-y$, and its phase space region $\mathbb{D}$ is defined by $\mathbb{D}=\left\{(x,y,\lambda)\in\mathbb{R}^{3}:0\leq-x^{2}+y^{2}\leq 1,y\geq 0\right\}$.
The critical points, their existence conditions, their eigenvalues and their stability conditions are given in Table 5, alongside the corresponding values of $\Omega_{\phi}$ and $w_{tot}$. We have the following critical points.
Figure 3: The stability region of the critical point $F_{2}$, for the case of scalar-field potential $V(\phi)=\frac{1}{8\beta}\left(1-e^{-\kappa\mu\;\phi}\right)^{2}$ and interaction rate $Q=\alpha H\rho_{m}$.
Point $x$ $y$ $\lambda$ $z$ Existence $E_{1}$ $E_{2}$ $E_{3}$ $E_{4}$ Stability $\Omega_{\phi}$ $w_{tot}$
$G_{1}$ $-\frac{2\mu}{\sqrt{6}}$ $\sqrt{1+\frac{2\mu^{2}}{3}}$ $2\mu$ 0 $\forall$ $\gamma$, $\mu$ $-3-4\mu^{2}$ $-2\mu^{2}-3$ $-2\mu^{2}$ $-2\mu^{2}$ stable node 1 $-1-\frac{4\mu^{2}}{3}$
$G_{2}$ $-\frac{2\mu}{\sqrt{6}}$ $\sqrt{1+\frac{2\mu^{2}}{3}}$ $2\mu$ 1 $\forall$ $\gamma$, $\mu$ $-sgn(\gamma)\infty$ $-2\mu^{2}-3$ $-2\mu^{2}$ $2\mu^{2}$ always saddle 1 $-1-\frac{4\mu^{2}}{3}$
$G_{3}$ $x_{c}$ $\sqrt{1-{x_{c}}^{2}}$ 0 $\frac{3}{3-\gamma}$ $|x_{c}|\leq\frac{1}{\sqrt{2}}$, $x_{c}\neq 0$ 00 $-\sqrt{6}\mu x_{c}$ $-3\left(1-\sqrt{1-{x_{c}}^{2}}\right)$ $-3\left(1+\sqrt{1-{x_{c}}^{2}}\right)$ stable $1-2{x_{c}}^{2}$ $-1$
with $\gamma<0$ for $\mu x_{c}>0$
Table 6: The critical points for the case of $V(\phi)=\frac{1}{8\beta}\left(1-e^{-\kappa\mu\;\phi}\right)^{2}$ and interaction rate $Q=\Gamma\rho_{m}$, alongside their eigenvalues and their existence and stability conditions. Additionally, we present the corresponding values of the dark-energy density parameter $\Omega_{\phi}$ and the total equation-of-state parameter $w_{tot}$.
  • Point $F_{1}$ exists always, and it corresponds to a dark energy dominated accelerating solution. Moreover, it is a stable node for $-\alpha<3+4\mu^{2}$ and hence it can describe the late-time universe.
  • The critical point $F_{2}$ corresponds to accelerating solutions for $\alpha<-1$, where $w_{tot}<-\frac{1}{3}$. We mention that at this point we have $\Omega_{m}/\Omega_{\phi}\neq 0$ corresponding to accelerating scaling solution, which is also stable in its stability region as shown in Table 5 and in Fig. 3, and thus it can alleviate the coincidence problem. Finally, one can see that for $\mu^{2}=-(\alpha+3)^{2}/4\alpha$ point $F_{1}$ corresponds to completely matter dominated solutions.
  • The curve of critical points $F_{3}$ has $\lambda=0$, and it was absent for the exponential and hyperbolic potentials. In general these points correspond to scaling accelerating solutions, while for $x_{c}=\pm\frac{1}{\sqrt{2}}$ they correspond to completely matter-dominated points. Since, one eigenvalue is zero and the other two are negative, they are normally hyperbolic critical points. By applying the center manifold theorem Leon and Saridakis (2010) we deduce that these critical points behave as stable points for $\mu x_{c}>0$. Thus, these points can alleviate the coincidence problem for the model parameter $\alpha=-3$.
III.3.2 Model II: $Q=\Gamma\rho_{m}$
In this case, using the dimensionless variables $x$, $y$, $\lambda$, $z$ of (19), (26) and (31), we result to the autonomous system
$x^{\prime}$ $=$ $-3x-\frac{\sqrt{6}}{2}\lambda y^{2}+\frac{3}{2}x\left(1-x^{2}-y^{2}\right)$ (43)
$-\frac{\gamma z\left(1+x^{2}-y^{2}\right)}{2x(1-z)},$
$y^{\prime}$ $=$ $-\frac{\sqrt{6}}{2}\lambda xy+\frac{3}{2}y\left(1-x^{2}-y^{2}\right),$ (44)
$\lambda^{\prime}$ $=$ $\frac{\sqrt{6}}{2}\left(\lambda^{2}-2\mu\lambda\right)x,$ (45)
$z^{\prime}$ $=$ $\frac{3}{2}z(1-z)\left(1-x^{2}-y^{2}\right),$ (46)
where $\gamma=\Gamma/H_{0}$. The above autonomous system is invariant under $y\longrightarrow-y$, and the phase space domain is defined as $\mathbb{D}=\left\{(x,y,\lambda,z)\in\mathbb{R}^{4}:0\leq-x^{2}+y^{2}\leq 1,y\geq 0,0\leq z\leq 1\right\}$. The properties and features of the critical points are summarized in Table 6. They are the following.
Point $x$ $v$ $w$ $\xi$ Existence Stability $\Omega_{\phi}$ $w_{tot}$
$I_{1}$ $\frac{2}{3}$ 0 0 $\frac{\sqrt{13}}{3}$ $\forall\alpha$ stable for $\alpha>-\frac{17}{3}$ 1 $-\frac{17}{9}$
[…]
$\cup\left(-3,\frac{-13+\sqrt{88}}{3}\right]$ otherwise saddle or unstable
$I_{3}$ $\frac{1}{3}$ 0 $\frac{\sqrt{10}}{3}$ 0 $\forall\alpha$ always saddle 1 $-\frac{11}{9}$
[…]
$\cup\left(-3,\frac{-10+\sqrt{19}}{3}\right]$
Table 7: The critical points for the case of $V(\phi)=c_{0}+c_{1}e^{\sqrt{\frac{2\kappa^{2}}{3}}\phi}+c_{2}e^{2\sqrt{\frac{2\kappa^{2}}{3}}\phi}$ and interaction rate $Q=\alpha H\rho_{m}$, alongside their existence and stability conditions. Furthermore, we present the corresponding values of the dark-energy density parameter $\Omega_{\phi}$ and the total equation-of-state parameter $w_{tot}$.
  • The critical point $G_{1}$ exists always, corresponding to dark-energy dominated super-accelerating solutions, with $H\longrightarrow\infty$, and it is always stable node. Thus, $G_{1}$ is a candidate for describing the late-time accelerated universe.
  • Point $G_{2}$ exists always, corresponding to dark-energy dominated accelerating solutions, with $H\longrightarrow 0$, and it is a saddle point.
  • The curves of critical points $G_{3}$ exist for $|x_{c}|\leq\frac{1}{\sqrt{2}}$ with $\gamma<0$. They are normally hyperbolic, and by applying the center manifold theorem in order to determine their stability Leon and Saridakis (2010) we deduce that $G_{3}$ are stable for $\mu x_{c}>0$. Note that in general, for $|x_{c}|<\frac{1}{\sqrt{2}}$, these points have $10<\Omega_{\phi}<$ 1, and thus they correspond to a late-time stable accelerating solution with $\Omega_{m}/\Omega_{\phi}\neq 0$, namely with a scaling nature that can alleviate the coincidence problem. Finally, in the limiting case where $x_{c}=\pm\frac{1}{\sqrt{2}}$ they lead to $\Omega_{\phi}=0$ and thus to $\Omega_{m}=1$, corresponding to matter-dominated solutions.
In summary, under the local interaction rate the potential $V(\phi)=\frac{1}{8\beta}\left(1-e^{-\kappa\mu\;\phi}\right)^{2}$ is able to accept late-time stable scaling accelerating solutions, which are not possible for the previous potentials, and that is why they were not studied in the literature. This is a novel result of the present work.
III.4 Potential: $V(\phi)=c_{0}+c_{1}e^{\sqrt{\frac{2\kappa^{2}}{3}}\phi}+c_{2}e^{2\sqrt{\frac{2\kappa^{2}}{3}}\phi}$
In this subsection we focus on the last potential function (15). In order to obtain the autonomous form for this potential case we introduce three new dimensionless variables $v$, $w$, $\xi$ as
$v=\frac{\kappa\sqrt{c_{0}}}{\sqrt{3}H},~{}w=\frac{\kappa\sqrt{c_{1}e^{\sqrt{\frac{2\kappa^{2}}{3}}\phi}}}{\sqrt{3}H},~{}\xi=\frac{\kappa\sqrt{c_{2}e^{2\sqrt{\frac{2\kappa^{2}}{3}}\phi}}}{\sqrt{3}H}.$ (47)
In terms of these variables, as well as $x$ defined in (19), the various cosmological parameters become
$\Omega_{m}=1+x^{2}-v^{2}-w^{2}-\xi^{2},$
$\Omega_{\phi}=-x^{2}+v^{2}+w^{2}+\xi^{2},$
$w_{tot}=-x^{2}-v^{2}-w^{2}-\xi^{2},$
$q=\frac{1}{2}\left[1-3(x^{2}+v^{2}+w^{2}+\xi^{2})\right].$ (48)
III.4.1 Model I: $Q=\alpha H\rho_{m}$
Using the above dimensionless variables we obtain the autonomous system as
$x^{\prime}$ $=$ $-\frac{3}{2}x\left(1+x^{2}+v^{2}+w^{2}+\xi^{2}\right)+w^{2}+2\xi^{2}$ (49)
$-\frac{\alpha\left(1+x^{2}-v^{2}-w^{2}-\xi^{2}\right)}{2x},$
$v^{\prime}$ $=$ $\frac{3}{2}v\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right),$ (50)
$w^{\prime}$ $=$ $w\left[x+\frac{3}{2}\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right)\right],$ (51)
$\xi^{\prime}$ $=$ $\xi\left[2x+\frac{3}{2}\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right)\right].$ (52)
This system is invariant under the transformations $v\longrightarrow-v$, $w\longrightarrow-w$ and $\xi\longrightarrow-\xi$ separately, and thus the physical region is defined as
$\mathbb{D}$ $=$ $\bigg{\{}(x,v,w,\xi)\in\mathbb{R}^{4}:0\leq-x^{2}+v^{2}+w^{2}+\xi^{2}\leq 1,$ (53)
$v\geq 0,w\geq 0,\xi\geq 0\bigg{\}}.$
The critical points, the existence and stability conditions as well as the corresponding values of $\Omega_{\phi}$ and $w_{tot}$ are summarized in Table 7, while the corresponding eigenvalues are displayed in Table 8. For this scenario we have the following critical points.
Point $E_{1}$ $E_{2}$ $E_{3}$ $E_{4}$
$I_{1}$ $-\frac{17}{3}-\alpha$ $-\frac{4}{3}$ $-\frac{2}{3}$ $-\frac{13}{3}$
$I_{2}$ $\frac{3+\alpha}{4}$ $\frac{3+\alpha}{2}$ $\frac{T^{\prime}-\sqrt{V^{\prime}}}{12(3+\alpha)}$ $\frac{T^{\prime}+\sqrt{V^{\prime}}}{12(3+\alpha)}$
$I_{3}$ $-\frac{11}{3}-\alpha$ $-\frac{1}{3}$ $-\frac{10}{3}$ $\frac{1}{3}$
$I_{4}$ $\frac{3+\alpha}{2}$ $-\frac{3+\alpha}{2}$ $\frac{R^{\prime}-\sqrt{S^{\prime}}}{12(3+\alpha)}$ $\frac{R^{\prime}+\sqrt{S^{\prime}}}{12(3+\alpha)}$
Table 8: The critical points and their associated eigenvalues of the linearised (Jacobian) matrix, for the case of scalar-field potential $V(\phi)=c_{0}+c_{1}e^{\sqrt{\frac{2\kappa^{2}}{3}}\phi}+c_{2}e^{2\sqrt{\frac{2\kappa^{2}}{3}}\phi}$ and interaction rate $Q=\alpha H\rho_{m}$. We have defined $T^{\prime}=-27+34\alpha+9\alpha^{2}$, $V^{\prime}=-11664-27891\alpha-16700\alpha^{2}-4362\alpha^{3}-540\alpha^{4}-27\alpha^{5}$, $R^{\prime}=-27+22\alpha+9\alpha^{2}$ and $S^{\prime}=-31347-57456\alpha-37514\alpha^{2}-11580\alpha^{3}-1755\alpha^{4}-108\alpha^{5}$.
  • The critical point $I_{1}$ exists for all $\alpha$, and it corresponds to a dark-energy dominated super-accelerated solution. It is stable in the region $\alpha>-\frac{17}{3}$ and thus it can describe the late-time phase of the universe.
  • Point $I_{2}$ in its stability region has $0\leq\Omega_{\phi}\leq 1$, and thus $0\leq\Omega_{m}\leq 1$. Hence, since $\Omega_{\phi}$ and $\Omega_{m}$ can be of the same order, this point can alleviate the coincidence problem Amendola (2000); Chimento et al. (2003); Cai and Wang (2005); Pavon and Zimdahl (2005); Huey and Wandelt (2006); Hu and Ling (2006); del Campo et al. (2008, 2009). Moreover, for $\alpha=\frac{-13-\sqrt{88}}{3}$ or $\alpha=\frac{-33+\sqrt{88}}{3}$, it has $\Omega_{m}=1$, and in this case it corresponds to a matter-dominated solution.
  • Point $I_{3}$ exists always, it is saddle, and it corresponds to a dark-energy dominated super-accelerating solution.
    Point $x$ $v$ $w$ $\xi$ $z$ Existence Stability $\Omega_{\phi}$ $w_{tot}$
    $J_{1}$ $\frac{1}{3}$ 0 $\frac{\sqrt{10}}{3}$ 0 0 $\forall\gamma$ always saddle 1 $-\frac{11}{9}$
    $J_{2}$ $\frac{2}{3}$ 0 0 $\frac{\sqrt{13}}{3}$ 0 $\forall\gamma$ always stable 1 $-\frac{17}{9}$
    $J_{3}$ $\frac{1}{3}$ 0 $\frac{\sqrt{10}}{3}$ 0 1 $\forall\gamma$ always saddle 1 $-\frac{11}{9}$
    $J_{4}$ $\frac{2}{3}$ 0 0 $\frac{\sqrt{13}}{3}$ 1 $\forall\gamma$ always saddle 1 $-\frac{17}{9}$
    $J_{5}$ $x_{c}$ $\sqrt{1-{x_{c}}^{2}}$ 0 0 $\ \ \ \frac{3}{3-\gamma}$ $\ \ \ 0<|x_{c}|\leq\frac{1}{\sqrt{2}}$ stable for $-\frac{1}{\sqrt{2}}\leq x_{c}<0$ $\ \ \ 1-2{x_{c}}^{2}$ $-1$
    with $\gamma<0$
    Table 9: The critical points for the case of $V(\phi)=c_{0}+c_{1}e^{\sqrt{\frac{2\kappa^{2}}{3}}\phi}+c_{2}e^{2\sqrt{\frac{2\kappa^{2}}{3}}\phi}$ and interaction rate $Q=\Gamma\rho_{m}$, alongside their existence and stability conditions. Furthermore, we present the corresponding values of the dark-energy density parameter $\Omega_{\phi}$ and the total equation-of-state parameter $w_{tot}$.
  • Point $I_{4}$ has $0\leq\Omega_{\phi}\leq 1$ and $0\leq\Omega_{m}\leq 1$, and corresponds to accelerating solutions. However, since it is saddle it cannot offer an alleviation for the coincidence problem. Finally, in the limiting case where $\alpha=\frac{-10\pm\sqrt{19}}{3}$ we have $\Omega_{m}=1$, corresponding to a matter-dominated solution.
III.4.2 Model II: $Q=\Gamma\rho_{m}$
In this case, using the dimensionless variables $v$, $w$, $\xi$ defined in (47), the variable $z$ defined in (26) and the dimensionless variable $x$ from (19), the autonomous system can be written as
$x^{\prime}$ $=$ $-\frac{3}{2}x\left(1+x^{2}+v^{2}+w^{2}+\xi^{2}\right)+w^{2}+2\xi^{2}$ (54)
$-\frac{\gamma\left(1+x^{2}-v^{2}-w^{2}-\xi^{2}\right)z}{2x(1-z)},$
$v^{\prime}$ $=$ $\frac{3}{2}v\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right),$ (55)
$w^{\prime}$ $=$ $w\left[x+\frac{3}{2}\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right)\right],$ (56)
$\xi^{\prime}$ $=$ $\xi\left[2x+\frac{3}{2}\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right)\right],$ (57)
$z^{\prime}$ $=$ $\frac{3}{2}z(1-z)\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right),$ (58)
where $\gamma=\Gamma/H_{0}$ is the dimensionless coupling parameter. The system is invariant under the transformations $v\longrightarrow-v$, $w\longrightarrow-w$ and $\xi\longrightarrow-\xi$ separately, and thus the physical region is defined as
$\mathbb{D}$ $=$ $\bigg{\{}(x,v,w,\xi,z)\in\mathbb{R}^{5}:0\leq-x^{2}+v^{2}+w^{2}+\xi^{2}\leq 1,$ (59)
$v\geq 0,w\geq 0,\xi\geq 0,0\leq z\leq 1\bigg{\}}.$
The critical points, the existence and stability conditions as well as the corresponding values of $\Omega_{\phi}$ and $w_{tot}$ are summarized in Table 9, while the corresponding eigenvalues are displayed in Table 10. For this scenario we have the following critical points.
Point $E_{1}$ $E_{2}$ $E_{3}$ $E_{4}$ $E_{5}$
$J_{1}$ $-\frac{11}{3}$ $-\frac{1}{3}$ $-\frac{10}{3}$ $\frac{1}{3}$ $-\frac{1}{3}$
$J_{2}$ $-\frac{17}{3}$ $-\frac{4}{3}$ $-\frac{2}{3}$ $-\frac{13}{3}$ $-\frac{4}{3}$
$J_{3}$ $-sgn(\gamma)\infty$ $-\frac{1}{3}$ $-\frac{10}{3}$ $\frac{1}{3}$ $\frac{1}{3}$
$J_{4}$ $-sgn(\gamma)\infty$ $-\frac{4}{3}$ $-\frac{2}{3}$ $-\frac{13}{3}$ $\frac{4}{3}$
$J_{5}$ 00 $x_{c}$ $2x_{c}$ $-3(1+\sqrt{1-{x_{c}}^{2}})$ $-3(1-\sqrt{1-{x_{c}}^{2}})$
Table 10: The critical points and their associated eigenvalues of the linearised (Jacobian) matrix, for the case of scalar-field potential $V(\phi)=c_{0}+c_{1}e^{\sqrt{\frac{2\kappa^{2}}{3}}\phi}+c_{2}e^{2\sqrt{\frac{2\kappa^{2}}{3}}\phi}$ and interaction rate $Q=\Gamma\rho_{m}$.
  • The critical point $J_{1}$ does not depend on the parameter $\gamma$, it exists always and it is saddle, corresponding to super-accelerated dark-energy dominated solution, with $H\longrightarrow\infty$.
  • Point $J_{2}$ exists always and it is stable, corresponding to super-accelerated dark-energy dominated solution, with $H\longrightarrow\infty$. Thus, $J_{2}$ can describe the late-time evolution of the universe.
  • Point $J_{3}$ exists always, it is saddle, and it also corresponds to super-accelerated dark-energy dominated solution, but with $H\longrightarrow 0$.
    Figure 4: Summary plot of the phase-space behavior and the critical points for the interaction models $Q=\alpha H\rho_{m}$ and $Q=\Gamma\rho_{m}$, considering all the examined potentials. Each row corresponds to a specific potential and color pattern (magenta, green, yellow, pink), and each column to a specific interaction model and shade pattern (dark shade, light shade). We only focus on late-time attractors. Black circles correspond to the late-time dark-energy dominated accelerating attractors, while red triangles indicate late-time accelerating scaling attractors. For the global interaction model $Q=\alpha H\rho_{m}$, all potentials can provide accelerating scaling attractors, while for local interaction rate $Q=\Gamma\rho_{m}$, only two potentials, namely $V(\phi)=\frac{1}{8\beta}\left(1-e^{-\kappa\mu\;\phi}\right)^{2}$ and $V(\phi)=c_{0}+c_{1}e^{\sqrt{\frac{2\kappa^{2}}{3}}\phi}+c_{2}e^{2\sqrt{\frac{2\kappa^{2}}{3}}\phi}$ can produce accelerating scaling attractors. Finally, note that all scaling solutions can transform to matter-dominated solution in the limiting case of the parameter region (see text).
  • Point $J_{4}$ exist always, it is saddle, and it also corresponds to a super-accelerated dark-energy dominated solution, with $H\longrightarrow 0$.
  • The curve of critical points $J_{5}$ corresponds to accelerating solutions. Since these points are normally hyperbolic, by applying the center manifold theorem we deduce that $J_{5}$ points are stable for $-\frac{1}{\sqrt{2}}\leq x_{c}<0$. Additionally, note that for this curve of critical points we have $0\leq\Omega_{\phi}\leq 1$ and $0\leq\Omega_{m}\leq 1$, and consequently $\Omega_{m}/\Omega_{\phi}\neq 0$, and hence $J_{5}$ can alleviate the coincidence problem. Finally, in the case $x_{c}=\pm\frac{1}{\sqrt{2}}$ one has $\Omega_{m}=1$, corresponding to a completely matter-dominated solution.
IV Summary and Conclusions
Cosmological scenarios which include interactions between dark matter and dark energy sectors have received significant attention in the scientific community since they exhibit many appealing features, such as the solution of the cosmic coincidence problem and the alleviation of $H_{0}$ and $S_{8}$ tensions. On the other hand, phantom cosmology has also attracted a significant amount of research, since it can efficiently describe the phantom crossing.
In the present work we performed a detailed investigation of interacting phantom cosmology, by applying the powerful method of dynamical system analysis, since this approach allows to extract information about the global features of the scenario, independently of the initial conditions or the specific evolution of the Universe. The novel ingredient of our work is the examination of new potentials, which indeed lead to new results even for the same interaction forms previously studied in the literature.
We considered two well-studied interaction forms, namely the global one $Q=\alpha H\rho_{m}$ and the local one $Q=\Gamma\rho_{m}$. Additionally, we considered four widely-used potentials for the phantom scalar field, namely A) $V(\phi)=V_{0}e^{-\kappa\lambda\phi}$, B) $V(\phi)=V_{0}\cosh(\kappa\eta\phi)$, C) $V(\phi)=\left(1-e^{-\kappa\mu\;\phi}\right)^{2}/8\beta$ and D) $V(\phi)=c_{0}+c_{1}e^{\sqrt{2\kappa^{2}/3}\phi}+c_{2}e^{2\sqrt{2\kappa^{2}/3}\phi}$, where the last two are the new ones that have not been studied before in the framework of interacting phantom cosmology. In all cases, we extracted the critical points and we examined their stability.
Our analysis extracted saddle matter-dominated points, stable dark-energy dominated points, and scaling accelerating solutions, that can attract the Universe at late times. As we discussed, some of the accelerating scaling attractors obtained for the last two potentials, when the local interaction is considered, are totally new. For completeness, in Fig. 4 we present the summary diagram that provides the qualitative behavior of the examined interacting phantom scenarios for different potentials.
In summary, our analysis revealed that interacting phantom cosmology can describe the Universe evolution very efficiently, namely the transition from a matter-dominated saddle point to a stable dark-energy dominated accelerated universe, or a stable accelerating phase in which dark-energy and dark matter co-exist, and hence alleviating the cosmic coincidence problem. These features make interacting phantom cosmology a good candidate for the description of nature.
V Acknowledgments
SH acknowledges the financial support from the University Grants Commission (UGC), Govt. of India (NTA Ref. No: 201610019097). SP and TS acknowledge the financial support from the Department of Science and Technology (DST), Govt. of India under the Scheme “Fund for Improvement of S&T Infrastructure (FIST)” (File No. SR/FST/MS-I/2019/41). ENS acknowledges the contribution of the LISA CosWG, and of COST Actions CA18108 ”Quantum Gravity Phenomenology in the multi-messenger approach” and CA21136” Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)”.