### v2

# Interacting phantom dark energy: new accelerating scaling attractors

Abstract
We perform a detailed investigation of interacting phantom cosmology, by applying the powerful method of dynamical system analysis. We consider two well-studied interaction forms, namely one global and one local one, while the novel ingredient of our work is the examination of new potentials for the phantom field. Our analysis shows the existence of saddle matter-dominated points, stable dark-energy dominated points, and scaling accelerating solutions, that can attract the Universe at late times. As we show, some of the stable accelerating scaling attractors, in which dark matter and dark energy can co-exist, alleviating the cosmic coincidence problem, are totally new, even for the previously studied interaction rates, and arise purely from the novel potential forms.
pacs:
98.80.-k, 95.36.+x, 95.35.+d, 98.80.Es
I Introduction
Dark matter (DM) and dark energy (DE) are two main ingredients of the universe, constituting of 95% of its total energy budget. The DM sector is responsible for the observed structure formation, while the DE sector is in charge of the late time accelerated expansion [1, 2]. The nature, origin and dynamics of both DE and DM are not yet clearly understood, nevertheless the $\Lambda$-Cold Dark Matter ($\Lambda$CDM) cosmological model, based on General Relativity, matches quite well with a large span of the astronomical datasets in which DE is a simple cosmological constant and DM is a cold pressureless sector. However, $\Lambda$CDM faces many challenges at both theoretical and observational levels [3, 4, 5], and among possible improvements is the consideration of a time varying nature in the DE sector.
In general, DE is assumed to evolve independently and interact only with gravity. However, there is no fundamental reason to exclude possible interactions between DE and other fields, especially with the DM sector whose microscopic nature is also unknown. The interaction between DE and DM can alleviate the cosmic coincidence problem [6, 7, 8, 9, 10, 11, 12, 13], while it can potentially alleviate the current cosmological tensions too [3, 4, 5, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Hence, over the last years interacting DM-DE scenarios have gained significant attention in the literature for their attractive features [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91] (for reviews see [92, 93]).
The interacting DM-DE models, widely known as Interacting DE (IDE) or Coupled DE, induce a richer dynamics in the universe evolution, recovering the non-interacting scenarios as special cases. The heart of the IDE models is the interaction function which characterizes the rate of energy and/or momentum transfer between the dark components, and consequently it affects the evolution of the universe. In order to investigate the dynamics of interacting cosmologies, one can confront them with observational datasets from various astronomical surveys and extract constraints on the involved functions and parameters. However, one can also apply the powerful approach of dynamical systems, which allows to extract analytical information independently of the initial conditions. The techniques of dynamical system analysis have been widely applied in the context of non-interacting and interacting dynamics [94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134].
In the present work we are interested in performing a dynamical-system analysis of interacting cosmology considering a phantom dark energy [135, 136, 137, 138, 139, 140, 141, 142]. In particular, we consider various interacting phantom DE scenarios, characterized by four different potentials and two specific interacting functions, one corresponding to the global interaction rate and the other corresponding to the local interaction rate. The examination of new potentials even for the known and studied cases of interaction form reveals new accelerating scaling attractors that were never studied in the literature. This is an important result, since it was believed that in general, due to the phantom nature of the acceleration, accelerating scaling solutions in phantom cosmology were very rare.
The manuscript is structured as follows. In section II we present the basic equations of general IDE scenarios and we construct the various specific models. In section III we perform the detailed phase space analysis extracting all possible solutions. Finally, section IV is devoted to the conclusions.
II Phantom Interacting Dark Energy
In this section we briefly present the scenario of phantom interacting dark energy, providing the main equations. We consider a flat Friedmann-Lemaître-Robertson-Walker (FLRW) line element given by
$ds^{2}=-dt^{2}+a^{2}(t)\left[dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})\right],$ (1)
where $a(t)$ is the expansion scale factor of the universe. Concerning the gravitational theory we consider standard general relativity, and we additionally consider the DM sector of the universe corresponding to a perfect fluid. Concerning the dark energy sector, we attribute it to a phantom scalar field $\phi$ with action
$S_{\phi}=\int d^{4}x\sqrt{-g}\left[\frac{1}{2}(\nabla\phi)^{2}-V(\phi)\right],$ (2)
where $V(\phi)$ is the potential. In the case of FLRW geometry, the energy density $\rho_{\phi}$ and pressure $p_{\phi}$ of the phantom fluid are given by
$\rho_{\phi}=-\frac{\dot{\phi}^{2}}{2}+V(\phi),$ (3)
$p_{\phi}=-\frac{\dot{\phi}^{2}}{2}-V(\phi),$ (4)
where a dot denotes derivative with respect to the cosmic time. Additionally, the Friedmann equations are written as
$3H^{2}=\kappa^{2}(\rho_{m}+\rho_{\phi})$ (5)
$2\dot{H}+3H^{2}=-\kappa^{2}(p_{m}+p_{\phi}),$ (6)
where $H=\dot{a}/a$ is the Hubble function (dots denote time-derivatives) $\kappa^{2}=8\pi G$ is the gravitational constant, and $\rho_{m}$, $p_{m}$ are respectively the energy density and pressure of the DM fluid.
We mention here that in phantom cosmology the kinetic term is negative, however the sign of the total energy of the phantom field depends on its potential too. Hence, if one chooses potentials for which $V(\phi)>\dot{\phi}^{2}/2$ holds, then $\rho_{\phi}$ remains positive. However, even if the total phantom energy becomes negative, it does not lead to pathologies, since the matter energy density in (5) compensates this situation and thus $H^{2}$ remains always positive, as it should be. In other words, the fact that in phantom cosmology one can always impose a specific potential, and/or interaction form, that will make everything well-behaved and self-consistent (this is the case in many modified gravity theories, e.g. Horndeski gravity [143], where in principle they can lead to negative $H^{2}$, however one chooses the involved functions in a suitable way in order to always have $H^{2}>0$).
In summary, at the classical level the negativity of $\rho_{\phi}$ is not a problem, and all energy conditions are the same with the case of canonical fields. Nevertheless, issues might arise concerning the quantum field theory description of phantom fields, since unbounded-from-below states may lead to problems. However, in the literature there have been various attempts in constructing a phantom theory consistent with the basic requirements of quantum field theory, for example considering that the phantom fields arise as effective descriptions of more fundamental canonical theories [144, 145].
The key idea of interacting cosmology is that although the total energy content of the universe is conserved (as a requirement of the diffeomorphism invariance) the separate DE and DM sectors are not conserved, but mutually interact. In particular, their conservation equations take the general form
$\dot{\rho}_{m}+3H(1+w_{m})\rho_{m}=-Q,$ (7)
$\dot{\rho}_{\phi}+3H(1+w_{\phi})\rho_{\phi}=Q,$ (8)
where $w_{m}=p_{m}/\rho_{m}$, $w_{\phi}=p_{\phi}/\rho_{\phi}$ are respectively the barotropic equation of state parameter of DM and DE, and $Q$ is the interaction function which characterizes the transfer of energy between the dark sectors. For $Q>0$, the energy flow occurs from DM to DE, while for $Q<0$ energy flows in the reverse direction, namely from DE to DM. Notice that eqn. (8) can alternatively be expressed in a Klein-Gordon-like form, namely
$\ddot{\phi}+3H\dot{\phi}-\frac{\partial V(\phi)}{\partial\phi}=-\frac{Q}{\dot{\phi}}.$ (9)
In this way, one can see that if the interaction function $Q(t)$ and the potential $V(\phi)$ of the scalar field are prescribed, then the evolution of the interacting fluids can be determined in principle.
In summary, the choice of both functions $Q(t)$ and $V(\phi)$ is the starting point for any analysis. In particular, using the conservation equations, one can find the evolution laws of the dark fluids as
[…]
In the following we focus on two interacting functions. The first interaction model is based on the most well-studied interaction form, namely [92, 93]
$\text{Model I}:\ \ \ Q=\alpha H\rho_{m},$ (12)
where $\alpha$ is a dimensionless parameter. In this scenario the interaction rate is proportional to the energy density of DM which is expected to be the case since interactions in physics and chemistry are typically stronger if the number density of the involved particles increases. Nevertheless, the interaction rate is proportional to the Hubble function too, which although very convenient at the calculation stage raises the criticism that the DM-DE interaction rate, which is expected to depend on local physics, depends on a global feature of the whole Universe, namely the Hubble function [146, 147, 148]. Hence, in this work we will additionally consider a second interaction function of the local form [146]
$\text{Model II}:\ \ \ Q=\Gamma\rho_{m},$ (13)
where $\Gamma$ is the dimensionfull coupling parameter. Finally, concerning the potential choices, we consider two forms that have been widely used in the literature, namely the well known exponential potential [94, 6, 149, 98, 150, 134] and the hyperbolic potential [114], given by
$V(\phi)=V_{0}e^{-\kappa\lambda\phi},$ (14)
$V(\phi)=V_{0}\cosh(\kappa\eta\phi),$ (15)
where $V_{0}>0$, and $\lambda$, $\eta$ are dimensionless free parameters. The third potential in this series is assumed to be [151]
$V(\phi)=\frac{1}{8\beta}\left(1-e^{-\kappa\mu\phi}\right)^{2},$ (16)
with $\mu$, $\beta$ dimensionless parameters. Let us mention here that such type of potential is not purely phenomenological, since similar forms may be derived in the context of $R^{2}$ gravity  [151]. Finally, we will also consider the following potential [152]
$V(\phi)=c_{0}+c_{1}e^{\sqrt{\frac{2\kappa^{2}}{3}}\phi}+c_{2}e^{2\sqrt{\frac{2\kappa^{2}}{3}}\phi},$ (17)
with $c_{0}$,$c_{1}$,$c_{2}$, being the non-negative model parameters. Similarly to the previous potential, this potential choice can be inspired from $R+R^{2}+\Lambda$ gravity [152], and hence it may lead to interesting cosmological implications. We mention that although the above potentials have been inspired by higher-order theories of gravity, such as the $R^{2}$ one, in the present analysis we handle them as scalar-field potentials in the framework of General Relativity. Note that the last two potentials have not been studied in detail in the framework of phantom cosmology.
In summary, the background evolution of the scenario at hand is determined by the two Friedmann equations (5),(6), together with the conservation equations (7),(8) and the above choices for $Q$ and $V(\phi)$. As usual, one can introduce the dark matter and dark energy density parameters as
$\Omega_{\phi}$$\equiv$ $\frac{\kappa^{2}\rho_{\phi}}{3H^{2}}$ (18)
$\Omega_{m}$$\equiv$ $\frac{\kappa^{2}\rho_{m}}{3H^{2}}\,,$ (19)
as well as the total, effective, equation-of-state parameter as
$w_{tot}$$\equiv$ $\frac{p_{\phi}+p_{m}}{\rho_{\phi}+\rho_{m}}=\frac{-\frac{1}{2}\dot{\phi}^{2}-V(\phi)+w_{m}\rho_{m}}{-\frac{1}{2}\dot{\phi}^{2}+V(\phi)+\rho_{m}}\,.$ (20)
III Dynamical system analysis
In this section we will perform a detailed phase-space and stability analysis of interacting phantom cosmology. This approach allows one to obtain information on the global dynamics and the behavior of a cosmological scenario, independently of the initial conditions and the specific universe evolution [94]. As usual, one first transforms the equations into an autonomous dynamical system, and extracts its critical points. Then, for each critical point one determines its stability by examining the sign of the eigenvalues of the perturbation matrix. The dynamical system analysis of interacting phantom cosmology has been performed in various articles [23, 98, 103, 153, 126, 134], however, it was restricted in specific potential choices. In this work we will perform it considering new potential forms and interaction models. In the following we focus on the case $w_{m}=0$, namely DM is cold, and we examine the various potentials and interaction forms separately.
Point $x$ $y$ Existence $E_{1}$ $E_{2}$ Stability $\Omega_{\phi}$ $w_{tot}$
$A_{1}$ $-\frac{\lambda}{\sqrt{6}}$ $\sqrt{1+\frac{\lambda^{2}}{6}}$ $\forall$ $\alpha$, $\lambda$ $-(3+\frac{\lambda^{2}}{2})$ $-(\lambda^{2}+\alpha+3)$ stable node for $-\alpha<3+\lambda^{2}$ $1$ $-1-\frac{\lambda^{2}}{3}$
for $-3<\alpha<0$
$A_{2}$ $\frac{3+\alpha}{\sqrt{6}\lambda}$ $\sqrt{-\frac{\alpha}{3}-\frac{(3+\alpha)^{2}}{6\lambda^{2}}}$ $\lambda^{2}\geq\frac{(3+\alpha)^{2}}{-\alpha}$ $\frac{R+\sqrt{S}}{4(3+\alpha)\lambda^{2}}$ $\frac{R-\sqrt{S}}{4(3+\alpha)\lambda^{2}}$ stable node for $\alpha<-3$ $-\frac{\alpha\lambda^{2}+(\alpha+3)^{2}}{3\lambda^{2}}$ $\frac{\alpha}{3}$
and for $\alpha<-3$ with $-(\alpha+3)>\lambda^{2}\geq\frac{(\alpha+3)^{2}}{-\alpha}$
$-(\alpha+3)\geq\lambda^{2}\geq\frac{(\alpha+3)^{2}}{-\alpha}$ (see Fig. 1)
Table 1: The critical points for the case of scalar-field potential $V(\phi)=V_{0}e^{-\kappa\lambda\phi}$ and interaction rate $Q=\alpha H\rho_{m}$, alongside their existence, eigenvalues at these points and stability conditions. Additionally, we present the corresponding values of the dark-energy density parameter $\Omega_{\phi}$ and the total equation-of-state parameter $w_{tot}$. We have defined $S=-1944\lambda^{2}-3240\alpha\lambda^{2}-2160\alpha^{2}\lambda^{2}-720\alpha^{3}\lambda^{2}-120\alpha^{4}\lambda^{2}-8\alpha^{5}\lambda^{2}-567\lambda^{4}-1404\alpha\lambda^{4}-882\alpha^{2}\lambda^{4}-204\alpha^{3}\lambda^{4}-15\alpha^{4}\lambda^{4}-180\alpha\lambda^{6}-72\alpha^{2}\lambda^{6}-4\alpha^{3}\lambda^{6}+4\alpha^{2}\lambda^{8}$ and $R=-9\lambda^{2}+6\alpha\lambda^{2}+3\alpha^{2}\lambda^{2}+2\alpha\lambda^{4}$.
III.1 Potential $V(\phi)=V_{0}e^{-\kappa\lambda\phi}$
In this subsection we focus on the analysis of the case of the exponential potential (14) and the two interaction models, one corresponding to the global interaction (12), and one corresponding to the local interaction (13).
III.1.1 Model I: $Q=\alpha H\rho_{m}$
In order to transform the cosmological equations into an autonomous form we introduce the dimensionless variables
$x=\frac{\kappa\dot{\phi}}{\sqrt{6}H},~{}~{}y=\frac{\kappa\sqrt{V}}{\sqrt{3}H}.$ (21)
Using these variables we can express the density parameters as
$\Omega_{\phi}=-x^{2}+y^{2},~{}~{}~{}~{}~{}\Omega_{m}=1+x^{2}-y^{2},$ (22)
while for the total equation-of-state parameter we have
$w_{tot}=-x^{2}-y^{2}.$ (23)
Finally, it proves convenient to introduce the deceleration parameter $q\equiv-1-\frac{\dot{H}}{H^{2}}$, which in the case of dust matter becomes simply $q=\frac{3w_{tot}+1}{2}$, and thus in terms of the dimensionless variables:
$q=\frac{1}{2}\left[1-3(x^{2}+y^{2})\right].$ (24)
In terms of the dimensionless variables (21), the autonomous system takes the form:
$x^{\prime}$ $=$ $-3x-\frac{\sqrt{6}}{2}\lambda y^{2}+\frac{3}{2}x\left(1-x^{2}-y^{2}\right)$ (25)
$-\frac{\alpha}{2x}\left(1+x^{2}-y^{2}\right),$
$y^{\prime}$ $=$ $-\frac{\sqrt{6}}{2}\lambda xy+\frac{3}{2}y\left(1-x^{2}-y^{2}\right),$ (26)
where primes denote derivatives with respect to $N\equiv\ln a$. The above system is invariant under the transformation $y\longrightarrow-y$, therefore the phase space is defined as
$\mathbb{D}=\left\{(x,y)\in\mathbb{R}^{2}:0\leq-x^{2}+y^{2}\leq 1,y\geq 0\right\}.$ (27)
Let us make here the following comment. As the denominator in the evolution equation (25) contains $x$ explicitly, the system exhibits a singularity along the line $x=0$, which implies that the system becomes ill defined along the line $x=0$. Now, as $x=0$ marks a singular line, we can partition the domain $\mathbb{D}$ into two regions: one for $x>0$ and the other for $x<0$, denoted as $\mathbb{D}_{+}=\mathbb{D}\setminus\{x\leq 0\}$ and $\mathbb{D}_{-}=\mathbb{D}\setminus\{x\geq 0\}$, respectively. After applying a re-parametrization, as outlined in Appendix A.1, one can use blow-up techniques [154] to investigate the asymptotic dynamics of the trajectories near the singular line $x=0$ of the system (25)-(26) in the regions $\mathbb{D}_{+}$ and $\mathbb{D}_{-}$. Subsequently, the nature of the trajectories near the singular line $x=0$ of the system (25)-(26) can be characterized by blow-down. This itself is a subject for further investigation. Therefore, in the present work we exclude $x=0$ from our discussion.
In order to extract the critical points, we impose $x^{\prime}=0$ and $y^{\prime}=0$. The critical points corresponding to an expanding Universe (namely having $y>0$), their existence conditions, their eigenvalues and their stability conditions are given in Table 1, alongside the corresponding values of the physically interesting quantities $\Omega_{\phi}$ and $w_{\rm tot}$. As we observe, for this scenario we have the following critical points.
Figure 1: The stability region of critical point $A_{2}$, for the case of scalar-field potential $V(\phi)=V_{0}e^{-\kappa\lambda\phi}$ and interaction rate $Q=\alpha H\rho_{m}$.
  • The critical point $A_{1}$ exists for all values of the parameters $\alpha$ and $\lambda$. It represents the dark-energy dominated ($\Omega_{\phi}=1$) solution with total equation of state $-1-\frac{\lambda^{2}}{3}$, and thus it corresponds to a phantom accelerating universe. It is a stable node when the condition $-\alpha<3+\lambda^{2}$ is satisfied, and as a late-time attractor this point can describe the late-time universe.
  • The critical point $A_{2}$ exists in the parameter region shown in Table 1. Since the eigenvalues of the corresponding linearized matrix are complicated, the stability conditions have to be determined numerically, and are presented in Fig. 1. When $\alpha<-1$, it corresponds to an accelerating universe. Moreover, at $A_{2}$ we have $\Omega_{m}/\Omega_{\phi}\neq 0$, and thus this point represents late-time accelerating scaling solution. Hence, in this case, the coincidence problem can be alleviated, as reported earlier in [98]. Additionally, we mention that the critical points $A_{2}$ can represent purely matter dominated solutions (namely having $\Omega_{m}=1$) for the parameters $\lambda$ and $\alpha$ satisfying $\lambda^{2}=-(\alpha+3)^{2}/\alpha$. Thus, depending on the parameter values, point $A_{2}$ can correspond to the above possibilities. As we shall see below, for different potential choices such late-time stable accelerating scaling solutions can be obtained in a larger parameter region compared to this case.
We close this subsection by mentioning that due to the non-compactness of the phase space in the $x-y$ direction discussed above, we should examine whether there are critical points at infinity. Hence, to complete the analysis, we compactify the phase space $\mathbb{D}$ by employing the Poincaré compactification technique [155, 154, 156, 121] and study the qualitative behaviour at the critical points at infinity. The details are provided in Appendix A.1. Our analysis reveals that the system possesses two critical points at infinity, namely $A^{\pm\infty}\left(\pm\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}},0\right)$, which exhibit unstable behavior, and thus they cannot attract the Universe at late times. Additionally, we can see that at these critical points, $q\longrightarrow-\infty$. However, note that this is not considered as a Big Rip, since it is obtained at infinite rather than at finite time (see the classifications of singularities in a phantom Universe, in [138]).
Point $x$ $y$ $z$ Existence $E_{1}$ $E_{2}$ $E_{3}$ Stability $\Omega_{\phi}$ $w_{tot}$
$B_{1}$ $-\frac{\lambda}{\sqrt{6}}$ $\sqrt{1+\frac{\lambda^{2}}{6}}$ 0 $\forall$ $\gamma$, $\lambda$ $-(3+\lambda^{2})$ $-(\frac{\lambda^{2}}{2}+3)$ $-\frac{\lambda^{2}}{2}$ always stable node 1 $-1-\frac{\lambda^{2}}{3}$
$B_{2}$ $-\frac{\lambda}{\sqrt{6}}$ $\sqrt{1+\frac{\lambda^{2}}{6}}$ 1 $\forall$ $\gamma$, $\lambda$ $-sgn(\gamma)\infty$ $-(\frac{\lambda^{2}}{2}+3)$ $\frac{\lambda^{2}}{2}$ always saddle 1 $-1-\frac{\lambda^{2}}{3}$
Table 2: The critical points for the case of scalar-field potential $V(\phi)=V_{0}e^{-\kappa\lambda\phi}$ and interaction rate $Q=\Gamma\rho_{m}$, alongside their eigenvalues and their existence and stability conditions. Additionally, we present the corresponding values of the dark-energy density parameter $\Omega_{\phi}$ and the total equation-of-state parameter $w_{tot}$.
III.1.2 Model II: $Q=\Gamma\rho_{m}$
In this case, apart from the dimensionless variables (21) we need one more, namely
$z=\frac{H_{0}}{H+H_{0}},$ (28)
with $H_{0}$ the value of the Hubble function at present. Hence, the autonomous system becomes:
$x^{\prime}$ $=$ $-3x-\frac{\sqrt{6}}{2}\lambda y^{2}+\frac{3}{2}x\left(1-x^{2}-y^{2}\right)$ (29)
$-\frac{\gamma z\left(1+x^{2}-y^{2}\right)}{2x(1-z)},$
$y^{\prime}$ $=$ $-\frac{\sqrt{6}}{2}\lambda xy+\frac{3}{2}y\left(1-x^{2}-y^{2}\right),$ (30)
$z^{\prime}$ $=$ $\frac{3}{2}z(1-z)\left(1-x^{2}-y^{2}\right),$ (31)
where $\gamma=\Gamma/H_{0}$ is the dimensionless coupling parameter. It is clear that the autonomous system (29)-(31) is invariant under the transformation $y\longrightarrow-y$, and therefore the physical region is defined as
$\mathbb{D}=\bigg{\{}(x,y,z)\in\mathbb{R}^{3}:0\leq-x^{2}+y^{2}\leq 1,y\geq 0,0\leq z\leq 1\bigg{\}}.$ (32)
Similar to the previous subsection, here we also observe that the explicit dependence of the denominator in equation (29) on $x$ leads to a singularity at $x=0$ plane, rendering the above system ill-defined at this plane. This issue can be addressed in an asymptotic way by dividing the entire phase space $\mathbb{D}$ into two separate domains, namely, $\mathbb{D}_{+}=\mathbb{D}\setminus\{x\leq 0\}$ and $\mathbb{D}_{-}=\mathbb{D}\setminus\{x\geq 0\}$, and by using re-parametrization and suitable blow-up transformations. Similarly, equation (29) features $(1-z)$ in the denominator, resulting in a singularity at the $z=1$ plane, which makes the system ill-defined at this location too. To resolve this, the system can be regularized by introducing the factor $(1-z)$. However, the analysis of the regularized dynamical system and the technique utilized in this article are equivalent (see [99, 157, 121, 134, 84]), and therefore, the regularized system does not add any new physical implications in this context. Note that the same methodology can be applied to all autonomous systems in the following where the dynamical variables are in the denominators.
The critical points, their existence conditions, their eigenvalues and their stability conditions are given in Table 2, alongside the values of $\Omega_{\phi}$ and $w_{tot}$. As we observe, for this scenario we have the following critical points.
  • The critical point $B_{1}$ exists always and corresponds to dark-energy dominated solutions. Since the total equation-of-state parameter is always less than $-1$, it corresponds to a phantom universe. Since, all the eigenvalues evaluated at $B_{1}$ are negative, it is always stable node. Thus, this point depicts a dark-energy dominated accelerated solution. However, note that it corresponds to $H\longrightarrow\infty$, as expected for a phantom, super-accelerating (i.e. with $\dot{H}>0$) universe.
  • As before, point $B_{2}$ corresponds to a dark-energy dominated solution, and exists always. It is always saddle and it has $H\longrightarrow 0$. Note that since it differs from point $B_{1}$ only in the $z$ value, both $B_{1}$ and $B_{2}$ have the same $\Omega_{\phi}$ and $w_{tot}$.
Finally, since the phase space is not compact in the $x-y$ directions, we expect to have critical points at infinity. In order to provide a complete analysis we apply the Poincaré compactification technique to examine the stability at these critical points. This analysis shows that the system has two points at infinity, namely $B^{\pm\infty}\left(\pm\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}},0,0\right)$, both of which display unstable behavior, and thus they cannot attract the Universe at late times. The details are presented in Appendix A.2. At these points one can observe that $q\longrightarrow-\infty$.
Finally, we mention that this local interaction model was investigated in [98], considering the exponential potential, and the absence of accelerating scaling attractors was also reported.
Point $x$ $y$ $\lambda$ Existence $E_{1}$ $E_{2}$ $E_{3}$ Stability $\Omega_{\phi}$ $w_{tot}$
$C_{1}$ $-\frac{\eta}{\sqrt{6}}$ $\sqrt{1+\frac{\eta^{2}}{6}}$ $\eta$ $\forall$ $\alpha$, $\eta$ $-(3+\alpha+\eta^{2})$ $-(3+\frac{\eta^{2}}{2})$ $-2\eta^{2}$ stable node for $-\alpha<3+\eta^{2}$ 1 $-1-\frac{\eta^{2}}{3}$
for $-3<\alpha<0$
$C_{2}$ $\frac{3+\alpha}{\sqrt{6}\eta}$   $\sqrt{-\frac{\alpha}{3}-\frac{(3+\alpha)^{2}}{6\eta^{2}}}$   $\eta$   $\eta^{2}\geq\frac{(3+\alpha)^{2}}{-\alpha}$ $\frac{P+\sqrt{S}}{4(3+\alpha)\eta^{2}}$ $\frac{P-\sqrt{S}}{4(3+\alpha)\eta^{2}}$ $2(3+\alpha)$ stable node for $\alpha<-3$ $-\frac{\alpha\eta^{2}+(\alpha+3)^{2}}{3\eta^{2}}$ $\frac{\alpha}{3}$
and for $\alpha<-3$ $-(\alpha+3)>\eta^{2}\geq\frac{(\alpha+3)^{2}}{-\alpha}$
$-(\alpha+3)\geq\eta^{2}\geq\frac{(\alpha+3)^{2}}{-\alpha}$ (see Fig. 2)
$C_{3}$ $\frac{\eta}{\sqrt{6}}$ $\sqrt{1+\frac{\eta^{2}}{6}}$ $-\eta$ $\forall$ $\alpha$, $\eta$ $-(3+\alpha+\eta^{2})$ $-(3+\frac{\eta^{2}}{2})$ $-2\eta^{2}$ stable node for $-\alpha<3+\eta^{2}$ 1 $-1-\frac{\eta^{2}}{3}$
for $-3<\alpha<0$
$C_{4}$ $-\frac{3+\alpha}{\sqrt{6}\eta}$ $\sqrt{-\frac{\alpha}{3}-\frac{(3+\alpha)^{2}}{6\eta^{2}}}$ $-\eta$ $\eta^{2}\geq\frac{(3+\alpha)^{2}}{-\alpha}$ $\frac{P+\sqrt{S}}{4(3+\alpha)\eta^{2}}$ $\frac{P-\sqrt{S}}{4(3+\alpha)\eta^{2}}$ $2(3+\alpha)$ stable node for $\alpha<-3$ $-\frac{\alpha\eta^{2}+(\alpha+3)^{2}}{3\eta^{2}}$ $\frac{\alpha}{3}$
and for $\alpha<-3$ $-(\alpha+3)>\eta^{2}\geq\frac{(\alpha+3)^{2}}{-\alpha}$
$-(\alpha+3)\geq\eta^{2}\geq\frac{(\alpha+3)^{2}}{-\alpha}$ (see Fig. 2)
Table 3: The critical points for the case of $V(\phi)=V_{0}\cosh{(\kappa\eta\phi)}$ and interaction rate $Q=\alpha H\rho_{m}$, alongside their existence, their eigenvalues and stability conditions. Additionally, we present the corresponding values of the dark-energy density parameter $\Omega_{\phi}$ and the total equation-of-state parameter $w_{tot}$. We have defined $S=-1944\eta^{2}-3240\alpha\eta^{2}-2160\alpha^{2}\eta^{2}-720\alpha^{3}\eta^{2}-120\alpha^{4}\eta^{2}-8\alpha^{5}\eta^{2}-567\eta^{4}-1404\alpha\eta^{4}-882\alpha^{2}\eta^{4}-204\alpha^{3}\eta^{4}-15\alpha^{4}\eta^{4}-180\alpha\eta^{6}-72\alpha^{2}\eta^{6}-4\alpha^{3}\eta^{6}+4\alpha^{2}\eta^{8}$ and $P=-9\eta^{2}+6\alpha\eta^{2}+3\alpha^{2}\eta^{2}+2\alpha\eta^{4}$.
III.2 Potential: $V(\phi)=V_{0}\cosh{(\kappa\eta\phi)}$
In this subsection we focus on the phase space analysis of the autonomous system for the hyperbolic potential (15) considering the global interaction rate (12) and the local interaction rate (13).
III.2.1 Model I: $Q=\alpha H\rho_{m}$
In order to transform the cosmological equations into an autonomous form, apart from the dimensionless variables (21) we need one more, namely
$\lambda=-\frac{V_{\phi}}{\kappa V},$ (33)
where a $\phi$-subscript denotes derivative with respect to $\phi$. If $\lambda(\phi)$ is invertible, and one can extract the form $\phi=\phi(\lambda)$, then the extraction of an autonomous system can be obtained by simply calculating the parameter $\Xi=\frac{VV_{\phi\phi}}{{V_{\phi}}^{2}}$. Using the potential $V(\phi)=V_{0}\cosh{(\kappa\eta\phi)}$, we find $\lambda=-\eta\tanh{(\kappa\eta\phi)}$ and $\Xi=\frac{\eta^{2}}{\lambda^{2}}$ Therefore, the range of $\lambda$ is $[-|\eta|,|\eta|]$. Finally, the autonomous system takes the form:
$x^{\prime}$ $=$ $-3x-\frac{\sqrt{6}}{2}\lambda y^{2}+\frac{3}{2}x\left(1-x^{2}-y^{2}\right)$ (34)
$-\frac{\alpha}{2x}\left(1+x^{2}-y^{2}\right),$
$y^{\prime}$ $=$ $-\frac{\sqrt{6}}{2}\lambda xy+\frac{3}{2}y\left(1-x^{2}-y^{2}\right),$ (35)
$\lambda^{\prime}$ $=$ $\sqrt{6}\left(\lambda^{2}-\eta^{2}\right)x.$ (36)
Since the above autonomous system is invariant under the transformation $y\longrightarrow-y$, the phase space domain $\mathbb{D}$ is defined as $\mathbb{D}=\left\{(x,y,\lambda)\in\mathbb{R}^{3}:0\leq-x^{2}+y^{2}\leq 1,y\geq 0,-|\eta|\leq\lambda\leq|\eta|\right\}$. The presence of $x$ in the denominator of equation (34) introduces a singularity at $x=0$ plane, causing the system to become ill-defined at this plane, and this can be resolved in a similar fashion as discussed in the previous subsections.
The critical points, their existence conditions, their eigenvalues and their stability conditions are given in Table 3, alongside the corresponding values of $\Omega_{\phi}$ and $w_{tot}$. For this scenario we have the following critical points.
  • The critical point $C_{1}$ exists for all $\alpha$ and $\eta$ and it corresponds to an accelerating universe, and it is stable in the region $-\alpha<3+\eta^{2}$. Moreover, at this point we have $\Omega_{\phi}=1$, which implies complete dark energy domination. Therefore, it describes the late-time dark-energy dominated accelerated universe.
  • The critical point $C_{2}$ describes an accelerating universe, since in their stability region, i.e. for $\alpha<-1$, we have $w_{tot}<-1/3$. The stability region is graphically shown in Fig. 2. As we observe, in this case we find almost similar stability region as in Fig. 1 for the exponential potential, which was expected since the present hyperbolic cosinus potential is related to the exponential potential.
    Again, we notice that at $C_{2}$ the ratio of $\Omega_{m}$ and $\Omega_{\phi}$ is non zero (namely $\Omega_{m}/\Omega_{\phi}\neq 0$), and thus this point represents a late-time accelerating scaling solution, which is stable in the aforementioned specific parameter region. Thus, for this case, the coincidence problem can be alleviated. Finally, one can further notice that $C_{2}$ can represent completely matter dominated solutions for a specific parameter space satisfying $\eta^{2}=-(\alpha+3)^{2}/\alpha$.
  • The critical point $C_{3}$ is stable for $-\alpha<3+\eta^{2}$, and the values of $\Omega_{\phi}$ and $w_{\rm tot}$ are $1$ and $-1-\frac{\eta^{2}}{3}<-1$ respectively, therefore this point represents dark-energy dominated accelerated solutions. Hence, point $C_{3}$, similar to $C_{1}$, is a candidate for the late-time universe.
  • The critical point $C_{4}$ has $w_{\rm tot}=\frac{\alpha}{3}<-\frac{1}{3}$ for $\alpha<-1$, and thus it corresponds to accelerated solutions. The stability region is graphically shown in Fig. 2, which is the same with point $C_{2}$ since they have the same eigenvalues. At this critical point $\Omega_{m}/\Omega_{\phi}\neq 0$ and therefore it corresponds to late-time accelerating stable solutions that can alleviate the coincidence problem. Finally, completely matter-dominated solutions are obtained for $\eta^{2}=-(\alpha+3)^{2}/\alpha$.
Finally, investigating the qualitative behaviour at the critical points at infinity (see Appendix B.1) reveals that the phase space domain $\mathbb{D}$ contains four such critical points, namely $C_{1}^{\pm\infty}\left(\pm\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}},0,0\right)$ and $C_{2}^{\pm\infty}\left(\pm\sqrt{\frac{3}{8}},\sqrt{\frac{3}{8}},\mp\frac{1}{2},0\right)$. The first two points are unstable, while the remaining two exhibit saddle behavior. At these points, we deduce that $q\longrightarrow-\infty$.
Figure 2: The stability region of the critical points $C_{2}$ and $C_{4}$, for the case of scalar-field potential $V(\phi)=V_{0}\cosh{(\kappa\eta\phi)}$ and interaction rate $Q=\alpha H\rho_{m}$.
III.2.2 Model II: $Q=\Gamma\rho_{m}$
Point $x$ $y$ $\lambda$ $z$ Existence $E_{1}$ $E_{2}$ $E_{3}$ $E_{4}$ Stability $\Omega_{\phi}$ $w_{tot}$
$D_{1}$ $-\frac{\eta}{\sqrt{6}}$   $\sqrt{1+\frac{\eta^{2}}{6}}$   $\eta$   0   $\forall$ $\gamma$, $\eta$ $-(3+\eta^{2})$ $-(\frac{\eta^{2}}{2}+3)$ $-2\eta^{2}$ $-\frac{\eta^{2}}{2}$ stable node 1 $-1-\frac{\eta^{2}}{3}$
$D_{2}$ $-\frac{\eta}{\sqrt{6}}$   $\sqrt{1+\frac{\eta^{2}}{6}}$   $\eta$   1   $\forall$ $\gamma$, $\eta$ $-sgn(\gamma)\infty$ $-(\frac{\eta^{2}}{2}+3)$ $-2\eta^{2}$ $\frac{\eta^{2}}{2}$ always saddle 1 $-1-\frac{\eta^{2}}{3}$
$D_{3}$ $\frac{\eta}{\sqrt{6}}$   $\sqrt{1+\frac{\eta^{2}}{6}}$   $-\eta$   0   $\forall$ $\gamma$, $\eta$ $-(3+\eta^{2})$ $-(\frac{\eta^{2}}{2}+3)$ $-2\eta^{2}$ $-\frac{\eta^{2}}{2}$ stable node 1 $-1-\frac{\eta^{2}}{3}$
$D_{4}$ $\frac{\eta}{\sqrt{6}}$   $\sqrt{1+\frac{\eta^{2}}{6}}$   $-\eta$   1   $\forall$ $\gamma$, $\eta$ $-sgn(\gamma)\infty$ $-(\frac{\eta^{2}}{2}+3)$ $-2\eta^{2}$ $\frac{\eta^{2}}{2}$ always saddle 1 $-1-\frac{\eta^{2}}{3}$
Table 4: The critical points for the case of $V(\phi)=V_{0}\cosh{(\kappa\eta\phi)}$ and interaction rate $Q=\Gamma\rho_{m}$, alongside their eigenvalues and their existence and stability conditions. Additionally, we present the corresponding values of the dark-energy density parameter $\Omega_{\phi}$ and the total equation-of-state parameter $w_{tot}$.
Using the dimensionless variables $x$, $y$ defined in (21) alongside the dimensionless variables $z$ and $\lambda$ of (28) and (33), we obtain the autonomous system as
$x^{\prime}$ $=$ $-3x-\frac{\sqrt{6}}{2}\lambda y^{2}+\frac{3}{2}x\left(1-x^{2}-y^{2}\right)$ (37)
$-\frac{\gamma z\left(1+x^{2}-y^{2}\right)}{2x(1-z)},$
$y^{\prime}$ $=$ $-\frac{\sqrt{6}}{2}\lambda xy+\frac{3}{2}y\left(1-x^{2}-y^{2}\right),$ (38)
$\lambda^{\prime}$ $=$ $\sqrt{6}\left(\lambda^{2}-\eta^{2}\right)x,$ (39)
$z^{\prime}$ $=$ $\frac{3}{2}z(1-z)\left(1-x^{2}-y^{2}\right),$ (40)
where $\gamma=\Gamma/H_{0}$ is the dimensionless coupling parameter. The above autonomous system is invariant under the transformation $y\longrightarrow-y$, and the phase space $\mathbb{D}$ is given by
$\mathbb{D}=\bigg{\{}(x,y,\lambda,z)\in\mathbb{R}^{4}:0\leq-x^{2}+y^{2}\leq 1,y\geq 0,$
$-|\eta|\leq\lambda\leq|\eta|,0\leq z\leq 1\bigg{\}}.$ (41)
Since $x$ and $(1-z)$ appear in the denominator of the evolution equation (37), the system exhibits singularities at $x=0$ and $z=1$ surfaces, making it ill-defined along these surfaces.
The critical points, their existence conditions, their stability conditions, and the values of $\Omega_{\phi}$ and $w_{tot}$, are given in Table 4. We have the following critical points.
  • The critical point $D_{1}$ exists always and it corresponds to super-accelerated dark-energy domination, exhibiting $H\longrightarrow\infty$. It is always stable and it can describe the Universe at late times.
  • The critical point $D_{2}$ exists always, corresponding to dark energy dominated accelerated solutions, with $H\longrightarrow 0$, and it is always saddle.
  • Point $D_{3}$ exists always, and it corresponds to dark-energy dominated accelerated solutions, with $H\longrightarrow\infty$. Since it is stable, it is also a suitable candidate for late-time evolution of the universe.
  • Point $D_{4}$ exists always, it is saddle, and it corresponds to dark-energy dominated accelerated solutions with $H\longrightarrow 0$.
Finally, with the help of the the Poincaré compactification it follows that the system encounters two critical points at infinity, namely $D^{+\infty}\Bigl{\{}\left(y_{1},y_{2},y_{3},y_{4},y_{5}\right)\in\mathbb{S}^{4}:y_{5}=0,y_{4}=0,y_{1}^{2}=y_{2}^{2},y_{1}>0,y_{2}>0\Bigr{\}}$, and $D^{-\infty}\Bigl{\{}\left(y_{1},y_{2},y_{3},y_{4},y_{5}\right)\in\mathbb{S}^{4}:y_{5}=0,y_{4}=0,y_{1}^{2}=y_{2}^{2},y_{1}<0,y_{2}>0\Bigr{\}}$, where $q\longrightarrow-\infty$. The analysis reveals that these critical points exhibit unstable behavior. The details of the analysis are presented in Appendix B.2.
In summary, the hyperbolic potential with the local interaction rate leads to results similar to the exponential potential, since the two potentials are related. Thus, in this scenario, late-time scaling attractors are not possible.
III.3 Potential: $V(\phi)=\frac{1}{8\beta}\left(1-e^{-\kappa\mu\;\phi}\right)^{2}$
In this subsection we focus on the phase space analysis of the autonomous system for the potential (16), considering as usual both the global interaction rate (12) and the local interaction rate (13).
III.3.1 Model I: $Q=\alpha H\rho_{m}$
Point $x$ $y$ $\lambda$ Existence $E_{1}$ $E_{2}$ $E_{3}$ Stability $\Omega_{\phi}$ $w_{tot}$
$F_{1}$ $-\frac{2\mu}{\sqrt{6}}$ $\sqrt{1+\frac{2\mu^{2}}{3}}$ $2\mu$ $\forall$ $\alpha$, $\mu$ $-(3+\alpha+4\mu^{2})$ $-3-2\mu^{2}$ $-2\mu^{2}$ stable node for $-\alpha<3+4\mu^{2}$ 1 $-1-\frac{4\mu^{2}}{3}$
for $-3<\alpha<0$
$F_{2}$ $\frac{3+\alpha}{2\sqrt{6}\mu}$ $\sqrt{-\frac{\alpha}{3}-\frac{(3+\alpha)^{2}}{24\mu^{2}}}$ $2\mu$ $\mu^{2}\geq\frac{(3+\alpha)^{2}}{-4\alpha}$ $\frac{P+\sqrt{S}}{4(3+\alpha)\mu^{2}}$ $\frac{P-\sqrt{S}}{4(3+\alpha)\mu^{2}}$ $\frac{1}{2}(3+\alpha)$ stable node for $\alpha<-3$ $-\frac{4\alpha\mu^{2}+(\alpha+3)^{2}}{12\mu^{2}}$ $\frac{\alpha}{3}$
and for $\alpha<-3$ $-\frac{(\alpha+3)}{4}>\mu^{2}\geq\frac{(\alpha+3)^{2}}{-4\alpha}$
$-\frac{(\alpha+3)}{4}\geq\mu^{2}\geq\frac{(\alpha+3)^{2}}{-4\alpha}$ (see Fig. 3)
$F_{3}$ $x_{c}$ $\sqrt{1-x_{c}^{2}}$ 00 $\forall\mu$, $\alpha=-3$, 00 $-6$ $-\sqrt{6}\mu x_{c}$ stable for $\mu>0$ and $x_{c}>0$ $1-2x_{c}^{2}$ $-1$
$|x_{c}|\leq\frac{1}{\sqrt{2}}$ with $x_{c}\neq 0$ or for $\mu<0$ and $x_{c}<0$
Table 5: The critical points for the case of $V(\phi)=\frac{1}{8\beta}\left(1-e^{-\kappa\mu\;\phi}\right)^{2}$ and interaction rate $Q=\alpha H\rho_{m}$, alongside their existence, eigenvalues and stability conditions. Furthermore, we present the corresponding values of the dark-energy density parameter $\Omega_{\phi}$ and the total equation-of-state parameter $w_{tot}$. We have defined $S=-486\mu^{2}-810\alpha\mu^{2}-540\alpha^{2}\mu^{2}-180\alpha^{3}\mu^{2}-30\alpha^{4}\mu^{2}-2\alpha^{5}\mu^{2}-567\mu^{4}-1404\alpha\mu^{4}-882\alpha^{2}\mu^{4}-204\alpha^{3}\mu^{4}-15\alpha^{4}\mu^{4}-720\alpha\mu^{6}-288\alpha^{2}\mu^{6}-16\alpha^{3}\mu^{6}+24\alpha^{2}\mu^{8}$ and $P=-9\mu^{2}+6\alpha\mu^{2}+3\alpha^{2}\mu^{2}+8\alpha\mu^{4}$.
In this case we derive the autonomous system in terms of the dimensionless variables $x$, $y$, $\lambda$, defined in (21) and (33). As we can see, the expressions of the parameters $\lambda=-\frac{V_{\phi}}{\kappa V}$ and $\Xi=\frac{VV_{\phi\phi}}{V_{\phi}^{2}}$ take the form $\lambda=-2\mu\frac{e^{-\kappa\mu\phi}}{1-e^{-\kappa\mu\phi}}$ ($\mu\neq 0$) and $\Xi=\frac{\mu}{\lambda}+\frac{1}{2}$, respectively. Therefore, the autonomous system becomes:
$x^{\prime}$ $=$ $-3x-\frac{\sqrt{6}}{2}\lambda y^{2}+\frac{3}{2}x\left(1-x^{2}-y^{2}\right)$ (42)
$-\frac{\alpha}{2x}\left(1+x^{2}-y^{2}\right),$
$y^{\prime}$ $=$ $-\frac{\sqrt{6}}{2}\lambda xy+\frac{3}{2}y\left(1-x^{2}-y^{2}\right),$ (43)
$\lambda^{\prime}$ $=$ $\frac{\sqrt{6}}{2}\left(\lambda^{2}-2\mu\lambda\right)x.$ (44)
The above system is invariant under $y\longrightarrow-y$, and its phase space region $\mathbb{D}$ is defined by $\mathbb{D}=\left\{(x,y,\lambda)\in\mathbb{R}^{3}:0\leq-x^{2}+y^{2}\leq 1,y\geq 0\right\}$. The explicit presence of $x$ in the denominator of equation (42) leads to a singularity at $x=0$ plane, where the system is no longer well-defined.
The critical points, their existence conditions, their eigenvalues and their stability conditions are given in Table 5, alongside the corresponding values of $\Omega_{\phi}$ and $w_{tot}$. We have the following critical points.
Figure 3: The stability region of the critical point $F_{2}$, for the case of scalar-field potential $V(\phi)=\frac{1}{8\beta}\left(1-e^{-\kappa\mu\;\phi}\right)^{2}$ and interaction rate $Q=\alpha H\rho_{m}$.
Point $x$ $y$ $\lambda$ $z$ Existence $E_{1}$ $E_{2}$ $E_{3}$ $E_{4}$ Stability $\Omega_{\phi}$ $w_{tot}$
$G_{1}$ $-\frac{2\mu}{\sqrt{6}}$ $\sqrt{1+\frac{2\mu^{2}}{3}}$ $2\mu$ 0 $\forall$ $\gamma$, $\mu$ $-3-4\mu^{2}$ $-2\mu^{2}-3$ $-2\mu^{2}$ $-2\mu^{2}$ stable node 1 $-1-\frac{4\mu^{2}}{3}$
$G_{2}$ $-\frac{2\mu}{\sqrt{6}}$ $\sqrt{1+\frac{2\mu^{2}}{3}}$ $2\mu$ 1 $\forall$ $\gamma$, $\mu$ $-sgn(\gamma)\infty$ $-2\mu^{2}-3$ $-2\mu^{2}$ $2\mu^{2}$ always saddle 1 $-1-\frac{4\mu^{2}}{3}$
$G_{3}$ $x_{c}$ $\sqrt{1-{x_{c}}^{2}}$ 0 $\frac{3}{3-\gamma}$ $|x_{c}|\leq\frac{1}{\sqrt{2}}$, $x_{c}\neq 0$ 00 $-\sqrt{6}\mu x_{c}$ $-3\left(1-\sqrt{1-{x_{c}}^{2}}\right)$ $-3\left(1+\sqrt{1-{x_{c}}^{2}}\right)$ stable $1-2{x_{c}}^{2}$ $-1$
with $\gamma<0$ for $\mu x_{c}>0$
Table 6: The critical points for the case of $V(\phi)=\frac{1}{8\beta}\left(1-e^{-\kappa\mu\;\phi}\right)^{2}$ and interaction rate $Q=\Gamma\rho_{m}$, alongside their eigenvalues and their existence and stability conditions. Additionally, we present the corresponding values of the dark-energy density parameter $\Omega_{\phi}$ and the total equation-of-state parameter $w_{tot}$.
  • Point $F_{1}$ exists always, and it corresponds to a dark energy dominated accelerating solution. Moreover, it is a stable node for $-\alpha<3+4\mu^{2}$ and hence it can describe the late-time universe.
  • The critical point $F_{2}$ corresponds to accelerating solutions for $\alpha<-1$, where $w_{tot}<-\frac{1}{3}$. We mention that at this point we have $\Omega_{m}/\Omega_{\phi}\neq 0$ corresponding to accelerating scaling solution, which is also stable in its stability region as shown in Table 5 and in Fig. 3, and thus it can alleviate the coincidence problem. Finally, one can see that for $\mu^{2}=-(\alpha+3)^{2}/4\alpha$ point $F_{1}$ corresponds to completely matter dominated solutions.
  • The curve of critical points $F_{3}$ has $\lambda=0$, and it was absent for the exponential and hyperbolic potentials. In general these points correspond to scaling accelerating solutions, while for $x_{c}=\pm\frac{1}{\sqrt{2}}$ they correspond to completely matter-dominated points. Since, one eigenvalue is zero and the other two are negative, they are normally hyperbolic critical points. By applying the center manifold theorem [103] we deduce that these critical points behave as stable points for $\mu x_{c}>0$. Thus, these points can alleviate the coincidence problem for the model parameter $\alpha=-3$.
Finally, from the critical point analysis at infinity (see Appendix C.1), we deduce that the phase space domain $\mathbb{D}$ includes four such critical points, namely $E_{1}^{\pm\infty}\left(\pm\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}},0,0\right)$, and $E_{2}^{\pm\infty}\left(\pm\sqrt{\frac{2}{7}},\sqrt{\frac{2}{7}},\mp\sqrt{\frac{3}{7}},0\right)$. The first two points are unstable, while the last two are saddles. At these critical points, it follows that $q\longrightarrow-\infty$.
III.3.2 Model II: $Q=\Gamma\rho_{m}$
In this case, using the dimensionless variables $x$, $y$, $\lambda$, $z$ of (21), (28) and (33), we result to the autonomous system
$x^{\prime}$ $=$ $-3x-\frac{\sqrt{6}}{2}\lambda y^{2}+\frac{3}{2}x\left(1-x^{2}-y^{2}\right)$ (45)
$-\frac{\gamma z\left(1+x^{2}-y^{2}\right)}{2x(1-z)},$
$y^{\prime}$ $=$ $-\frac{\sqrt{6}}{2}\lambda xy+\frac{3}{2}y\left(1-x^{2}-y^{2}\right),$ (46)
$\lambda^{\prime}$ $=$ $\frac{\sqrt{6}}{2}\left(\lambda^{2}-2\mu\lambda\right)x,$ (47)
$z^{\prime}$ $=$ $\frac{3}{2}z(1-z)\left(1-x^{2}-y^{2}\right),$ (48)
where $\gamma=\Gamma/H_{0}$. The above autonomous system is invariant under $y\longrightarrow-y$, and the phase space domain is defined as $\mathbb{D}=\left\{(x,y,\lambda,z)\in\mathbb{R}^{4}:0\leq-x^{2}+y^{2}\leq 1,y\geq 0,0\leq z\leq 1\right\}$. The dynamical system has two singular surfaces, namely $x=0$ and $z=1$, in which it is ill-defined. The properties and features of the critical points are summarized in Table 6. They are the following.
Point $x$ $v$ $w$ $\xi$ Existence Stability $\Omega_{\phi}$ $w_{tot}$
$I_{1}$ $\frac{2}{3}$ 0 0 $\frac{\sqrt{13}}{3}$ $\forall\alpha$ stable for $\alpha>-\frac{17}{3}$ 1 $-\frac{17}{9}$
[…]
$\cup\left(-3,\frac{-13+\sqrt{88}}{3}\right]$ otherwise saddle or unstable
$I_{3}$ $\frac{1}{3}$ 0 $\frac{\sqrt{10}}{3}$ 0 $\forall\alpha$ always saddle 1 $-\frac{11}{9}$
[…]
$\cup\left(-3,\frac{-10+\sqrt{19}}{3}\right]$
Table 7: The critical points for the case of $V(\phi)=c_{0}+c_{1}e^{\sqrt{\frac{2\kappa^{2}}{3}}\phi}+c_{2}e^{2\sqrt{\frac{2\kappa^{2}}{3}}\phi}$ and interaction rate $Q=\alpha H\rho_{m}$, alongside their existence and stability conditions. Furthermore, we present the corresponding values of the dark-energy density parameter $\Omega_{\phi}$ and the total equation-of-state parameter $w_{tot}$.
  • The critical point $G_{1}$ exists always, corresponding to dark-energy dominated super-accelerating solutions, with $H\longrightarrow\infty$, and it is always stable node. Thus, $G_{1}$ is a candidate for describing the late-time accelerated universe.
  • Point $G_{2}$ exists always, corresponding to dark-energy dominated accelerating solutions, with $H\longrightarrow 0$, and it is a saddle point.
  • The curves of critical points $G_{3}$ exist for $|x_{c}|\leq\frac{1}{\sqrt{2}}$ with $\gamma<0$. They are normally hyperbolic, and by applying the center manifold theorem in order to determine their stability [103] we deduce that $G_{3}$ are stable for $\mu x_{c}>0$. Note that in general, for $|x_{c}|<\frac{1}{\sqrt{2}}$, these points have $10<\Omega_{\phi}<$ 1, and thus they correspond to a late-time stable accelerating solution with $\Omega_{m}/\Omega_{\phi}\neq 0$, namely with a scaling nature that can alleviate the coincidence problem. Finally, in the limiting case where $x_{c}=\pm\frac{1}{\sqrt{2}}$ they lead to $\Omega_{\phi}=0$ and thus to $\Omega_{m}=1$, corresponding to matter-dominated solutions.
Lastly, the system encounters two critical points at infinity, namely $F^{+\infty}\Bigl{\{}\left(y_{1},y_{2},y_{3},y_{4},y_{5}\right)\in\mathbb{S}^{4}:y_{5}=0,y_{4}=0,y_{1}^{2}=y_{2}^{2},y_{1}>0,y_{2}>0\Bigr{\}}$ and $F^{-\infty}\Bigl{\{}\left(y_{1},y_{2},y_{3},y_{4},y_{5}\right)\in\mathbb{S}^{4}:y_{5}=0,y_{4}=0,y_{1}^{2}=y_{2}^{2},y_{1}<0,y_{2}>0\Bigr{\}}$, where $q\longrightarrow-\infty$. These critical points exhibit unstable behavior. The details of the analysis are presented in Appendix C.2.
In summary, under the local interaction rate the potential $V(\phi)=\frac{1}{8\beta}\left(1-e^{-\kappa\mu\;\phi}\right)^{2}$ is able to accept late-time stable scaling accelerating solutions, which are not possible for the previous potentials, and that is why they were not studied in the literature. This is a novel result of the present work.
III.4 Potential: $V(\phi)=c_{0}+c_{1}e^{\sqrt{\frac{2\kappa^{2}}{3}}\phi}+c_{2}e^{2\sqrt{\frac{2\kappa^{2}}{3}}\phi}$
In this subsection we focus on the last potential function (17). In order to obtain the autonomous form for this potential case we introduce three new dimensionless variables $v$, $w$, $\xi$ as
$v=\frac{\kappa\sqrt{c_{0}}}{\sqrt{3}H},~{}w=\frac{\kappa\sqrt{c_{1}e^{\sqrt{\frac{2\kappa^{2}}{3}}\phi}}}{\sqrt{3}H},~{}\xi=\frac{\kappa\sqrt{c_{2}e^{2\sqrt{\frac{2\kappa^{2}}{3}}\phi}}}{\sqrt{3}H}.$ (49)
In terms of these variables, as well as $x$ defined in (21), the various cosmological parameters become
$\Omega_{m}=1+x^{2}-v^{2}-w^{2}-\xi^{2},$
$\Omega_{\phi}=-x^{2}+v^{2}+w^{2}+\xi^{2},$
$w_{tot}=-x^{2}-v^{2}-w^{2}-\xi^{2},$
$q=\frac{1}{2}\left[1-3(x^{2}+v^{2}+w^{2}+\xi^{2})\right].$ (50)
III.4.1 Model I: $Q=\alpha H\rho_{m}$
Using the above dimensionless variables we obtain the autonomous system as
$x^{\prime}$ $=$ $-\frac{3}{2}x\left(1+x^{2}+v^{2}+w^{2}+\xi^{2}\right)+w^{2}+2\xi^{2}$ (51)
$-\frac{\alpha\left(1+x^{2}-v^{2}-w^{2}-\xi^{2}\right)}{2x},$
$v^{\prime}$ $=$ $\frac{3}{2}v\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right),$ (52)
$w^{\prime}$ $=$ $w\left[x+\frac{3}{2}\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right)\right],$ (53)
$\xi^{\prime}$ $=$ $\xi\left[2x+\frac{3}{2}\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right)\right].$ (54)
This system is invariant under the transformations $v\longrightarrow-v$, $w\longrightarrow-w$ and $\xi\longrightarrow-\xi$ separately, and thus the physical region is defined as
$\mathbb{D}$ $=$ $\bigg{\{}(x,v,w,\xi)\in\mathbb{R}^{4}:0\leq-x^{2}+v^{2}+w^{2}+\xi^{2}\leq 1,$ (55)
$v\geq 0,w\geq 0,\xi\geq 0\bigg{\}}.$
The system exhibits a singular surface at $x=0$, where it becomes ill-defined. The critical points, the existence and stability conditions as well as the corresponding values of $\Omega_{\phi}$ and $w_{tot}$ are summarized in Table 7, while the corresponding eigenvalues are displayed in Table 8. For this scenario we have the following critical points.
Point $E_{1}$ $E_{2}$ $E_{3}$ $E_{4}$
$I_{1}$ $-\frac{17}{3}-\alpha$ $-\frac{4}{3}$ $-\frac{2}{3}$ $-\frac{13}{3}$
$I_{2}$ $\frac{3+\alpha}{4}$ $\frac{3+\alpha}{2}$ $\frac{T^{\prime}-\sqrt{V^{\prime}}}{12(3+\alpha)}$ $\frac{T^{\prime}+\sqrt{V^{\prime}}}{12(3+\alpha)}$
$I_{3}$ $-\frac{11}{3}-\alpha$ $-\frac{1}{3}$ $-\frac{10}{3}$ $\frac{1}{3}$
$I_{4}$ $\frac{3+\alpha}{2}$ $-\frac{3+\alpha}{2}$ $\frac{R^{\prime}-\sqrt{S^{\prime}}}{12(3+\alpha)}$ $\frac{R^{\prime}+\sqrt{S^{\prime}}}{12(3+\alpha)}$
Table 8: The critical points and their associated eigenvalues of the linearised (Jacobian) matrix, for the case of scalar-field potential $V(\phi)=c_{0}+c_{1}e^{\sqrt{\frac{2\kappa^{2}}{3}}\phi}+c_{2}e^{2\sqrt{\frac{2\kappa^{2}}{3}}\phi}$ and interaction rate $Q=\alpha H\rho_{m}$. We have defined $T^{\prime}=-27+34\alpha+9\alpha^{2}$, $V^{\prime}=-11664-27891\alpha-16700\alpha^{2}-4362\alpha^{3}-540\alpha^{4}-27\alpha^{5}$, $R^{\prime}=-27+22\alpha+9\alpha^{2}$ and $S^{\prime}=-31347-57456\alpha-37514\alpha^{2}-11580\alpha^{3}-1755\alpha^{4}-108\alpha^{5}$.
  • The critical point $I_{1}$ exists for all $\alpha$, and it corresponds to a dark-energy dominated super-accelerated solution. It is stable in the region $\alpha>-\frac{17}{3}$ and thus it can describe the late-time phase of the universe.
  • Point $I_{2}$ in its stability region has $0\leq\Omega_{\phi}\leq 1$, and thus $0\leq\Omega_{m}\leq 1$. Hence, since $\Omega_{\phi}$ and $\Omega_{m}$ can be of the same order, this point can alleviate the coincidence problem [6, 7, 8, 9, 10, 11, 12, 13]. Moreover, for $\alpha=\frac{-13-\sqrt{88}}{3}$ or $\alpha=\frac{-33+\sqrt{88}}{3}$, it has $\Omega_{m}=1$, and in this case it corresponds to a matter-dominated solution.
  • Point $I_{3}$ exists always, it is saddle, and it corresponds to a dark-energy dominated super-accelerating solution.
    Point $x$ $v$ $w$ $\xi$ $z$ Existence Stability $\Omega_{\phi}$ $w_{tot}$
    $J_{1}$ $\frac{1}{3}$ 0 $\frac{\sqrt{10}}{3}$ 0 0 $\forall\gamma$ always saddle 1 $-\frac{11}{9}$
    $J_{2}$ $\frac{2}{3}$ 0 0 $\frac{\sqrt{13}}{3}$ 0 $\forall\gamma$ always stable 1 $-\frac{17}{9}$
    $J_{3}$ $\frac{1}{3}$ 0 $\frac{\sqrt{10}}{3}$ 0 1 $\forall\gamma$ always saddle 1 $-\frac{11}{9}$
    $J_{4}$ $\frac{2}{3}$ 0 0 $\frac{\sqrt{13}}{3}$ 1 $\forall\gamma$ always saddle 1 $-\frac{17}{9}$
    $J_{5}$ $x_{c}$ $\sqrt{1-{x_{c}}^{2}}$ 0 0 $\ \ \ \frac{3}{3-\gamma}$ $\ \ \ 0<|x_{c}|\leq\frac{1}{\sqrt{2}}$ stable for $-\frac{1}{\sqrt{2}}\leq x_{c}<0$ $\ \ \ 1-2{x_{c}}^{2}$ $-1$
    with $\gamma<0$
    Table 9: The critical points for the case of $V(\phi)=c_{0}+c_{1}e^{\sqrt{\frac{2\kappa^{2}}{3}}\phi}+c_{2}e^{2\sqrt{\frac{2\kappa^{2}}{3}}\phi}$ and interaction rate $Q=\Gamma\rho_{m}$, alongside their existence and stability conditions. Furthermore, we present the corresponding values of the dark-energy density parameter $\Omega_{\phi}$ and the total equation-of-state parameter $w_{tot}$.
  • Point $I_{4}$ has $0\leq\Omega_{\phi}\leq 1$ and $0\leq\Omega_{m}\leq 1$, and corresponds to accelerating solutions. However, since it is saddle it cannot offer an alleviation for the coincidence problem. Finally, in the limiting case where $\alpha=\frac{-10\pm\sqrt{19}}{3}$ we have $\Omega_{m}=1$, corresponding to a matter-dominated solution.
Finally, we observe that the system possesses two critical points at infinity, namely $G^{\pm\infty}\Bigl{\{}\left(y_{1},y_{2},y_{3},y_{4},y_{5}\right)\in\mathbb{S}^{4}:y_{5}=0,y_{4}>0,y_{3}>0,y_{2}>0,y_{1}=\pm\frac{1}{\sqrt{2}}\Bigr{\}}$, exhibiting unstable behavior and having $q\longrightarrow-\infty$. The details of the analysis are presented in Appendix D.1.
III.4.2 Model II: $Q=\Gamma\rho_{m}$
In this case, using the dimensionless variables $v$, $w$, $\xi$ defined in (49), the variable $z$ defined in (28) and the dimensionless variable $x$ from (21), the autonomous system can be written as
$x^{\prime}$ $=$ $-\frac{3}{2}x\left(1+x^{2}+v^{2}+w^{2}+\xi^{2}\right)+w^{2}+2\xi^{2}$ (56)
$-\frac{\gamma\left(1+x^{2}-v^{2}-w^{2}-\xi^{2}\right)z}{2x(1-z)},$
$v^{\prime}$ $=$ $\frac{3}{2}v\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right),$ (57)
$w^{\prime}$ $=$ $w\left[x+\frac{3}{2}\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right)\right],$ (58)
$\xi^{\prime}$ $=$ $\xi\left[2x+\frac{3}{2}\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right)\right],$ (59)
$z^{\prime}$ $=$ $\frac{3}{2}z(1-z)\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right),$ (60)
where $\gamma=\Gamma/H_{0}$ is the dimensionless coupling parameter. The system is invariant under the transformations $v\longrightarrow-v$, $w\longrightarrow-w$ and $\xi\longrightarrow-\xi$ separately, and thus the physical region is defined as
$\mathbb{D}$ $=$ $\bigg{\{}(x,v,w,\xi,z)\in\mathbb{R}^{5}:0\leq-x^{2}+v^{2}+w^{2}+\xi^{2}\leq 1,$ (61)
$v\geq 0,w\geq 0,\xi\geq 0,0\leq z\leq 1\bigg{\}}.$
There are two singular surfaces in the dynamical system, at $x=0$ and $z=1$ respectively, at which the system is ill-defined.
The critical points, the existence and stability conditions as well as the corresponding values of $\Omega_{\phi}$ and $w_{tot}$ are summarized in Table 9, while the corresponding eigenvalues are displayed in Table 10. For this scenario we have the following critical points.
Point $E_{1}$ $E_{2}$ $E_{3}$ $E_{4}$ $E_{5}$
$J_{1}$ $-\frac{11}{3}$ $-\frac{1}{3}$ $-\frac{10}{3}$ $\frac{1}{3}$ $-\frac{1}{3}$
$J_{2}$ $-\frac{17}{3}$ $-\frac{4}{3}$ $-\frac{2}{3}$ $-\frac{13}{3}$ $-\frac{4}{3}$
$J_{3}$ $-sgn(\gamma)\infty$ $-\frac{1}{3}$ $-\frac{10}{3}$ $\frac{1}{3}$ $\frac{1}{3}$
$J_{4}$ $-sgn(\gamma)\infty$ $-\frac{4}{3}$ $-\frac{2}{3}$ $-\frac{13}{3}$ $\frac{4}{3}$
$J_{5}$ 00 $x_{c}$ $2x_{c}$ $-3(1+\sqrt{1-{x_{c}}^{2}})$ $-3(1-\sqrt{1-{x_{c}}^{2}})$
Table 10: The critical points and their associated eigenvalues of the linearised (Jacobian) matrix, for the case of scalar-field potential $V(\phi)=c_{0}+c_{1}e^{\sqrt{\frac{2\kappa^{2}}{3}}\phi}+c_{2}e^{2\sqrt{\frac{2\kappa^{2}}{3}}\phi}$ and interaction rate $Q=\Gamma\rho_{m}$.
  • The critical point $J_{1}$ does not depend on the parameter $\gamma$, it exists always and it is saddle, corresponding to super-accelerated dark-energy dominated solution, with $H\longrightarrow\infty$.
  • Point $J_{2}$ exists always and it is stable, corresponding to super-accelerated dark-energy dominated solution, with $H\longrightarrow\infty$. Thus, $J_{2}$ can describe the late-time evolution of the universe.
  • Point $J_{3}$ exists always, it is saddle, and it also corresponds to super-accelerated dark-energy dominated solution, but with $H\longrightarrow 0$.
    Figure 4: Summary plot of the phase-space behavior and the critical points for the interaction models $Q=\alpha H\rho_{m}$ and $Q=\Gamma\rho_{m}$, considering all the examined potentials. Each row corresponds to a specific potential and color pattern (magenta, green, yellow, pink), and each column to a specific interaction model and shade pattern (dark shade, light shade). We only focus on late-time attractors. Black circles correspond to the late-time dark-energy dominated accelerating attractors, while red triangles indicate late-time accelerating scaling attractors. For the global interaction model $Q=\alpha H\rho_{m}$, all potentials can provide accelerating scaling attractors, while for local interaction rate $Q=\Gamma\rho_{m}$, only two potentials, namely $V(\phi)=\frac{1}{8\beta}\left(1-e^{-\kappa\mu\;\phi}\right)^{2}$ and $V(\phi)=c_{0}+c_{1}e^{\sqrt{\frac{2\kappa^{2}}{3}}\phi}+c_{2}e^{2\sqrt{\frac{2\kappa^{2}}{3}}\phi}$ can produce accelerating scaling attractors. Finally, note that all scaling solutions can transform to matter-dominated solution in the limiting case of the parameter region (see text).
  • Point $J_{4}$ exist always, it is saddle, and it also corresponds to a super-accelerated dark-energy dominated solution, with $H\longrightarrow 0$.
  • The curve of critical points $J_{5}$ corresponds to accelerating solutions. Since these points are normally hyperbolic, by applying the center manifold theorem we deduce that $J_{5}$ points are stable for $-\frac{1}{\sqrt{2}}\leq x_{c}<0$. Additionally, note that for this curve of critical points we have $0\leq\Omega_{\phi}\leq 1$ and $0\leq\Omega_{m}\leq 1$, and consequently $\Omega_{m}/\Omega_{\phi}\neq 0$, and hence $J_{5}$ can alleviate the coincidence problem. Finally, in the case $x_{c}=\pm\frac{1}{\sqrt{2}}$ one has $\Omega_{m}=1$, corresponding to a completely matter-dominated solution.
Lastly, the Poincaré compactification of the system results in two critical points at infinity, given by, $J^{\pm\infty}\Bigl{\{}\left(y_{1},y_{2},y_{3},y_{4},y_{5},y_{6}\right)\in\mathbb{S}^{5}:y_{6}=0,y_{5}=0,y_{4}>0,y_{3}>0,y_{2}>0,y_{1}=\pm\frac{1}{\sqrt{2}}\Bigr{\}}$, which exhibit unstable behavior, and having $q\longrightarrow-\infty$. The details are presented in Appendix D.2.
IV Summary and Conclusions
Cosmological scenarios which include interactions between dark matter and dark energy sectors have received significant attention in the scientific community since they exhibit many appealing features, such as the solution of the cosmic coincidence problem and the alleviation of $H_{0}$ and $S_{8}$ tensions. On the other hand, phantom cosmology has also attracted a significant amount of research, since it can efficiently describe the phantom crossing.
In the present work we performed a detailed investigation of interacting phantom cosmology, by applying the powerful method of dynamical system analysis, since this approach allows to extract information about the global features of the scenario, independently of the initial conditions or the specific evolution of the Universe. The novel ingredient of our work is the examination of new potentials, which indeed lead to new results even for the same interaction forms previously studied in the literature.
We considered two well-studied interaction forms, namely the global one $Q=\alpha H\rho_{m}$ and the local one $Q=\Gamma\rho_{m}$. Additionally, we considered four widely-used potentials for the phantom scalar field, namely A) $V(\phi)=V_{0}e^{-\kappa\lambda\phi}$, B) $V(\phi)=V_{0}\cosh(\kappa\eta\phi)$, C) $V(\phi)=\left(1-e^{-\kappa\mu\;\phi}\right)^{2}/8\beta$ and D) $V(\phi)=c_{0}+c_{1}e^{\sqrt{2\kappa^{2}/3}\phi}+c_{2}e^{2\sqrt{2\kappa^{2}/3}\phi}$, where the last two are the new ones that have not been studied before in the framework of interacting phantom cosmology. In all cases, we extracted the critical points and we examined their stability.
Our analysis extracted saddle matter-dominated points, stable dark-energy dominated points, and scaling accelerating solutions, that can attract the Universe at late times. As we discussed, some of the accelerating scaling attractors obtained for the last two potentials, when the local interaction is considered, are totally new. Furthermore, performing a detailed analysis at infinity it has been shown that there are no stable critical points at infinity, and thus they cannot attract the Universe at late times. For completeness, in Fig. 4 we present the summary diagram that provides the qualitative behavior of the examined interacting phantom scenarios for different potentials.
In summary, our analysis revealed that interacting phantom cosmology can describe the Universe evolution very efficiently, namely the transition from a matter-dominated saddle point to a stable dark-energy dominated accelerated universe, or a stable accelerating phase in which dark-energy and dark matter co-exist, and hence alleviating the cosmic coincidence problem. These features make interacting phantom cosmology a good candidate for the description of nature.
V Acknowledgments
The authors would like to thank an anonymous referee for important comments that significantly improved the quality of the manuscript. SH acknowledges the financial support from the University Grants Commission (UGC), Govt. of India (NTA Ref. No: 201610019097). SP and TS acknowledge the financial support from the Department of Science and Technology (DST), Govt. of India under the Scheme “Fund for Improvement of S&T Infrastructure (FIST)” (File No. SR/FST/MS-I/2019/41). ENS acknowledges the contribution of the LISA CosWG, and of COST Actions CA18108 ”Quantum Gravity Phenomenology in the multi-messenger approach” and CA21136” Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)”.
Appendix A Potential $V(\phi)=V_{0}e^{-\kappa\lambda\phi}$
A.1 Model I: $Q=\alpha H\rho_{m}$
The phase space domain of the autonomous system (25)-(26) is not compact in the $x-y$ direction. Thus, one may expect that there are critical points at infinity. Since the autonomous system is ill-defined at $x=0$ we divide the phase-space domain $\mathbb{D}$ into two parts: the first one for $x>0$ and the second one for $x<0$, i.e. we remove $x=0$ singular line from our phase space and we resort to Poincaré compactification technique [155, 154, 156, 121] to determine the qualitative behavior of the critical points at infinity.
In the part of the phase space where $x>0$, we use a re-parametrization, i.e., we introduce a new time variable $\tau$, defined by $dN=xd\tau$, and the system (25)-(26) simplifies as
$\frac{dx}{d\tau}$ $=$ $-3x^{2}-\frac{\sqrt{6}}{2}\lambda xy^{2}+\frac{3}{2}x^{2}\left(1-x^{2}-y^{2}\right)$ (62)
$-\frac{\alpha}{2}\left(1+x^{2}-y^{2}\right),$
$\frac{dy}{d\tau}$ $=$ $-\frac{\sqrt{6}}{2}\lambda x^{2}y+\frac{3}{2}xy\left(1-x^{2}-y^{2}\right).$ (63)
The vector field associated with the above system is defined as $(P,Q)$ where $P=-3x^{2}-\frac{\sqrt{6}}{2}\lambda xy^{2}+\frac{3}{2}x^{2}\left(1-x^{2}-y^{2}\right)-\frac{\alpha}{2}\left(1+x^{2}-y^{2}\right)$ and $Q=-\frac{\sqrt{6}}{2}\lambda x^{2}y+\frac{3}{2}xy\left(1-x^{2}-y^{2}\right)$. Now, we project this vector field on the upper hemisphere $\mathbb{S}^{2+}$ using the central projection $y_{1}=\frac{x}{\Delta x}$, $y_{2}=\frac{y}{\Delta x}$ and $y_{3}=\frac{1}{\Delta x}$, where $\Delta x=\sqrt{1+x^{2}+y^{2}}$. The stability analysis is performed in three different local charts, namely $\left(U_{i},\phi_{i}\right)$, where $i=1,2,3$, $U_{i}=\left\{(y_{1},y_{2},y_{3})\in\mathbb{S}^{2}:y_{i}>0\right\}$, and the local map $\phi_{i}:U_{i}\longrightarrow\mathbb{R}^{2}$ is defined as $\phi_{i}\left(y_{1},y_{2},y_{3}\right)=\left(\frac{y_{n}}{y_{i}},\frac{y_{m}}{y_{i}}\right)$ for $n<m$ and $n,m\neq i$.
In the first chart, we set $\phi_{1}\left(y_{1},y_{2},y_{3}\right)=(u,v)$, reducing the system (62)-(63) as
$\frac{du}{d\tau}$ $=$ $3uv^{2}+\frac{\sqrt{6}}{2}\lambda uv\left(u^{2}-1\right)$ (64)
$-\frac{\alpha}{2}uv^{2}\left(1-u^{2}+v^{2}\right),$
$\frac{dv}{d\tau}$ $=$ $3v^{3}+\frac{\sqrt{6}}{2}\lambda u^{2}v^{2}+\frac{3}{2}v\left(1+u^{2}-v^{2}\right)$ (65)
$\frac{\alpha}{2}v^{3}\left(1-u^{2}+v^{2}\right).$
Since, $\frac{du}{d\tau}=0$, $\frac{dv}{d\tau}=0$ at $v=0$, the critical points at infinity are $\{(u,0):u\in\mathbb{R}\}$, which correspond to the equator circle of $\mathbb{S}^{2}$ with $y_{1}>0$. For our phase-space domain, the point at infinity is the point $A^{+\infty}\left(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}},0\right)$, which further indicates the point $(1,0)$ in the $u-v$ plane. The eigenvalues at point $(1,0)$ are $3$ and 00, concluding that it is an unstable point i.e. $A^{+\infty}$ is unstable.
For the second chart we consider $\phi_{2}\left(y_{1},y_{2},y_{3}\right)=(u,v)$, which leads to the system
$\frac{du}{d\tau}$ $=$ $-3u^{2}v^{2}+\frac{\sqrt{6}}{2}\lambda uv\left(u^{2}-1\right)$ (66)
$-\frac{\alpha}{2}v^{2}\left(u^{2}+v^{2}-1\right),$
$\frac{dv}{d\tau}$ $=$ $\frac{\sqrt{6}}{2}\lambda u^{2}v^{2}-\frac{3}{2}uv\left(v^{2}-u^{2}-1\right).$ (67)
Note that here the curve $\{(u,0):u\in\mathbb{R}\}$ also corresponds to points at infinity, which is the equator circle of $\mathbb{S}^{2}$ with $y_{2}>0$. Once again, the point $A^{+\infty}\left(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}},0\right)$ on the equator circle is the point at infinity for our physical domain. The point $A^{+\infty}$ associates the point $(1,0)$ on the $u-v$ plane. The point $(1,0)$ is also unstable since one eigenvalue is $2$ and the other is zero. Thus, $A^{+\infty}$ exhibits unstable behavior.
In the third chart, if we assume $\phi_{2}\left(y_{1},y_{2},y_{3}\right)=(u,v)$, the dynamical system (62)-(63) becomes
$\frac{du}{d\tau}$ $=$ $P(u,v),$ (68)
$\frac{dv}{d\tau}$ $=$ $Q(u,v),$ (69)
which is identical to the system (62)-(63) and hence it cannot provide any new solutions at infinity.
When we deal with the phase-space domain $\mathbb{D}$ with $x<0$, we use the re-parametrization $dN=-xd\tau$, transforming the system (25)-(26) as
$\frac{dx}{d\tau}$ $=$ $3x^{2}+\frac{\sqrt{6}}{2}\lambda xy^{2}-\frac{3}{2}x^{2}\left(1-x^{2}-y^{2}\right)$ (70)
$+\frac{\alpha}{2}\left(1+x^{2}-y^{2}\right),$
$\frac{dy}{d\tau}$ $=$ $\frac{\sqrt{6}}{2}\lambda x^{2}y-\frac{3}{2}xy\left(1-x^{2}-y^{2}\right).$ (71)
By applying the same procedure used in the previous paragraphs, in this case we also extract a critical point at infinity, namely $A^{-\infty}\left(-\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}},0\right)$, which exhibits unstable behavior.
In summary, the original system (25)-(26) has two points at infinity, namely $A^{\pm\infty}\left(\pm\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}},0\right)$, which correspond to unstable solutions, and thus they cannot attract the Universe at late times.
A.2 Model II: $Q=\Gamma\rho_{m}$
The dynamical system (29)-(31) exhibits a singularity at $x=0$, making the system ill-defined there. Thus, in the part of the domain $\mathbb{D}$ where $x>0$, we utilize the re-parametrization $dN=x(1-z)d\tau$, simplifying the system (29)-(31) to the form
$\frac{dx}{d\tau}$ $=$ $-3x^{2}(1-z)+\frac{3}{2}x^{2}\left(1-x^{2}-y^{2}\right)(1-z)$ (72)
$-\frac{\sqrt{6}}{2}\lambda xy^{2}(1-z)-\frac{\gamma z\left(1+x^{2}-y^{2}\right)}{2},$
$\frac{dy}{d\tau}$ $=$ $\left[-\frac{\sqrt{6}}{2}\lambda x^{2}y+\frac{3}{2}xy\left(1-x^{2}-y^{2}\right)\right](1-z),$ (73)
$\frac{dz}{d\tau}$ $=$ $\frac{3}{2}xz(1-z)^{2}\left(1-x^{2}-y^{2}\right).$ (74)
The vector field related to the system mentioned above is defined as $(P,Q,R)$, in which $P=-3x^{2}(1-z)-\frac{\sqrt{6}}{2}\lambda xy^{2}(1-z)+\frac{3}{2}x^{2}\left(1-x^{2}-y^{2}\right)(1-z)-\frac{\gamma z\left(1+x^{2}-y^{2}\right)}{2}$, $Q=\left[-\frac{\sqrt{6}}{2}\lambda x^{2}y+\frac{3}{2}xy\left(1-x^{2}-y^{2}\right)\right](1-z)$ and $R=\frac{3}{2}xz(1-z)^{2}\left(1-x^{2}-y^{2}\right)$. Now, we proceed by projecting this vector field onto the upper three-sphere $\mathbb{S}^{3+}$ through central projection $y_{1}=\frac{x}{\Delta x}$, $y_{2}=\frac{y}{\Delta x}$, $y_{3}=\frac{z}{\Delta x}$ and $y_{4}=\frac{1}{\Delta x}$, where $\Delta x=\sqrt{1+x^{2}+y^{2}+z^{2}}$. The stability analysis is conducted in four distinct local charts, denoted as $\left(U_{i},\phi_{i}\right)$, where $i=1,2,3,4$ and $U_{i}=\left\{\left(y_{1},y_{2},y_{3},y_{4}\right)\in\mathbb{S}^{3}:y_{i}>0\right\}$, while the local map $\phi_{i}:U_{i}\longrightarrow\mathbb{R}^{3}$ is defined by $\phi_{i}\left(y_{1},y_{2},y_{3},y_{4}\right)=\left(\frac{y_{n}}{y_{i}},\frac{y_{m}}{y_{i}},\frac{y_{p}}{y_{i}}\right)$ for $n<m<p$ and $n,m,p\neq i$.
For the first chart, taking $\phi_{1}\left(y_{1},y_{2},y_{3},y_{3}\right)=(u,v,w)$, the dynamical system (72)-(74) becomes
$\frac{du}{d\tau}$ $=$ $3uw^{3}(w-v)+\frac{\gamma uvw^{3}(w^{2}+1-u^{2})}{2}$ (75)
$+\frac{\sqrt{6}}{2}uw^{2}(w-v)(u^{2}-1),$
$\frac{dv}{d\tau}$ $=$ $\frac{3}{2}vw(w-v)\left(1+u^{2}+w^{2}\right)$ (76)
$+\frac{\sqrt{6}}{2}\lambda u^{2}vw^{2}(w-v)+\frac{\gamma v^{2}w^{3}\left(w^{2}+1-u^{2}\right)}{2}$
$+\frac{3}{2}v(w-v)^{2}\left(w^{2}-1-u^{2}\right),$
$\frac{dw}{d\tau}$ $=$ $\frac{\sqrt{6}}{2}\lambda u^{2}w^{3}(w-v)+\frac{\gamma vw^{4}\left(w^{2}+1-u^{2}\right)}{2}$ (77)
$+\frac{3}{2}w^{2}(w-v)\left(1+u^{2}+w^{2}\right).$
As $\frac{du}{dt}=0$, $\frac{dv}{dt}=-\frac{3}{2}v^{3}\left(1+u^{2}\right)$ and $\frac{dw}{dt}=0$ at $w=0$, the critical points at infinity are given by $\left\{(u,0,0):u\in\mathbb{R}\right\}$, corresponding to the circle $\mathbb{S}^{1}$ with $y_{1}>0$. Taking into account the phase space domain, the critical point at infinity is the point $B^{+\infty}\left(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}},0,0\right)$, implying the point $(1,0,0)$ in the $(u,v,w)$ coordinate system. Linear stability analysis fails to predict the stability of the point $(1,0,0)$, since all eigenvalues at this point are zero. Nevertheless, at $v=0$ plane we have $\frac{du}{d\tau}=\frac{1}{2}uw^{3}\left(6w+\sqrt{6}\lambda\left(u^{2}-1\right)\right)$ and $\frac{dw}{d\tau}=\frac{1}{2}w^{3}\left(3\left(1+w^{2}\right)+u^{2}(3+\sqrt{6}\lambda w)\right)$. Hence, close to $w=0$, we have $\frac{dw}{d\tau}\cong\frac{3}{2}w^{3}\left(1+u^{2}\right)$, which states that $w$ is increasing for $w>0$ and decreasing for $w<0$. As a result, point $(1,0,0)$ exhibits unstable behavior and the corresponding point $B^{+\infty}$ is also unstable.
In the second chart we consider $\phi_{2}\left(y_{1},y_{2},y_{3},y_{4}\right)=(u,v,w)$, leading to the system
$\frac{du}{d\tau}$ $=$ $\frac{3}{2}u^{2}w(w-v-1)\left(w^{2}-u^{2}-1\right)-3u^{2}w^{3}(w-v)$ (78)
$+\frac{\sqrt{6}}{2}\lambda uw^{2}\left(u^{2}-1\right)-\frac{\gamma vw^{3}\left(w^{2}+u^{2}-1\right)}{2},$
$\frac{dv}{d\tau}$ $=$ $\frac{3}{2}uv(1-w)(w-v)\left(w^{2}-u^{2}-1\right)$ (79)
$+\frac{\sqrt{6}}{2}\lambda u^{2}vw^{2}(w-v),$
$\frac{dw}{d\tau}$ $=$ $\frac{\sqrt{6}}{2}\lambda u^{2}w^{3}-\frac{3}{2}uw^{2}(w-v)\left(w^{2}-u^{2}-1\right).$ (80)
Similarly, the points at infinity are described by the line $\left\{(u,0,0):u\in\mathbb{R}\right\}$, representing the circle $\mathbb{S}^{1}$ with $y_{1}>0$. Thus, the critical point at infinity is $\left(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}},0,0\right)$, which already has been obtained in the first chart.
In the third chart we consider the local map $\phi_{3}\left(y_{1},y_{2},y_{3},y_{4}\right)=(u,v,w)$. Therefore, the system (29)-(31) in terms of $u$, $v$ and $w$ can be expressed as
$\frac{du}{d\tau}$ $=$ $\frac{3}{2}u^{2}(w-1)\left(w^{2}-u^{2}-v^{2}\right)-\frac{\sqrt{6}}{2}\lambda uv^{2}w^{2}\left(w-1\right)$ (81)
$-3u^{2}w^{3}(w-1)-\frac{\gamma w^{3}\left(w^{2}+u^{2}-v^{2}\right)}{2},$
$\frac{dv}{d\tau}$ $=$ $\frac{3}{2}uv(w-1)\left(w^{2}-u^{2}-v^{2}\right)$ (82)
$-\frac{\sqrt{6}}{2}\lambda u^{2}vw^{2}(w-1),$
$\frac{dw}{d\tau}$ $=$ $-\frac{3}{2}uw(w-1)^{2}\left(w^{2}-u^{2}-v^{2}\right).$ (83)
As we can see, for this scenario, no critical points exist at infinity.
In the fourth chart, if we consider $\phi_{4}\left(y_{1},y_{2},y_{3},y_{4}\right)=(u,v,w)$, the dynamical system becomes:
$\frac{du}{d\tau}=P(u,v,w),$ (84)
$\frac{dv}{d\tau}=Q(u,v,w),$ (85)
$\frac{dw}{d\tau}=R(u,v,w).$ (86)
This is the same as the system (72)-(74), and it does not yield any new solutions at infinity.
When working with the phase-space domain $\mathbb{D}$ where $x<0$, we introduce a re-parametrization, defined by $dN=-x(1-z)d\tau$, which transforms the system (29)-(31) as
$\frac{dx}{d\tau}$ $=$ $3x^{2}(1-z)-\frac{3}{2}x^{2}\left(1-x^{2}-y^{2}\right)(1-z)$ (87)
$+\frac{\sqrt{6}}{2}\lambda xy^{2}(1-z)+\frac{\gamma z\left(1+x^{2}-y^{2}\right)}{2},$
$\frac{dy}{d\tau}$ $=$ $\left[\frac{\sqrt{6}}{2}\lambda x^{2}y-\frac{3}{2}xy\left(1-x^{2}-y^{2}\right)\right](1-z),$ (88)
$\frac{dz}{d\tau}$ $=$ $-\frac{3}{2}xz(1-z)^{2}\left(1-x^{2}-y^{2}\right).$ (89)
Following the same approach as in the previous paragraphs, we observe that there exists a critical point at infinity, namely $B^{-\infty}\left(-\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}},0,0\right)$, which exhibits unstable behavior.
In conclusion, the original system (29)-(31) has two points at infinity, namely $B^{\pm\infty}\left(\pm\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}},0,0\right)$, both representing unstable solutions, and thus they cannot attract the Universe at late times.
Appendix B Potential $V(\phi)=V_{0}\cosh{(\kappa\eta\phi)}$
B.1 Model I: $Q=\alpha H\rho_{m}$
For the phase-space domain $\mathbb{D}_{+}=\mathbb{D}\setminus\{x\leq 0\}$, the re-parametrization, defined as $dN=xd\tau$, transforms the systems (34)-(36) into the following:
$\frac{dx}{d\tau}$ $=$ $-3x^{2}-\frac{\sqrt{6}}{2}\lambda xy^{2}+\frac{3}{2}x^{2}\left(1-x^{2}-y^{2}\right)$ (90)
$-\frac{\alpha}{2}\left(1+x^{2}-y^{2}\right),$
$\frac{dy}{d\tau}$ $=$ $-\frac{\sqrt{6}}{2}\lambda x^{2}y+\frac{3}{2}xy\left(1-x^{2}-y^{2}\right),$ (91)
$\frac{d\lambda}{d\tau}$ $=$ $\sqrt{6}\left(\lambda^{2}-\eta^{2}\right)x^{2}.$ (92)
The vector field corresponding to the system described above is represented by the components $P$, $Q$ and $R$. These are defined as: $P=-3x^{2}-\frac{\sqrt{6}}{2}\lambda xy^{2}+\frac{3}{2}x^{2}\left(1-x^{2}-y^{2}\right)-\frac{\alpha}{2}\left(1+x^{2}-y^{2}\right)$, $Q=-\frac{\sqrt{6}}{2}\lambda x^{2}y+\frac{3}{2}xy\left(1-x^{2}-y^{2}\right)$ and $R=\sqrt{6}\left(\lambda^{2}-\eta^{2}\right)x^{2}$. We project the vector field onto the upper three-sphere $\mathbb{S}^{+}$ via central projection which is defined by the following relations for $y_{1}$, $y_{2}$, $y_{3}$ and $y_{4}$: $y_{1}=\frac{x}{\Delta x}$, $y_{2}=\frac{y}{\Delta x}$, $y_{3}=\frac{\lambda}{\Delta x}$ and $y_{4}=\frac{1}{\Delta x}$, where $\Delta x=\sqrt{1+x^{2}+y^{2}+\lambda^{2}}$. The stability analysis is then carried out in four distinct local charts, $\left(U_{i},\phi_{i}\right)$, labeled with $i=1,2,3,4$, and the sets $U_{i}$ are given by $U_{i}=\left\{\left(y_{1},y_{2},y_{3},y_{4}\right)\in\mathbb{S}^{3}:y_{i}>0\right\}$. The local map $\phi_{i}:U_{i}\longrightarrow\mathbb{R}^{3}$ is given by the relation $\phi_{i}\left(y_{1},y_{2},y_{3},y_{4}\right)=\left(\frac{y_{n}}{y_{i}},\frac{y_{m}}{y_{i}},\frac{y_{p}}{y_{i}}\right)$, with the indices $n$, $m$, $p$ satisfying $n<m<p$, and are all distinct from $i$.
For the first chart, by setting $\phi_{1}\left(y_{1},y_{2},y_{3},y_{4}\right)=(u,v,w)$, the dynamical system (90)-(92) transforms as
$\frac{du}{d\tau}$ $=$ $3uw^{2}+\frac{\alpha}{2}uw^{2}\left(1-u^{2}+w^{2}\right)$ (93)
$+\frac{\sqrt{6}}{2}uv\left(u^{2}-1\right),$
$\frac{dv}{d\tau}$ $=$ $\frac{3}{2}v\left(1+u^{2}+w^{2}\right)+\frac{\alpha}{2}vw^{2}\left(1-u^{2}+w^{2}\right)$ (94)
$+\frac{\sqrt{6}}{2}u^{2}v^{2}+\sqrt{6}\left(v^{2}-\eta^{2}w^{2}\right),$
$\frac{dw}{d\tau}$ $=$ $\frac{3}{2}w\left(1+u^{2}+w^{2}\right)+\frac{\alpha}{2}w^{3}\left(1-u^{2}+w^{2}\right)$ (95)
$+\frac{\sqrt{6}}{2}u^{2}vw.$
At $w=0$ we have $\frac{du}{d\tau}=\frac{\sqrt{6}}{2}uv\left(u^{2}-1\right)$, $\frac{dv}{d\tau}=\frac{3}{2}uv\left(\sqrt{6}uv+1+u^{2}\right)$ and $\frac{dw}{d\tau}=0$, which indicates that the points at infinity of the above system are $\left\{(u,0,0):u\in\mathbb{R}\right\}$, $\left\{(0,v,0):v\in\mathbb{R}\right\}$, $\left(1,-\sqrt{\frac{2}{3}},0\right)$ and $\left(-1,\sqrt{\frac{2}{3}},0\right)$. From the perspective of our phase-space domain, the critical points at infinity which lie on $\mathbb{S}^{3}$ are $C_{1}^{+\infty}\left(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}},0,0\right)$, and $C_{2}^{+\infty}\left(\sqrt{\frac{3}{8}},\sqrt{\frac{3}{8}},-\frac{1}{2},0\right)$, indicating respectively the points $(1,0,0)$ and $\left(1,-\sqrt{\frac{2}{3}},0\right)$ in the $(u,v,w)$ coordinate system. The eigenvalues at $(1,0,0)$ are $3,3,0$, concluding it is unstable in nature. Similarly, the eigenvalues evaluated at $\left(1,-\sqrt{\frac{2}{3}},0\right)$ are $-3,-2,2$, implying that this point is saddle. Therefore, $C_{1}^{+\infty}$ exhibits unstable nature, while $C_{2}^{+\infty}$ exhibits saddle behavior.
In the second chart we find two points at infinity, namely, $C_{1}^{+\infty}\left(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}},0,0\right)$, and $C_{2}^{+\infty}\left(\sqrt{\frac{3}{8}},\sqrt{\frac{3}{8}},-\frac{1}{2},0\right)$, which we have already seen in the first chart. Moreover, we cannot detect any points at infinity in the third chart for our phase space domain. Finally, the dynamical system in the fourth chart is completely analogous to the system (90)-(92), which does not lead to any new solutions at infinity.
In the phase-space domain $\mathbb{D}_{-}=\mathbb{D}\setminus\{x\geq 0\}$, the re-parametrization, defined as $dN=-xd\tau$, leads to a simplified version of the system of equations (34)-(31), namely
$\frac{dx}{d\tau}$ $=$ $3x^{2}+\frac{\sqrt{6}}{2}\lambda xy^{2}-\frac{3}{2}x^{2}\left(1-x^{2}-y^{2}\right)$ (96)
$+\frac{\alpha}{2}\left(1+x^{2}-y^{2}\right),$
$\frac{dy}{d\tau}$ $=$ $\frac{\sqrt{6}}{2}\lambda x^{2}y-\frac{3}{2}xy\left(1-x^{2}-y^{2}\right),$ (97)
$\frac{d\lambda}{d\tau}$ $=$ $-\sqrt{6}\left(\lambda^{2}-\eta^{2}\right)x^{2}.$ (98)
Following the same reasoning as in the previous paragraphs, we extract two critical points at infinity, namely $C_{1}^{-\infty}\left(-\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}},0,0\right)$ and $C_{2}^{-\infty}\left(-\sqrt{\frac{3}{8}},\sqrt{\frac{3}{8}},\frac{1}{2},0\right)$, and we deduce that $C_{1}^{-\infty}$ is unstable, while $C_{2}^{-\infty}$ is saddle.
To summarize, the original system (34)-(36) has four points at infinity, i.e. $C_{1}^{\pm\infty}\left(\pm\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}},0,0\right)$, and $C_{2}^{\pm\infty}\left(\pm\sqrt{\frac{3}{8}},\sqrt{\frac{3}{8}},\mp\frac{1}{2},0\right)$, which exhibit unstable dynamics, and thus they cannot attract the Universe at late times.
B.2 Model II: $Q=\Gamma\rho_{m}$
Making use of the re-parametrization, defined as $dN=x(1-z)d\tau$ for the phase-space domain $\mathbb{D}_{+}=\mathbb{D}\setminus\{x\leq 0\}$, the system of equations (37)-(40) becomes
$\frac{dx}{d\tau}$ $=$ $-3x^{2}(1-z)+\frac{3}{2}x^{2}\left(1-x^{2}-y^{2}\right)(1-z)$ (99)
$-\frac{\sqrt{6}}{2}\lambda xy^{2}(1-z)-\frac{\gamma z\left(1+x^{2}-y^{2}\right)}{2},$
$\frac{dy}{d\tau}$ $=$ $\left[-\frac{\sqrt{6}}{2}\lambda x^{2}y+\frac{3}{2}xy\left(1-x^{2}-y^{2}\right)\right](1-z),$ (100)
$\frac{d\lambda}{d\tau}$ $=$ $\sqrt{6}\left(\lambda^{2}-\eta^{2}\right)x^{2}(1-z),$ (101)
$\frac{dz}{d\tau}$ $=$ $\frac{3}{2}xz(1-z)^{2}\left(1-x^{2}-y^{2}\right).$ (102)
The vector field corresponding to the system is represented by the components $P$, $Q$, $R$ and $S$, defined as: $P=-3x^{2}(1-z)+\frac{3}{2}x^{2}\left(1-x^{2}-y^{2}\right)(1-z)-\frac{\sqrt{6}}{2}\lambda xy^{2}(1-z)-\frac{\gamma z\left(1+x^{2}-y^{2}\right)}{2}$, $Q=\left[-\frac{\sqrt{6}}{2}\lambda x^{2}y+\frac{3}{2}xy\left(1-x^{2}-y^{2}\right)\right](1-z)$, $R=\sqrt{6}\left(\lambda^{2}-\eta^{2}\right)x^{2}(1-z)$ and $S=\frac{3}{2}xz(1-z)^{2}\left(1-x^{2}-y^{2}\right)$. We start by projecting the vector field onto the upper four dimensional sphere $\mathbb{S}^{4+}$ using central projection, defined by the following relations for $y_{1}$, $y_{2}$, $y_{3}$, $y_{4}$, $y_{5}$: $y_{1}=\frac{x}{\Delta x}$, $y_{2}=\frac{y}{\Delta x}$, $y_{3}=\frac{\lambda}{\Delta x}$, $y_{4}=\frac{z}{\Delta x}$ and $y_{5}=\frac{1}{\Delta x}$, where $\Delta x=\sqrt{1+x^{2}+y^{2}+\lambda^{2}+z^{2}}$. The stability of the system is studied within five distinct local charts $\left(U_{i},\phi_{i}\right)$ indexed by $i=1,2,3,4,5$, where the sets $U_{i}$ are defined as $U_{i}=\left\{\left(y_{1},y_{2},y_{3},y_{4},y_{5}\right)\in\mathbb{S}^{4}:y_{i}>0\right\}$. The local map $\phi_{i}:U_{i}\longrightarrow\mathbb{R}^{4}$ is given by $\phi_{i}\left(y_{1},y_{2},y_{3},y_{4},y_{5}\right)=\left(\frac{y_{n}}{y_{i}},\frac{y_{m}}{y_{i}},\frac{y_{p}}{y_{i}},\frac{y_{q}}{y_{i}}\right)$, where $n$, $m$, $p$ and $q$ are distinct indices such that $n<m<p<q$, and none of $n,m,p,q$ are equal to $i$.
For the first chart, by selecting $\phi_{1}\left(y_{1},y_{2},y_{3},y_{4},y_{5}\right)=(u,v,w,r)$, the dynamical system described by equations (99)-(102) is modified as
$\frac{du}{d\tau}$ $=$ $3ur^{3}(r-w)+\frac{\gamma uwr^{3}(1-u^{2}+r^{2})}{2}$ (103)
$+\frac{\sqrt{6}}{2}uvr\left(u^{2}-1\right)(r-w),$
$\frac{dv}{d\tau}$ $=$ $\frac{3}{2}vr(r-w)\left(1+u^{2}+r^{2}\right)$ (104)
$+\frac{\sqrt{6}}{2}u^{2}v^{2}r(r-w)+\frac{\gamma vwr^{3}\left(1-u^{2}+r^{2}\right)}{2}$
$+\sqrt{6}r(r-w)\left(v^{2}-\eta^{2}r^{2}\right),$
$\frac{dw}{d\tau}$ $=$ $\frac{3}{2}wr(r-w)\left(1+u^{2}+r^{2}\right)$ (105)
$+\frac{\sqrt{6}}{2}u^{2}vwr(r-w)+\frac{\gamma w^{2}r^{3}\left(1-u^{2}+r^{2}\right)}{2}$
$+\frac{3}{2}w(r-w)^{2}\left(r^{2}-1-u^{2}\right),$
$\frac{dr}{d\tau}$ $=$ $\frac{3}{2}r^{2}(r-w)\left(1+u^{2}+r^{2}\right)+\frac{\sqrt{6}}{2}u^{2}vr^{2}(r-w)$ (106)
$+\frac{\gamma wr^{4}\left(1-u^{2}+r^{2}\right)}{2}.$
At $r=0$, the system evolves as $\frac{du}{d\tau}=0$, $\frac{dv}{d\tau}=0$, $\frac{dw}{d\tau}=-\frac{3}{2}w^{3}\left(1+u^{2}\right)$ and $\frac{dr}{d\tau}=0$. This indicates that the critical points at infinity are given by $\left\{(u,v,0,0):u\in\mathbb{R},v\in\mathbb{R}\right\}$. From the perspective of our phase-space domain, the critical points at infinity, located on the sphere $\mathbb{S}^{4}$, are $D^{+\infty}\Bigl{\{}\left(y_{1},y_{2},y_{3},y_{4},y_{5}\right)\in\mathbb{S}^{4}:y_{5}=0,y_{4}=0,y_{1}^{2}=y_{2}^{2},y_{1}>0,y_{2}>0\Bigr{\}}$. At the points $(u,v,0,0)$ all eigenvalues are zero, which suggests that a linear stability analysis does not provide conclusive information about stability. However, on the $w=0$ surface the system evolves as $\frac{du}{d\tau}=3r^{4}u+\sqrt{\frac{3}{2}}r^{2}uv\left(u^{2}-1\right)$, $\frac{dv}{d\tau}=\frac{1}{2}r^{2}\left(3v\left(1+r^{2}+u^{2}\right)+\sqrt{6}v^{2}\left(2+u^{2}\right)-2\sqrt{6}r^{2}\eta^{2}\right)$, $\frac{dw}{d\tau}=0$ and $\frac{dr}{d\tau}=\frac{1}{2}r^{3}\left(3+3r^{2}+u^{2}(3+\sqrt{6}v)\right)$. From these we observe that $r$ increases when $r>0$ with $v>0$, and decreases when $r<0$ with $v>0$. As a result, $D^{+\infty}$ exhibits unstable behavior.
The second and third charts reveal the same point at infinity, namely $D^{+\infty}$, as found in the first chart. Furthermore, the fourth chart does not uncover any additional critical points at infinity within our phase space domain. Finally, the fifth chart yields a dynamical system similar to (99)-(102), without generating any new solutions at infinity.
By using the re-parametrization $dN=-x(1-z)d\tau$ in the phase-space domain $\mathbb{D}_{-}=\mathbb{D}\setminus\{x\geq 0\}$, we obtain the following simplified form of the system (37)-(40), namely,
$\frac{dx}{d\tau}$ $=$ $3x^{2}(1-z)-\frac{3}{2}x^{2}\left(1-x^{2}-y^{2}\right)(1-z)$ (107)
$+\frac{\sqrt{6}}{2}\lambda xy^{2}(1-z)+\frac{\gamma z\left(1+x^{2}-y^{2}\right)}{2},$
$\frac{dy}{d\tau}$ $=$ $\left[\frac{\sqrt{6}}{2}\lambda x^{2}y-\frac{3}{2}xy\left(1-x^{2}-y^{2}\right)\right](1-z),$ (108)
$\frac{d\lambda}{d\tau}$ $=$ $-\sqrt{6}\left(\lambda^{2}-\eta^{2}\right)x^{2}(1-z),$ (109)
$\frac{dz}{d\tau}$ $=$ $-\frac{3}{2}xz(1-z)^{2}\left(1-x^{2}-y^{2}\right).$ (110)
Continuing with the same approach as in previous sections, we identify the critical points at infinity as $D^{-\infty}\Bigl{\{}\left(y_{1},y_{2},y_{3},y_{4},y_{5}\right)\in\mathbb{S}^{4}:y_{5}=0,y_{4}=0,y_{1}^{2}=y_{2}^{2},\\ y_{1}<0,y_{2}>0\Bigr{\}}$. This suggests that $D^{-\infty}$ also exhibits unstable characteristics.
In conclusion, the original system (37)-(40) has two points at infinity: $D^{+\infty}$ and $D^{-\infty}$, both of which exhibit unstable dynamics, and thus they cannot be the late-time solutions of the Universe.
Appendix C Potential $V(\phi)=\frac{1}{8\beta}\left(1-e^{-\kappa\mu\;\phi}\right)^{2}$
C.1 Model I: $Q=\alpha H\rho_{m}$
For the system of equations (42)-(44), the re-parametrization $dN=xd\tau$ in the phase-space domain $\mathbb{D}_{+}=\mathbb{D}\setminus\{x\leq 0\}$ yields the following:
$\frac{dx}{d\tau}$ $=$ $-3x^{2}-\frac{\sqrt{6}}{2}\lambda xy^{2}+\frac{3}{2}x^{2}\left(1-x^{2}-y^{2}\right)$ (111)
$-\frac{\alpha}{2}\left(1+x^{2}-y^{2}\right),$
$\frac{dy}{d\tau}$ $=$ $-\frac{\sqrt{6}}{2}\lambda x^{2}y+\frac{3}{2}xy\left(1-x^{2}-y^{2}\right),$ (112)
$\frac{d\lambda}{d\tau}$ $=$ $\frac{\sqrt{6}}{2}\left(\lambda^{2}-2\mu\lambda\right)x^{2}.$ (113)
The components $P$, $Q$ and $R$ represent the vector field corresponding to the system outlined above. These are given by $P=-3x^{2}-\frac{\sqrt{6}}{2}\lambda xy^{2}+\frac{3}{2}x^{2}\left(1-x^{2}-y^{2}\right)-\frac{\alpha}{2}\left(1+x^{2}-y^{2}\right)$, $Q=-\frac{\sqrt{6}}{2}\lambda x^{2}y+\frac{3}{2}xy\left(1-x^{2}-y^{2}\right)$ and $R=\frac{\sqrt{6}}{2}\left(\lambda^{2}-2\mu\lambda\right)x^{2}$. We begin by projecting the vector field onto the upper three dimensional sphere $\mathbb{S}^{3+}$ using central projection, defined by the following relations for $y_{1}$, $y_{2}$, $y_{3}$ and $y_{4}$: $y_{1}=\frac{x}{\Delta x}$, $y_{2}=\frac{y}{\Delta x}$, $y_{3}=\frac{\lambda}{\Delta x}$ and $y_{4}=\frac{1}{\Delta x}$. The stability is then analyzed within four distinct local charts $\left(U_{i},\phi_{i}\right)$ with indices $i=1,2,3,4$, where the sets $U_{i}$ are given by $U_{i}=\left\{\left(y_{1},y_{2},y_{3},y_{4}\right)\in\mathbb{S}^{3}:y_{i}>0\right\}$. The local map $\phi_{i}:U_{i}\longrightarrow\mathbb{R}^{3}$ is described by $\phi_{i}\left(y_{1},y_{2},y_{3},y_{4}\right)=\left(\frac{y_{n}}{y_{i}},\frac{y_{m}}{y_{i}},\frac{y_{p}}{y_{i}}\right)$, with the indices $n$, $m$ and $p$ satisfying $n<m<p$, with $n,m,p\neq i$.
In the case of the first chart, by choosing $\phi_{1}\left(y_{1},y_{2},y_{3},y_{4}\right)=(u,v,w)$ the dynamical system (111)-(113) becomes
$\frac{du}{d\tau}$ $=$ $3uw^{2}+\frac{\alpha}{2}uw^{2}\left(1-u^{2}+w^{2}\right)$ (114)
$+\frac{\sqrt{6}}{2}uv\left(u^{2}-1\right),$
$\frac{dv}{d\tau}$ $=$ $\frac{3}{2}v\left(1+u^{2}+w^{2}\right)+\frac{\alpha}{2}vw^{2}\left(1-u^{2}+w^{2}\right)$ (115)
$+\frac{\sqrt{6}}{2}u^{2}v^{2}+\frac{\sqrt{6}}{2}\left(v^{2}-2\mu vw\right),$
$\frac{dw}{d\tau}$ $=$ $\frac{3}{2}w\left(1+u^{2}+w^{2}\right)+\frac{\alpha}{2}w^{3}\left(1-u^{2}+w^{2}\right)$ (116)
$+\frac{\sqrt{6}}{2}u^{2}vw.$
At $w=0$, the system evolves according to $\frac{du}{d\tau}=\frac{\sqrt{6}}{2}uv\left(u^{2}-1\right)$, $\frac{dv}{d\tau}=\frac{1}{2}v\left(1+u^{2}\right)(\sqrt{6}v+3)$ and $\frac{dw}{d\tau}=0$. This suggests that the critical points at infinity are $\left\{(u,0,0):u\in\mathbb{R}\right\}$, $\left(0,-\sqrt{\frac{3}{2}},0\right)$, $\left(1,-\sqrt{\frac{3}{2}},0\right)$ and $\left(-1,-\sqrt{\frac{3}{2}},0\right)$. From the viewpoint of our phase space domain, the critical points at infinity located on $\mathbb{S}^{3}$ are $E_{1}^{+\infty}\left(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}},0,0\right)$ and $E_{2}^{+\infty}\left(\sqrt{\frac{2}{7}},\sqrt{\frac{2}{7}},-\sqrt{\frac{3}{7}},0\right)$, corresponding respectively to the points $(1,0,0)$ and $\left(1,-\sqrt{\frac{3}{2}},0\right)$ in $(u,v,w)$ coordinates. The eigenvalues for the point $(1,0,0)$ are $3,3$ and 00, which indicates an unstable nature. Additionally, at point $\left(1,-\sqrt{\frac{3}{2}},0\right)$ the eigenvalues are $-3,-3$ and $\frac{3}{2}$, suggesting that it is a saddle point. Consequently, $E_{1}^{+\infty}$ is unstable, and $E_{2}^{+\infty}$ behaves as a saddle point.
In the second chart we again find two points at infinity, namely $E_{1}^{+\infty}\left(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}},0,0\right)$ and $E_{2}^{+\infty}\left(\sqrt{\frac{2}{7}},\sqrt{\frac{2}{7}},-\sqrt{\frac{3}{7}},0\right)$, which were identified in the first chart. The third chart does not reveal any additional points at infinity in our phase-space domain. The fourth chart gives a dynamical system that is similar to the system (111)-(113), and therefore it does not provide any new solutions at infinity.
In the phase space domain $\mathbb{D}_{-}=\mathbb{D}\setminus\{x\geq 0\}$, we employ the re-parametrization $dN=-xd\tau$ and obtain a simplified version of the system (42)-(44) as
$\frac{dx}{d\tau}$ $=$ $3x^{2}+\frac{\sqrt{6}}{2}\lambda xy^{2}-\frac{3}{2}x^{2}\left(1-x^{2}-y^{2}\right)$ (117)
$+\frac{\alpha}{2}\left(1+x^{2}-y^{2}\right),$
$\frac{dy}{d\tau}$ $=$ $\frac{\sqrt{6}}{2}\lambda x^{2}y-\frac{3}{2}xy\left(1-x^{2}-y^{2}\right),$ (118)
$\frac{d\lambda}{d\tau}$ $=$ $-\frac{\sqrt{6}}{2}\left(\lambda^{2}-2\mu\lambda\right)x^{2}.$ (119)
Continuing with the same approach as in previous sections, we find two critical points at infinity: $E_{1}^{-\infty}\left(-\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}},0,0\right)$ and $E_{2}^{-\infty}\left(-\sqrt{\frac{2}{7}},\sqrt{\frac{2}{7}},\sqrt{\frac{3}{7}},0\right)$. This suggests that $E_{1}^{-\infty}$ is unstable, while $E_{2}^{-\infty}$ is saddle.
To conclude, the original system (42)-(44) has four points at infinity: $E_{1}^{\pm\infty}\left(\pm\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}},0,0\right)$, and $E_{2}^{\pm\infty}\left(\pm\sqrt{\frac{2}{7}},\sqrt{\frac{2}{7}},\mp\sqrt{\frac{3}{7}},0\right)$, all of which display unstable dynamics. Therefore, they cannot be the late-time solutions of the Universe.
C.2 Model II: $Q=\Gamma\rho_{m}$
After using a re-parametrization of the form $dN=x(1-z)d\tau$ in the phase-space domain $\mathbb{D}_{+}=\mathbb{D}\setminus\{x\leq 0\}$ for the system of equations (45)-(48), we get
$\frac{dx}{d\tau}$ $=$ $-3x^{2}(1-z)+\frac{3}{2}x^{2}\left(1-x^{2}-y^{2}\right)(1-z)$ (120)
$-\frac{\sqrt{6}}{2}\lambda xy^{2}(1-z)-\frac{\gamma z\left(1+x^{2}-y^{2}\right)}{2},$
$\frac{dy}{d\tau}$ $=$ $\left[-\frac{\sqrt{6}}{2}\lambda x^{2}y+\frac{3}{2}xy\left(1-x^{2}-y^{2}\right)\right](1-z),$ (121)
$\frac{d\lambda}{d\tau}$ $=$ $\frac{\sqrt{6}}{2}\left(\lambda^{2}-2\mu\lambda\right)x^{2}(1-z),$ (122)
$\frac{dz}{d\tau}$ $=$ $\frac{3}{2}xz(1-z)^{2}\left(1-x^{2}-y^{2}\right).$ (123)
The components $P$, $Q$, $R$ and $S$ represent the vector field of the above system, and are expressed as $P=-3x^{2}(1-z)+\frac{3}{2}x^{2}\left(1-x^{2}-y^{2}\right)(1-z)-\frac{\sqrt{6}}{2}\lambda xy^{2}(1-z)-\frac{\gamma z\left(1+x^{2}-y^{2}\right)}{2}$, $Q=\left[-\frac{\sqrt{6}}{2}\lambda x^{2}y+\frac{3}{2}xy\left(1-x^{2}-y^{2}\right)\right](1-z)$, $R=\frac{\sqrt{6}}{2}\left(\lambda^{2}-2\mu\lambda\right)x^{2}(1-z)$ and $S=\frac{3}{2}xz(1-z)^{2}\left(1-x^{2}-y^{2}\right)$. We first project the vector field onto the upper four dimensional sphere $\mathbb{S}^{4+}$ via central projection, where the relations for $y_{1}$, $y_{2}$, $y_{3}$, $y_{4}$ and $y_{5}$ are given by: $y_{1}=\frac{x}{\Delta x}$, $y_{2}=\frac{y}{\Delta x}$, $y_{3}=\frac{\lambda}{\Delta x}$, $y_{4}=\frac{z}{\Delta x}$ and $y_{5}=\frac{1}{\Delta x}$, where $\Delta x=\sqrt{1+x^{2}+y^{2}+\lambda^{2}+z^{2}}$. To analyze the stability, we use five distinct local charts $\left(U_{i},\phi_{i}\right)$, where $i=1,2,3,4,5$ and the sets $U_{i}$ are defined as $U_{i}=\left\{\left(y_{1},y_{2},y_{3},y_{4},y_{5}\right)\in\mathbb{S}^{4}:y_{i}>0\right\}$. The local map $\phi_{i}:U_{i}\longrightarrow\mathbb{R}^{4}$ is specified by $\phi_{i}\left(y_{1},y_{2},y_{3},y_{4},y_{5}\right)=\left(\frac{y_{n}}{y_{i}},\frac{y_{m}}{y_{i}},\frac{y_{p}}{y_{i}},\frac{y_{q}}{y_{i}}\right)$, where the indices $n$, $m$, $p$ and $q$ satisfy $n<m<p<q$, while $n,m,p,q\neq i$.
In the first chart, by defining $\phi_{1}\left(y_{1},y_{2},y_{3},y_{4},y_{5}\right)=(u,v,w,r)$, the dynamical system (120)-(123) takes the form
$\frac{du}{d\tau}$ $=$ $\frac{1}{2}ru\Bigl{[}6r^{3}+\sqrt{6}rv\left(u^{2}-1\right)-\sqrt{6}vw\left(u^{2}-1\right)$ (124)
$+\gamma r^{4}w+r^{2}w\left(\gamma-6-\gamma u^{2}\right)\Bigr{]},$
$\frac{dv}{d\tau}$ $=$ $\frac{1}{2}rv\Bigl{[}3r^{3}+(r-w)\left(3+u^{2}(3+\sqrt{6}v)\right)$ (125)
$+\gamma wr^{4}-wr^{2}\left(3+\gamma\left(u^{2}-1\right)\right)$
$+\sqrt{6}(r-w)(v-2\mu r)\Bigr{]},$
$\frac{dw}{d\tau}$ $=$ $\frac{1}{2}w\Bigl{[}6r^{4}+wr\left(3+u^{2}(3-\sqrt{6}v)\right)$ (126)
$-3w^{2}\left(1+u^{2}\right)+r^{2}\left(3w^{2}+\sqrt{6}u^{2}v\right)$
$+\gamma wr^{5}+wr^{3}\left(\gamma-9-\gamma u^{2}\right)\Bigr{]},$
$\frac{dr}{d\tau}$ $=$ $\frac{1}{2}r^{2}\Bigl{[}3r^{3}+(r-w)\left(3+u^{2}(3+\sqrt{6}v)\right)$ (127)
$+\gamma wr^{4}+wr^{2}\left(\gamma-3-\gamma u^{2}\right)\Bigr{]}.$
When $r=0$, the system evolves according to the set of equations $\frac{du}{d\tau}=0$, $\frac{dv}{d\tau}=0$, $\frac{dw}{d\tau}=-\frac{3}{2}w^{3}\left(1+u^{2}\right)$ and $\frac{dr}{d\tau}=0$. This suggests that the points at infinity are $\left\{(u,v,0,0):u\in\mathbb{R},v\in\mathbb{R}\right\}$. Regarding our phase-space domain, the critical points at infinity on the four-sphere $\mathbb{S}^{4}$ are given by $F^{+\infty}\Bigl{\{}\left(y_{1},y_{2},y_{3},y_{4},y_{5}\right)\in\mathbb{S}^{4}:y_{5}=0,y_{4}=0,y_{1}^{2}=y_{2}^{2},y_{1}>0,y_{2}>0\Bigr{\}}$. At the points $(u,v,0,0)$ all eigenvalues are zero, indicating that linear stability analysis is inconclusive. However, on the surface $w=0$, the system evolves according to the equations $\frac{du}{d\tau}=3r^{4}u+\sqrt{\frac{3}{2}}r^{2}uv\left(u^{2}-1\right)$, $\frac{dv}{d\tau}=\frac{1}{2}r^{2}v\left(3r^{2}+\left(1+u^{2}\right)(3+\sqrt{6}v)-2\sqrt{6}\mu r\right)$, $\frac{dw}{d\tau}=0$ and $\frac{dr}{d\tau}=\frac{1}{2}r^{3}\left(3+3r^{2}+u^{2}(3+\sqrt{6}v)\right)$. This leads to the conclusion that $r$ increases for $r>0$ and $v>0$, and decreases for $r<0$ and $v>0$. Consequently, $F^{+\infty}$ demonstrates unstable behavior.
In the second and third charts we identify $F^{+\infty}$ as the critical point at infinity for our phase space domain, which was already observed in the first chart. The fourth chart does not present any new points at infinity. The fifth chart leads to a dynamical system that resembles system (120)-(123), offering no new solutions at infinity, too.
By using the re-parametrization $dN=-x(1-z)d\tau$ in the phase-space domain $\mathbb{D}_{-}=\mathbb{D}\setminus\{x\geq 0\}$ through $dN=-x(1-z)d\tau$, we write the system (45)-(48) as
$\frac{dx}{d\tau}$ $=$ $3x^{2}(1-z)-\frac{3}{2}x^{2}\left(1-x^{2}-y^{2}\right)(1-z)$ (128)
$+\frac{\sqrt{6}}{2}\lambda xy^{2}(1-z)+\frac{\gamma z\left(1+x^{2}-y^{2}\right)}{2},$
$\frac{dy}{d\tau}$ $=$ $\left[\frac{\sqrt{6}}{2}\lambda x^{2}y-\frac{3}{2}xy\left(1-x^{2}-y^{2}\right)\right](1-z),$ (129)
$\frac{d\lambda}{d\tau}$ $=$ $-\frac{\sqrt{6}}{2}\left(\lambda^{2}-2\mu\lambda\right)x^{2}(1-z),$ (130)
$\frac{dz}{d\tau}$ $=$ $-\frac{3}{2}xz(1-z)^{2}\left(1-x^{2}-y^{2}\right).$ (131)
Following the same method as in the previous sections, we determine the critical points at infinity as $F^{-\infty}\Bigl{\{}\left(y_{1},y_{2},y_{3},y_{4},y_{5}\right)\in\mathbb{S}^{4}:y_{5}=0,y_{4}=0,y_{1}^{2}=y_{2}^{2},\\ y_{1}<0,y_{2}>0\Bigr{\}}$. This indicates that $F^{-\infty}$ is unstable.
In summary, the original system (45)-(48) exhibits two points at infinity, namelt $F^{+\infty}$ and $F^{-\infty}$, each of which demonstrates unstable behavior, and thus cannot be the late-time stages of the Universe.
Appendix D Potential $V(\phi)=c_{0}+c_{1}e^{\sqrt{\frac{2\kappa^{2}}{3}}\phi}+c_{2}e^{2\sqrt{\frac{2\kappa^{2}}{3}}\phi}$
D.1 Model I: $Q=\alpha H\rho_{m}$
By introducing the re-parametrization $dN=xd\tau$ in the phase space domain $\mathbb{D}_{+}=\mathbb{D}\setminus\{x\leq 0\}$, the system (51)-(54) becomes
$\frac{dx}{d\tau}$ $=$ $-\frac{3}{2}x^{2}\left(1+x^{2}+v^{2}+w^{2}+\xi^{2}\right)+xw^{2}+2x\xi^{2}$ (132)
$-\frac{\alpha\left(1+x^{2}-v^{2}-w^{2}-\xi^{2}\right)}{2},$
$\frac{dv}{d\tau}$ $=$ $\frac{3}{2}xv\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right),$ (133)
$\frac{dw}{d\tau}$ $=$ $xw\left[x+\frac{3}{2}\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right)\right],$ (134)
$\frac{d\xi}{d\tau}$ $=$ $x\xi\left[2x+\frac{3}{2}\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right)\right].$ (135)
The system outlined above is described by the vector field components $P$, $Q$, $R$ and $S$, which are given by: $P=-\frac{3}{2}x^{2}\left(1+x^{2}+v^{2}+w^{2}+\xi^{2}\right)+xw^{2}+2x\xi^{2}-\frac{\alpha\left(1+x^{2}-v^{2}-w^{2}-\xi^{2}\right)}{2}$, $Q=\frac{3}{2}xv\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right)$, $R=xw\left[x+\frac{3}{2}\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right)\right]$ and $S=x\xi\left[2x+\frac{3}{2}\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right)\right]$. The projection of the vector field onto the upper four dimensional sphere $\mathbb{S}^{4+}$ is performed using central projection, with the following relations defining $y_{1}$, $y_{2}$, $y_{3}$, $y_{4}$ and $y_{5}$: $y_{1}=\frac{x}{\Delta x}$, $y_{2}=\frac{v}{\Delta x}$, $y_{3}=\frac{w}{\Delta x}$, $y_{4}=\frac{\xi}{\Delta x}$ and $y_{5}=\frac{1}{\Delta x}$, where $\Delta x=\sqrt{1+x^{2}+v^{2}+w^{2}+\xi^{2}}$. For the stability analysis we consider five distinct local charts $\left(U_{i},\phi_{i}\right)$, indexed by $i=1,2,3,4,5$, with the sets $U_{i}$ defined as $U_{i}=\left\{\left(y_{1},y_{2},y_{3},y_{4},y_{5}\right)\in\mathbb{S}^{4}:y_{i}>0\right\}$. The corresponding local map $\phi_{i}:U_{i}\longrightarrow\mathbb{R}^{4}$ is described by $\phi_{i}\left(y_{1},y_{2},y_{3},y_{4},y_{5}\right)=\left(\frac{y_{n}}{y_{i}},\frac{y_{m}}{y_{i}},\frac{y_{p}}{y_{i}},\frac{y_{q}}{y_{i}}\right)$, where the indices $n$, $m$, $p$, $q$ follow the condition $n<m<p<q$, and no index among $n,m,p,q$ equals $i$.
By choosing $\phi_{1}\left(y_{1},y_{2},y_{3},y_{4},y_{5}\right)=\left(x_{1},x_{2},x_{3},x_{4}\right)$ for the first chart, the dynamical system (132)-(135) is simplified into
$\frac{dx_{1}}{d\tau}$ $=$ $\frac{1}{2}x_{1}x_{4}\Bigl{[}6x_{4}-2\left(x_{2}^{2}+2x_{3}^{2}\right)+\alpha x_{4}^{3}$ (136)
$+\alpha x_{4}\left(1-x_{1}^{2}-x_{2}^{2}-x_{3}^{2}\right)\Bigr{]},$
$\frac{dx_{2}}{d\tau}$ $=$ $\frac{1}{2}x_{2}x_{4}\Bigl{[}6x_{4}-2\left(x_{2}^{2}+2x_{3}^{2}-1\right)+\alpha x_{4}^{3}$ (137)
$+\alpha x_{4}\left(1-x_{1}^{2}-x_{2}^{2}-x_{3}^{2}\right)\Bigr{]},$
$\frac{dx_{3}}{d\tau}$ $=$ $\frac{1}{2}x_{3}x_{4}\Bigl{[}6x_{4}-2\left(x_{2}^{2}+2x_{3}^{2}-2\right)+\alpha x_{4}^{3}$ (138)
$+\alpha x_{4}\left(1-x_{1}^{2}-x_{2}^{2}-x_{3}^{2}\right)\Bigr{]},$
$\frac{dx_{4}}{d\tau}$ $=$ $\frac{1}{2}x_{4}\Bigl{[}3\left(1+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}\right)-2x_{4}\left(x_{2}^{2}+2x_{3}^{2}\right)$ (139)
$+\alpha x_{4}^{4}+\alpha x_{4}^{2}\left(3+\left(1-x_{2}^{2}-x_{2}^{2}-x_{3}^{2}\right)\right)\Bigr{]}.$
At $x_{4}=0$, the system behaves as $\frac{dx_{1}}{d\tau}=0$, $\frac{dx_{2}}{d\tau}=0$, $\frac{dx_{3}}{d\tau}=0$ and $\frac{dx_{4}}{d\tau}=0$. This leads to the conclusion that the points at infinity are $\left\{\left(x_{1},x_{2},x_{3},0\right):x_{1},x_{2},x_{3}\in\mathbb{R}\right\}$. From the viewpoint of our phase-space domain, the critical points at infinity, located on the sphere $\mathbb{S}^{4}$, are $G^{+\infty}\Bigl{\{}\left(y_{1},y_{2},y_{3},y_{4},y_{5}\right)\in\mathbb{S}^{4}:y_{5}=0,y_{4}>0,y_{3}>0,y_{2}>0,y_{1}=\frac{1}{\sqrt{2}}\Bigr{\}}$. At the points $\left(x_{1},x_{2},x_{3},0\right)$ three eigenvalues are zero, while one eigenvalue is positive, expressed as $\frac{3}{2}\left(1+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}\right)$. This implies that these points exhibit unstable behavior.
In the second, third and fourth charts, we again obtain points $G^{+\infty}$ at infinity, which were identified in the first chart. Similarly, the fifth chart results in a dynamical system similar to (132)-(135), without offering any new solutions at infinity.
By using the re-parametrization $dN=-xd\tau$ in the phase space domain $\mathbb{D}_{-}=\mathbb{D}\setminus\{x\geq 0\}$, we write the system (51)-(54) as
$\frac{dx}{d\tau}$ $=$ $\frac{3}{2}x^{2}\left(1+x^{2}+v^{2}+w^{2}+\xi^{2}\right)-xw^{2}-2x\xi^{2}$ (140)
$+\frac{\alpha\left(1+x^{2}-v^{2}-w^{2}-\xi^{2}\right)}{2},$
$\frac{dv}{d\tau}$ $=$ $-\frac{3}{2}xv\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right),$ (141)
$\frac{dw}{d\tau}$ $=$ $-xw\left[x+\frac{3}{2}\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right)\right],$ (142)
$\frac{d\xi}{d\tau}$ $=$ $-x\xi\left[2x+\frac{3}{2}\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right)\right].$ (143)
By applying the same approach used in previous sections, we extract the critical points at infinity as $G^{-\infty}\Bigl{\{}\left(y_{1},y_{2},y_{3},y_{4},y_{5}\right)\in\mathbb{S}^{4}:y_{5}=0,y_{4}>0,y_{3}>0,y_{2}>0,y_{1}=-\frac{1}{\sqrt{2}}\Bigr{\}}$. This suggests that $G^{-\infty}$ also exhibits unstable behavior.
In summary, the original system (51)-(54) has two points at infinity, namely, $G^{+\infty}$ and $G^{-\infty}$, both of which display unstable dynamics, and thus they cannot attract the Universe at late times.
D.2 Model II: $Q=\Gamma\rho_{m}$
The system of equations (56)-(60) is simplified by introducing the re-parametrization $dN=x(1-z)d\tau$ for the phase space domain $\mathbb{D}_{+}=\mathbb{D}\setminus\{x\leq 0\}$, as
$\frac{dx}{d\tau}$ $=$ $(1-z)\Bigl{[}-\frac{3}{2}x^{2}\left(1+x^{2}+v^{2}+w^{2}+\xi^{2}\right)+xw^{2}$ (144)
$+2x\xi^{2}\Bigr{]}-\frac{\alpha\left(1+x^{2}-v^{2}-w^{2}-\xi^{2}\right)}{2},$
$\frac{dv}{d\tau}$ $=$ $\frac{3}{2}xv(1-z)\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right),$ (145)
$\frac{dw}{d\tau}$ $=$ $xw\Bigl{[}x+\frac{3}{2}\Bigl{(}1-x^{2}-v^{2}-w^{2}$ (146)
$-\xi^{2}\Bigr{)}\Bigr{]}(1-z),$
$\frac{d\xi}{d\tau}$ $=$ $x\xi(1-z)\Bigl{[}2x+\frac{3}{2}\Bigl{(}1-x^{2}-v^{2}$ (147)
$-w^{2}-\xi^{2}\Bigr{)}\Bigr{]}.$
$\frac{dz}{d\tau}$ $=$ $\frac{3}{2}xz(1-z)^{2}\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right).$ (148)
The vector field for the above system is defined by the components $P$, $Q$, $R$, $S$ and $T$, which are specified as: $P=(1-z)\Bigl{[}-\frac{3}{2}x^{2}\left(1+x^{2}+v^{2}+w^{2}+\xi^{2}\right)+xw^{2}+2x\xi^{2}\Bigr{]}-\frac{\alpha\left(1+x^{2}-v^{2}-w^{2}-\xi^{2}\right)}{2}$, $Q=\frac{3}{2}xv(1-z)\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right)$, $R=xw\Bigl{[}x+\frac{3}{2}\Bigl{(}1-x^{2}-v^{2}-w^{2}-\xi^{2}\Bigr{)}\Bigr{]}(1-z)$, $S=x\xi(1-z)\Bigl{[}2x+\frac{3}{2}\Bigl{(}1-x^{2}-v^{2}-w^{2}-\xi^{2}\Bigr{)}\Bigr{]}$ and $T=\frac{3}{2}xz(1-z)^{2}\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right)$. We begin by projecting the vector field onto the upper five dimensional sphere $\mathbb{S}^{5+}$ using central projection, where the coordinates $y_{1}$, $y_{2}$, $y_{3}$, $y_{4}$, $y_{5}$, $y_{6}$ are related to $x$, $v$, $w$, $\xi$, $z$ through the following formulas: $y_{1}=\frac{x}{\Delta x}$, $y_{2}=\frac{v}{\Delta x}$, $y_{3}=\frac{w}{\Delta x}$, $y_{4}=\frac{\xi}{\Delta x}$, $y_{5}=\frac{z}{\Delta x}$ and $y_{6}=\frac{1}{\Delta x}$, where $\Delta x=\sqrt{1+x^{2}+v^{2}+w^{2}+\xi^{2}+z^{2}}$. The stability of the system is analyzed using six distinct local charts $\left(U_{i},\phi_{i}\right)$, where $i=1,2,3,4,5,6$, and the sets $U_{i}$ are defined by $U_{i}=\left\{\left(y_{1},y_{2},y_{3},y_{4},y_{5},y_{6}\right)\in\mathbb{S}^{5}:y_{i}>0\right\}$. The local map $\phi_{i}:U_{i}\longrightarrow\mathbb{R}^{5}$ is expressed as $\phi_{i}\left(y_{1},y_{2},y_{3},y_{4},y_{5},y_{6}\right)=\left(\frac{y_{n}}{y_{i}},\frac{y_{m}}{y_{i}},\frac{y_{p}}{y_{i}},\frac{y_{q}}{y_{i}},\frac{y_{r}}{y_{i}}\right)$, with the indices $n$, $m$, $p$, $q$, $r$ satisfying $n<m<p<q<r$, while $n,m,p,q,r\neq i$.
For the first chart, with the selection $\phi_{1}\left(y_{1},y_{2},y_{3},y_{4},y_{5},y_{6}\right)=(x_{1},x_{2},x_{3},x_{4},x_{5})$, the dynamical system described by (144)-(148) becomes
$\frac{dx_{1}}{d\tau}$ $=$ $\frac{1}{2}x_{5}^{2}x_{1}\Bigl{\{}2x_{5}\left(3x_{5}-x_{2}^{2}-2x_{3}^{2}\right)+x_{4}\Bigl{\{}2(x_{2}^{2}+2x_{3}^{2})$ (149)
$+\gamma x_{5}^{3}-x_{5}\left[6+\gamma\left(x_{1}^{2}+x_{2}^{2}+x_{3}^{2}-1\right)\right]\Bigr{\}}\Bigr{\}},$
$\frac{dx_{2}}{d\tau}$ $=$ $\frac{1}{2}x_{5}^{2}x_{2}\Bigl{\{}2x_{5}\left(1+3x_{5}-x_{2}^{2}-2x_{3}^{2}\right)$ (150)
$+x_{4}\Bigl{\{}2(x_{2}^{2}+2x_{3}^{2}-1)$
$+\gamma x_{5}^{3}-x_{5}\left[6+\gamma\left(x_{1}^{2}+x_{2}^{2}+x_{3}^{2}-1\right)\right]\Bigr{\}}\Bigr{\}},$
$\frac{dx_{3}}{d\tau}$ $=$ $\frac{1}{2}x_{5}^{2}x_{3}\Bigl{\{}2x_{5}\left(2+3x_{5}-x_{2}^{2}-2x_{3}^{2}\right)$ (151)
$+x_{4}\Bigl{\{}2(x_{2}^{2}+2x_{3}^{2}-2)$
$+\gamma x_{5}^{3}-x_{5}\left[6+\gamma\left(x_{1}^{2}+x_{2}^{2}+x_{3}^{2}-1\right)\right]\Bigr{\}}\Bigr{\}},$
$\frac{dx_{4}}{d\tau}$ $=$ $\frac{1}{2}x_{4}\Bigl{\{}2x_{5}^{3}\left(3x_{5}-x_{2}^{2}-2x_{3}^{2}\right)+$ (152)
$3x_{4}^{2}\left(x_{5}^{2}-1-x_{1}^{2}-x_{2}^{2}-x_{3}^{2}\right)$
$+x_{4}x_{5}\Bigl{\{}3\left(1+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}\right)+2x_{5}\left(x_{2}^{2}+2x_{3}^{2}\right)$
$+\gamma x_{5}^{4}-x_{5}^{2}\left[9+\gamma\left(x_{1}^{2}+x_{2}^{2}+x_{3}^{2}-1\right)\right]\Bigr{\}}\Bigr{\}},$
$\frac{dx_{5}}{d\tau}$ $=$ $\frac{1}{2}x_{5}^{2}\Bigl{\{}x_{5}\left(3x_{5}^{2}+3\left(1+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}\right)\right.$ (153)
$\left.-2x_{5}\left(x_{2}^{2}+2x_{3}^{2}\right)\right)$
$+x_{4}\Bigl{\{}-3\left(1+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}\right)+2x_{5}\left(x_{2}^{2}+2x_{3}^{2}\right)$
$+\gamma x_{5}^{4}-x_{5}^{2}\left[3+\gamma\left(x_{1}^{2}+x_{2}^{2}+x_{3}^{2}-1\right)\right]\Bigr{\}}\Bigr{\}}.$
For $x_{5}=0$, the system evolves according to the following differential equations: $\frac{dx_{1}}{d\tau}=0$, $\frac{dx_{2}}{d\tau}=0$, $\frac{dx_{3}}{d\tau}=0$, $\frac{dx_{4}}{d\tau}=-\frac{3}{2}x_{4}^{3}\left(1+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}\right)$ and $\frac{dx_{5}}{d\tau}=0$. This implies that the points at infinity are $\left\{\left(x_{1},x_{2},x_{3},0,0\right):x_{1},x_{2},x_{3}\in\mathbb{R}\right\}$. In the context of our phase-space domain, the critical points at infinity located on the sphere $\mathbb{S}^{5}$ are given by $J^{+\infty}\Bigl{\{}\left(y_{1},y_{2},y_{3},y_{4},y_{5},y_{6}\right)\in\mathbb{S}^{5}:y_{6}=0,y_{5}=0,y_{4}>0,y_{3}>0,y_{2}>0,y_{1}=\frac{1}{\sqrt{2}}\Bigr{\}}$. At the points $\left(x_{1},x_{2},x_{3},0,0\right)$, all eigenvalues are zero, implying that the linear stability analysis does not provide conclusive results regarding stability. However, on the surface where $x_{4}=0$, the system follows the dynamics $\frac{dx_{1}}{d\tau}=x_{1}x_{5}^{3}\left(3x_{5}-x_{2}^{2}-2x_{3}^{2}\right)$, $\frac{dx_{2}}{d\tau}=x_{2}x_{5}^{3}\left(1+3x_{5}-x_{2}^{2}-2x_{3}^{2}\right)$, $\frac{dx_{3}}{d\tau}=x_{3}x_{5}^{3}\left(2+3x_{5}-x_{2}^{2}-2x_{3}^{2}\right)$, $\frac{dx_{4}}{d\tau}=0$ and $\frac{dx_{5}}{d\tau}=\frac{1}{2}x_{5}^{3}\left(3x_{5}^{2}+3\left(1+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}\right)-2x_{5}\left(x_{2}^{2}+2x_{3}^{2}\right)\right)$, which suggest that near $x_{5}=0$, we have $\frac{dx_{5}}{d\tau}\cong\frac{3}{2}x_{5}^{3}\left(1+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}\right)$, showing that $x_{5}$ is decreasing for $x_{5}<0$, and increasing for $x_{5}>0$. Therefore, $F^{+\infty}$ exhibits unstable behavior.
In the second, third and fourth charts, we find $J^{+\infty}$ (which has already been found as a critical point at infinity in the first chart) as the critical point at infinity. The fifth chart does not introduce any new critical points at infinity within our phase-space domain. The sixth chart leads to a dynamical system similar to (144)-(148), without providing any new solutions at infinity, too.
In the phase-space domain $\mathbb{D}_{-}=\mathbb{D}\setminus\{x\geq 0\}$, the re-parametrization $dN=-x(1-z)d\tau$, transforms the system of equations
(56)-(60) as
$\frac{dx}{d\tau}$ $=$ $(1-z)\Bigl{[}\frac{3}{2}x^{2}\left(1+x^{2}+v^{2}+w^{2}+\xi^{2}\right)-xw^{2}$ (154)
$-2x\xi^{2}\Bigr{]}+\frac{\alpha\left(1+x^{2}-v^{2}-w^{2}-\xi^{2}\right)}{2},$
$\frac{dv}{d\tau}$ $=$ $-\frac{3}{2}xv(1-z)\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right),$
$\frac{dw}{d\tau}$ $=$ $-xw\Bigl{[}x+\frac{3}{2}\Bigl{(}1-x^{2}-v^{2}-w^{2}$ (155)
$-\xi^{2}\Bigr{)}\Bigr{]}(1-z),$
$\frac{d\xi}{d\tau}$ $=$ $-x\xi(1-z)\Bigl{[}2x+\frac{3}{2}\Bigl{(}1-x^{2}-v^{2}$ (156)
$-w^{2}-\xi^{2}\Bigr{)}\Bigr{]},$
$\frac{dz}{d\tau}$ $=$ $-\frac{3}{2}xz(1-z)^{2}\left(1-x^{2}-v^{2}-w^{2}-\xi^{2}\right).$ (157)
Continuing with the same methodology as in previous sections, the critical points at infinity are found to be: $J^{-\infty}\Bigl{\{}\left(y_{1},y_{2},y_{3},y_{4},y_{5},y_{6}\right)\in\mathbb{S}^{5}:y_{6}=0,y_{5}=0,y_{4}>0,y_{3}>0,y_{2}>0,y_{1}=-\frac{1}{\sqrt{2}}\Bigr{\}}$. This suggests that $J^{-\infty}$ also displays unstable behavior.
In summary, the original system (56)-(60) has two points at infinity, namely $J^{+\infty}$ and $J^{-\infty}$, which are both unstable, and therefore they cannot be the late-time solutions of the Universe.